### v1

# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

Abstract
Retrieval-augmented generation (RAG) extends large language models (LLMs) with external data sources to enhance factual correctness and domain coverage. Modern RAG pipelines rely on large datastores, leading to system challenges in latency-sensitive deployments, especially when limited GPU memory is available. To address these challenges, we propose TeleRAG, an efficient inference system that reduces RAG latency with minimal GPU memory requirements. The core innovation of TeleRAG is lookahead retrieval, a prefetching mechanism that anticipates required data and transfers it from CPU to GPU in parallel with LLM generation. By leveraging the modularity of RAG pipelines, the inverted file index (IVF) search algorithm and similarities between queries, TeleRAG optimally overlaps data movement and computation. Experimental results show that TeleRAG reduces end-to-end RAG inference latency by up to 1.72× on average compared to state-of-the-art systems, enabling faster, more memory-efficient deployments of advanced RAG applications.
Equal contribution. Work done during internship at UW. Correspondence to: Chien-Yu Lin ¡cyulin@cs.washington.edu¿.
1 Introduction
(a) Typical pipeline stages of a RAG application.
(b) Different scenarios for RAG and the illustration of the proposed lookahead retrieval mechanism. It prefetches relevant data for retrieval from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU-CPU cooperation.
Figure 1: (a) Illustrations of RAG pipeline stages. (b) Overview of TeleRAG and comparison to baseline systems.
Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases[29, 13, 57]. During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations [63, 74, 50] and incorporate up-to-date or private information [36, 64].
Modern RAG applications share two key characteristics. (1) RAG applications are built as modular pipelines as shown in Figure 1(a), where a single query undergoes multiple rounds of LLM calls and retrievals, each of which serves different functions to improve overall output quality. For example, query transformation rounds [27, 100, 71, 10] generally occur before retrieval to refine the user’s query with LLMs (pre-retrieval generation). (2) RAG’s datastores are typically large, supported by recent works demonstrating that increasing the size of the datastore positively affects the performance of RAG applications [64, 50, 15, 33, 77].
These characteristics create significant challenges for efficient RAG inference, especially in latency-sensitive applications such as customer chatbots [18, 7, 83], financial analysis [62, 65], and emergency medical diagnosis [53, 31]. First, the latency of RAG systems is inherently increased due to multiple rounds of LLM generation and retrieval. Second, although GPU can speed up both LLM and retrieval, the combined memory demand of these processes often exceeds available GPU resources, making it expensive or infeasible to fully host both operations on GPU. Consequently, in local or small-scale deployments, which are common for RAG applications handling private or sensitive data, the retrieval datastore is frequently off-loaded to CPU memory to alleviate GPU constraints. However, while CPU off-loading resolves memory limitations, it significantly increases retrieval latency, diminishing overall system efficiency.
Addressing the latency challenge, several recent works have been proposed to accelerate RAG inference. These approaches include FPGA-based retrieval acceleration [41], reuse of KV-cache [44, 60, 92], and speculative document retrieval [87, 98]. However, none of these works directly address the substantial memory demands associated with retrieval, which remains a significant bottleneck for efficient RAG deployments.
In this paper, we identify a key opportunity in the widely adopted inverted file index (IVF) [80] retrieval method. IVF reduces retrieval latency by pre-clustering the datastore into sub-clusters, limiting the search space to relevant clusters of the retrieval query during runtime. This approach also offers a way to reduces GPU memory usage, as only a subset of clusters are needed for each retrieval query. For example, after identifying relevant clusters, their data can be dynamically transferred from CPU to GPU memory and accelerate the similarity searches on GPU. However, this approach incurs high CPU-GPU data transfer overhead, keeping the overall retrieval latency still high [49] (see §3).
Our proposal. To tackle latency and memory bottlenecks in RAG inference, we introduce a novel mechanism called lookahead retrieval. Lookahead retrieval predicts which subsets of IVF clusters will likely be accessed during the retrieval phase and proactively transfers them from CPU to GPU during the pre-retrieval generation stage.
The idea behind lookahead retrieval is based on the observation that queries before and after the pre-retrieval generation stage share substantial semantic overlap. Consequently, IVF clusters relevant to the initial input query are also likely to be relevant to the refined query. To validate this intuition, we rigorously analyze IVF cluster assignments across original queries and refined queries produced by the pre-retrieval generation step, evaluating on six representative RAG pipelines and three datasets. Our analysis confirms a strong similarity between these two query forms (see §3).
Leveraging this insight, we propose TeleRAG, an efficient inference system designed to optimize RAG latency while minimizing GPU memory consumption. TeleRAG employs lookahead retrieval to preemptively load relevant IVF clusters onto the GPU, effectively hiding CPU-GPU data transfer overhead during concurrent LLM generation. As illustrated in Figure 1(b), this approach significantly reduces retrieval latency without exceeding GPU memory constraints, enabling efficient execution of RAG applications in resource-limited environments.
Although TeleRAG significantly reduces latency by processing prefetched IVF clusters on the GPU, some relevant clusters might still be missed during prefetching. To maintain retrieval accuracy, TeleRAG concurrently executes similarity searches for these missed clusters on the CPU. The results from CPU searches are seamlessly merged with GPU results, ensuring complete retrieval precision. Thanks to the high cluster similarity between the initial and refined queries, the number of missed clusters is typically small, allowing TeleRAG to achieve low retrieval latency through efficient CPU-GPU coordination.
A key challenge in deploying lookahead retrieval is determining the optimal number of IVF clusters to prefetch. Prefetching too many clusters increases data-transfer overhead, whereas fetching too few clusters could result in higher retrieval latency due to increased CPU processing. To address this, we propose a profile-guided approach coupled with an analytical model that dynamically determines the ideal prefetch amount based on pipeline characteristics and hardware configurations. Our adaptive method consistently outperforms static configurations, delivering substantial latency improvements across diverse RAG setups and hardware environments.
Results summary. We evaluate TeleRAG using a 61GB Wikipedia-based datastore [5] across six popular RAG pipelines built with the Llama model family [82] (3B, 8B, and 13B). Remarkably, TeleRAG supports retrieval from a 61GB datastore alongside a Llama3-8B (16GB) LLM on a single RTX 4090 GPU (24GB memory), significantly outperforming the CPU retrieval baseline. Experiments conducted on an RTX 4090 and an H100 GPU [68] demonstrate speedups of up to 2.68x and 2.49x, with average speedups of 1.70x and 1.75x, respectively. These results highlight the capability of TeleRAG to efficiently handle large-scale RAG inference tasks within tight GPU memory constraints, confirming its practical value for resource-limited deployments.
In summary, we make the following key contributions:
  • Analyzing the correlation of the queries between the pre-retrieval generation and retrieval stages, revealing significant overlap in their corresponding IVF clusters.
  • Proposing lookahead retrieval, which prefetches likely IVF clusters to the GPU, and hides CPU-GPU data transfer time during pre-retrieval generation.
  • Developing TeleRAG, an efficient RAG inference system that integrates lookahead retrieval, resulting in significant acceleration of RAG with minimal GPU memory usage.
2 Background
Figure 2: Overview of RAG.
2.1 RAG
RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text [29, 13, 57]. The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations [63, 74, 50] and generate contents based on up-to-date or private information [36, 64].
RAG workflow. A basic RAG workflow involves several phases, including data store building, retrieval, and interactions with LLMs [24, 34, 23, 88, 84, 70, 28]. In order to build a data store, raw data in various formats is cleaned, converted to plain text, and divided into chunks. These chunks are then encoded into vectors using an embedding model and stored in a vector index, enabling efficient searching based on similarity. When a user provides a query, it is converted into a vector using the same embedding model, and the most similar chunks from the indexed database are retrieved. For a large-scale index, approximate algorithms such as the inverted file index (IVF) [80] (see §2.2 for detail) are commonly used to accelerate the search process. The retrieved chunks, along with the original user query, are combined and given as a prompt to the LLM. The LLM then generates a response that relies on the information provided in the retrieved documents. This approach enables dynamic and informative responses in both single- and multi-turn dialogues [57].
Modularity in modern RAG applications. However, naively integrating document retrieval into LLM generation can cause several issues. For example, the retrieval often struggles to find relevant content and may select irrelevant or mismatched chunks [27, 100, 61], and a single retrieval may not provide sufficient context, necessitating multiple retrieval rounds [29].
To solve these issues, most state-of-the-art RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  [29, 26, 34, 28, 23, 88]. Typically, they have the following types of steps (shown in Figure 2):
  • Pre-retrieval generation, used to assess whether retrieval is needed or to generate queries for retrieval. An example of a pre-retrieval technique is query transformation [61, 37, 27, 102, 93, 73, 72, 100, 58], which reformulates the original query to make it clearer and more suitable for the retrieval task.
  • Retrieval, used to identify relevant documents from the vector data store. This stage takes the output of pre-retrieval generation and generates data for the next stage.
  • Post-retrieval generation, generates response based on user query and retrieved documents. It can also perform additional process such as summarization [39, 51] or reranking[105, 81] on the retrieved documents.
  • Judgment, dynamically determines the execution flow. For example, it decides if more iteration is needed to enhance the response. Heuristics or LLMs can be used for judgement stage.
By proceeding through these functions, RAG applications can deliver more precise and contextually appropriate responses, significantly enhancing the capabilities of LLMs for various applications [45, 29].
2.2 Vector Index and Inverted File (IVF)
The vector index is a crucial component of RAG applications that retrieves similar items from large datasets of high-dimensional vectors. Given the query vector $x\in\mathrm{R}^{D}$ and vector database $Y=\{y_{0},\ldots,y_{N-1}\}\subset\mathrm{R}^{D}$, which comprises $N$ vectors, the vector index aims to find the $k$ nearest neighbors of $x$ from the database:
[…]
where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product [46].
To improve search efficiency, the inverted file index (IVF) algorithm is widely used for large-scale vector databases due to its simplicity and effectiveness. IVF partitions the data store into clusters and restricts searches to the most relevant clusters, reducing the computational cost. To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means [66] and partitions the database into $N_{c}$ clusters:
$\{C_{1},C_{2},\ldots,C_{N_{c}}\}=k\text{-means}(Y,N_{c}),$
where $C_{j}$ is the set of vectors assigned to the $j$-th cluster, and the cluster centroids $\{c_{1},c_{2},\ldots,c_{N_{c}}\}$ are obtained by taking the mean of each vectors in $\{C_{1},C_{2},\ldots,C_{N_{c}}\}$. Then, each database vector $y_{i}\in Y$ is assigned to the nearest cluster center:
$\text{Cluster}(y_{i})=\text{argmin}_{j=1:N_{c}}d(y_{i},c_{j}).$
With the trained centroids and cluster assignment, we can perform a more efficient index search. There are two steps involved in the IVF index searching. First, it will identify the nearest $L$ cluster centroids of a query vector $x$:
$\{c_{j_{1}},c_{j_{2}},\ldots,c_{j_{L}}\}=L\text{-argmin}_{j=1:C}\ d(x,c_{j}).$
This step is also referred to as the coarse-grained search in IVF, as it identifies candidates at the cluster level. Second, the fine-grained search is performed in the $L$ nearest clusters and identify $k$ closest vectors of $x$:
$k\text{-argmin}_{y_{i}\in\cup_{l=1}^{L}C_{j_{l}}}d(x,y_{i}).$
which involves sorting. In this way, IVF reduces search space and accelerates the retrieval process. Here, the hyperparameter $L$ from the first step is also referred as nprobe [25]. When nprobe is larger, the IVF retrieval will be more accurate as it search more clusters. However, the retrieve latency is longer as more computation and data are needed.
Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs. Open-source libraries offer efficient GPU implementations [46, 75]. However, IVF does not reduce the memory requirement for the index since the data for all clusters must be stored. As a result, IVF still requires a substantial GPU memory footprint, which becomes a bottleneck when GPU memory capacity is the constraint of RAG systems. Moreover, recent work reported that increasing the size of the data store positively affects RAG application performance [64, 50, 15, 33, 77], exacerbating this issue.
2.3 Latency-Sensitive RAG Use Cases
In this paper, we focus on reducing inference latency for a single query, a critical objective in latency-sensitive RAG applications. These applications span a range of real-world scenarios, such as customer chatbots [18, 7, 83], financial analysis [62, 65], autonomous driving [17, 94], and emergency medical diagnosis [53, 31], where users expect near-instantaneous responses. In each of these contexts, maintaining low latency per query is paramount for providing timely and accurate results.
Batching multiple queries, a common strategy for improving throughput, is not considered in this work. It is because many RAG applications involve user-specific private data [86, 55, 21, 89, 30] and serve relatively small user bases, making batching infeasible. For large-scale RAG services [2, 1, 10], although batching is feasible, it increases the latency for each individual queries [103, 97, 79] and therefore, is not ideal for the latency sensitive applications.
A straightforward way to reduce inference latency is to scale up GPU resources, allowing both the LLM and retrieval datastore to reside in GPU memory. However, this approach is often cost-prohibitive and could lead to significant hardware underutilization, particularly in the single-query scenarios that this work targets. Therefore, in this paper, we aim to improve single-query RAG latency without significantly increasing GPU memory usage, ensuring that both cloud-based and local deployments can satisfy stringent performance requirements without incurring excessive costs.
3 Analyzing Latency of RAG Pipelines
In this section, we analyze state-of-the-art RAG applications and identify their underlying systems challenges in achieving low latency. To conduct the analysis, we construct a 61GB vector index with the FAISS library [25], and set the IVF clusters to 4096 following common practice [11]. We use Llama-3-8B [6] as the LLM, and run our analysis on an H100 GPU with 80GB memory. §5.2 provides details about our experimental setup.
3.1 End-to-end Latency of RAG Pipelines
Figure 3: Latency breakdown of six RAG pipelines on NQ dataset [54].
Figure 4: The breakdown of memory consumption at GPU and CPU for two different strategies: CPU offloading and GPU-based retrieval. The dotted line indicates the memory capacity of a RTX4090 GPU, which is a common used GPU for local deployment.
We begin by analyzing the end-to-end latency of six representative RAG pipelines (see §5.1 for details) in two scenarios. In the first, the LLM runs on the GPU while retrieval is offloaded to the CPU, minimizing GPU memory usage. In the second, both the LLM and the vector index reside in GPU memory, enabling GPU-based retrieval for lower latency. If the GPU’s memory is insufficient, then CPU offloading is necessary.
We set the nprobe to 512, a common used setting under this index’s configuration (see §5.2 for details), to measure retrieval latency. Figure 3 shows the breakdown of end-to-end latency into LLM computation and retrieval. The results are the average of 512 random data from the NQ dataset [54]. We observe that CPU-based retrieval (first scenario) is the primary bottleneck, consuming about 56% of the total execution time on average, while GPU-based retrieval (second scenario) accounts for just 12%. On average, GPU retrieval is 12.5x faster than CPU retrieval, reducing overall latency by approximately 2.2x. Thus, accelerating retrieval on GPU is crucial for improving end-to-end latency.
However, GPU acceleration comes with a significant memory cost. Figure 4 shows the memory requirements for both GPU and CPU. As Figure 4 shows, putting both LLM weights and the retrieval index on the GPU requires around 77GB of memory, which exceeds the capacity of consumer GPUs like the RTX 4090 with 24GB. Thus, GPU acceleration on retrieval is often not feasible when running on a lower-end GPU or indexing with a large datastore. Note that the memory footprint of the LLM and the index can be even larger and further exacerbate the memory constraint.
In the rest of this section, we try to answer the following question: Is it possible to achieve the latency of GPU-based retrieval while using much less GPU memory?
3.2 Opportunities of GPU-accelerated Retrieval with Limited Memory
A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on-demand for each query, leveraging the IVF index structure that narrows the target search clusters (§2.2). While such a method enables faster search on the GPU, the bottleneck of retrieval is shifted to the data fetching.
To examine the performance, we reconstruct the IVF index in PyTorch [8], leveraging its high-performance GPU operator, and implement such an on-demand fetching retrieval system. We measure the retrieval latency with an RTX 4090 GPU and compare it with the baseline of CPU-based retrieval. Figure 5 shows the breakdown of retrieval time into data fetch and search time. We show the performance for different amounts of data fetched (indicated by the nprobe parameter). We observe that the fetch time dominates the latency due to limited PCIe bandwidth between the CPU and GPU (32 GB/s). Despite the fact that GPU retrieval is significantly faster than CPU retrieval, due to the fetch overhead, the effective latency is slower than CPU retrieval (around 3% slow down on average across nprobe’s). Therefore, we must hide the data fetching latency in order to achieve a meaningful speedup to following such approach.
Hiding the data fetch cost. Hiding the data fetch cost implies that CPU-to-GPU data fetch must be done before the retrieval operation is performed. However, predicting which data to retrieve in advance is challenging since the exact data requirements only become clear at retrieval time once we have the query produced by the pre-retrieval process (§2.1).
Fortunately, as we show next, the original form of a query before pre-retrieval can serve as a valuable hint for the clusters required to be retrieved at the end of the pre-retrieval stage. This hint can guide us in predicting the clusters that should be fetched, and the data transfer can be initiated in the background along with pre-retrieval to hide its latency.
Figure 5: Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
Dataset HyDE SubQ Iter IRG FLARE S-RAG
[27] [58] [58] [78] [43] [11]
NQ [54] 75.7% 66.2% 93.4% 85.7% 80.3% 100.0%
HotpotQA [91] 78.6% 66.6% 86.9% 91.2% 82.2% 100.0%
TriviaQA [47] 76.8% 65.3% 86.8% 87.5% 82.9% 100.0%
Table 1: IVF cluster overlapping rate between the input and output of the pre-retrieval generation. The nprobe is set to 512. Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
3.3 Overlapping of IVF Clusters
While the exact data to be retrieved is only known after the pre-retrieval generation is done, we observe that there are high similarities of the IVF clusters assignments between the queries at different stages.
Similarity of queries at different stages. During the pre-retrieval process of RAG pipelines (e.g., query transformation [61, 37, 27, 102, 93, 73, 72, 100, 58]), an LLM call is issued to refine an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$, which is then used for retrieval. This process often rewrites the query into a different format [27] or simplifies its complexities [58], which intuitively preserves the query’s core semantic content. Hence, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar. This similarity, in turn, suggests that the IVF clusters to which these queries would be assigned will overlap. Therefore, $q_{\text{in}}$ can serve as a valuable hint for predicting subsequent memory accesses.
Prediction coverage. To verify this hypothesis, we evaluate the average cluster coverage rate between prefetched clusters and clusters used for retrieval in three popular question-answer datasets (NQ [54], HotpotQA [91], and TriviaQA [47]) and six RAG pipelines. Table 1 shows the coverage when we prefetch $512$ clusters. From the table, we observe that IVF cluster overlap rates are consistently high across a range of datasets and pipelines. For instance, even the lowest reported values remain above 65% (for SubQ).
Opportunity. This data shows an opportunity for predicting required clusters, which can make it possible to hide data transfer overhead during LLM generation. In this paper, we aim to leverage this observation to accelerate the inference latency for RAG.
4 Design of TeleRAG
Figure 6: (a) The overview of TeleRAG’s lookahead retrieval and comparison to the baseline CPU-offloaded retrieval. (b) System design of TeleRAG. It first loads the entire index data into CPU memory and identifies the clusters ($C_{\text{in}}$) for the query before the transformation ($q_{\text{in}}$). It then transfers the data of $C_{\text{in}}$ to GPU memory while running the pre-retrieval stage and generating the retrieval query $q_{\text{out}}$. During the search, the overlapped clusters are searched on GPU, and the missed clusters are searched on CPU. The retrieval results are then merged and passed to the next stage.
Based on the high IVF cluster overlapping observation presented in § 3.3, we present TeleRAG, an efficient inference system for RAG that incorporates the idea of lookahead retrieval, which predicts and prefetches likely IVF cluster data to GPU and significantly accelerates the retrieval process with minimal GPU memory requirement. In this section, we describe our design overview (§4.1), an analytical model on finding the optimal prefetch data amount (§4.3), and the implementation details (§4.4).
4.1 Overview
Figure 6 presents the overview of TeleRAG with lookahead retrieval. In Figure 6, the input to the pre-retrieval stage is defined as $q_{\text{in}}$, and the output is denoted as $q_{\text{out}}$. In the subsequent retrieval stage, $q_{\text{out}}$ serves as the input query to perform the retrieval. We highlight the IVF clusters corresponding to $q_{\text{in}}$ with a red background ($C_{\text{in}}$), and those corresponding to $q_{\text{out}}$ with a blue background ($C_{\text{out}}$). Due to the semantic similarity of $q_{\text{in}}$ and $q_{\text{out}}$, there is significant overlap between $C_{\text{in}}$ and $C_{\text{out}}$3.3), marked with a purple background ($C_{\text{overlap}}$).
Given this, the lookahead retrieval technique in TeleRAG operates in three key steps. First, $C_{\text{in}}$ is computed, and their data is transferred to GPU memory during pre-retrieval generation, beginning with those clusters whose centroids are closest to $q_{\text{in}}$. This is possible because GPUs are equipped with Direct Memory Access (DMA) engines, which can handle memory transfers concurrently with compute operations. Additionally, LLM generation is typically a compute-intensive task that involves minimal data transfer between the device and host memory. Hence, lookahead retrieval can run seamlessly alongside LLM generation.
Second, at retrieval stage, the GPU performs a fast similarity search on the accurately predicted clusters ($C_{\text{overlap}}$) after determining $C_{\text{out}}$. Since similarity search essentially involves dot-product computations with high dimension embeddings, this is highly efficient on the GPU.
Third, simultaneously to the second step, the CPU processes the similarity search for remaining clusters of $C_{\text{out}}$ that were not prefetched and is stored on CPU memory. Although CPUs are generally slower at performing similarity search, lookahead retrieval significantly reduces the workload assigned to the CPU, thereby accelerating the overall retrieval process.
Additionally, TeleRAG optimizes the sorting of vector distances, which is required for the $k$-argmin operation, the final step of the IVF search where the top-$k$ indices are identified. Since each cluster could have more than thousands of embedding vectors, sorting is much faster on the GPU than on the CPU [9]. In order to perform the sorting on GPU, TeleRAG passes the CPU computed distances of $C_{\text{miss}}$ to GPU first and then perform the global sorting on the combined distances of $C_{\text{overlap}}$ and $C_{\text{miss}}$. Unlike the original vector data, the distance values (such as L2-norm or inner product [46]) are scalars per data point and thus, are much smaller in size. Hence, transferring them to the GPU incurs minimal overhead, allowing the GPU to achieve end-to-end speedup on the sorting process. This GPU-based sorting optimization is typically absent in CPU-based retrieval implementations [25] as they assume there is no GPU in the systems.
In summary, TeleRAG optimizes retrieval by leveraging lookahead retrieval to hide the data movement time for prefetching, GPU computation for efficient similarity search and sorting, and CPU processing for non-overlapping clusters, significantly accelerating the retrieval process.
4.2 Prefetching Design
Prefetch target. In the early design of TeleRAG, we use number of clusters as the prefetch target as it aligns with the nprobe setting of the IVF retrieval process. With this target, Despite we can configure the number clusters to prefetch to adapt to each pipeline’s characteristics, we found this target brings an unstable inference performance. It’s because the size of each cluster is often heavily skewed [59, 44], making data loading time for some clusters much longer than the others. To solve this, we use number of bytes, $b_{p}$, as the prefetch target. With $b_{p}$ as the prefeth target, it provides an upper bound on prefetching time at $\frac{b_{p}}{B}$, where $B$ stands for the CPU-GPU memory bandwidth.
To perform prefetching, we first sort the clusters by the distance between their centroids and the query embedding vector. Starting with the nearest cluster, we include clusters one by one until adding another would exceed the target $b_{p}$. We do not split clusters to completely fill $b_{p}$, ensuring a clear separation between GPU and CPU searches.
Prefetch for Multi-Round RAG. As discussed in §2.1, RAG pipelines often run through multiple rounds. Because each round still processes for the same input query, the IVF clusters used in retrieval are highly similar across rounds. In TeleRAG, we therefore perform a full prefetch only for the first round, then fetch only the additional IVF clusters needed for subsequent rounds, until reaching the memory budget. We currently flush the prefetched data for each new query, assuming no correlation across queries. We leave more advanced memory management and identification of hot clusters between queries for future work.
4.3 Finding the Optimal Prefetch Amount
A key challenge in TeleRAG is to balance the benefit of reducing retrieval latency by prefetching data against the overhead of CPU-GPU transfers. Prefetching more clusters reduces the subsequent retrieval time but can also extend the transfer phase; if it extends beyond the LLM generation window, we lose the advantage of overlap and potentially introduce additional delay. Nevertheless, even additional delay in transferring data can be worthwhile if it substantially reduces retrieval latency.
Mathematical model. We develop a mathematical model to guide the choice of the optimal amount of data to prefetch. Here, we follow the notations from §4.2 where $b_{p}$ is bytes to prefetch and $B$ is the CPU-GPU bandwidth. Optimal amount of data to prefetch is denoted as $b_{p}^{*}$.
To start, we let $t_{1}$ represent the combined time of prefetching and pre-retrieval LLM generation, and $t_{2}$ represents the retrieval time. We have:
$t_{1}$ $=\max(t_{\text{LLM}},t_{p}),$ (1)
$t_{2}$ $=\max(t_{c},t_{g}),$ (2)
where $t_{\text{LLM}}$ is the time for LLM generation, $t_{p}$ is prefetching time, $t_{c}$ is the CPU retrieval time, and $t_{g}$ is the GPU retrieval time. The objective is to minimize $t_{1}+t_{2}$.
Since prefetching time $t_{p}$ is proportional to $b_{p}$, we express $t_{1}$ as a piecewise function:
$t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$ (3)
From Eq. 3, if we prefetch fewer bytes than can be transferred during LLM generation, $t_{1}$ is effectively just $t_{\text{LLM}}$ because the transfer overlaps completely with LLM execution.
As shown in §3.1, GPU retrieval ($t_{g}$) is generally much faster than CPU retrieval ($t_{c}$). Thus, we assume $t_{c}\gg t_{g}$. As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed [40], such that:
$t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$ (4)
where $r_{\mathrm{miss}}$ is the miss rate (percentage of IVF clusters are not caught on GPU), $n_{\mathrm{probe}}$ is the total number of clusters to search, and $t_{cc}$ is the CPU time to search a single cluster. Increasing $b_{p}$ can only decrease or maintain the current miss rate, i.e., $\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\leq 0$. Moreover, because clusters are prefetched in order of descending likelihood, we assume $r_{\mathrm{miss}}$ is either linear or a concave up function [^2] An upward U-shaped function whose second derivative is positive. of $b_{p}$, i.e., $\frac{\mathrm{d}^{2}r_{\mathrm{miss}}}{\mathrm{d}b_{p}^{2}}\geq 0$.
We now examine two cases:
  • Case 1: $b_{p}^{*}\leq B\cdot t_{\text{LLM}}$. Here, $t_{1}=t_{\text{LLM}}$ is constant because prefetching is fully overlapped with LLM generation. Since increasing $b_{p}$ in this regime will not increase $t_{1}$ and cannot worsen the miss rate, pushing $b_{p}$ to the boundary $B\cdot t_{\text{LLM}}$ minimizes $t_{1}+t_{2}$. Hence,
    $b_{p}^{*}=B\cdot t_{\text{LLM}}.$ (5)
  • Case 2: $b_{p}^{*}>B\cdot t_{\text{LLM}}$. In this region, $t_{1}$ grows linearly with $b_{p}$, and we have: $\frac{\mathrm{d}^{2}}{\mathrm{d}b_{p}^{2}}(t_{1}+t_{2})=\frac{\mathrm{d}^{2}r_{\mathrm{miss}}}{\mathrm{d}b_{p}^{2}}\cdot n_{\mathrm{probe}}\cdot t_{cc}\geq 0.$ Therefore, $t_{1}+t_{2}$ is concave up, allowing at most one minimum. At the minimum point, we have:
    $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$ (6)
    From this, we obtain:
    $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$ (7)
    where $\Delta r_{\mathrm{miss}}$ is the decrement of the miss rate for this round. If $b_{p}^{*}$ is indeed larger than $B\cdot t_{\text{LLM}}$, it becomes the global minimum; otherwise, the solution reverts to Case 1.
In summary, our analysis shows that the optimal prefetch amount $b_{p}^{*}$ can only lie at one of two points: (1) Prefetch until LLM generation completes (i.e. $b_{p}=B\cdot t_{\text{LLM}}$). (2) A point determined by Eq. 7. However, under typical CPU-GPU bandwidth (e.g., 55 GB/s on PCIe 5), the time spent loading additional clusters often outweighs any retrieval latency reduction from lowering $r_{\mathrm{miss}}$. Consequently, the second scenario in Eq. 7 becomes nearly infeasible in practice. Therefore, on current hardware, prefetching exactly until LLM execution ends is generally the most effective choice.
Profiling-guided approach. Although $b_{p}^{*}$ can be derived from the analysis above, it depends on knowing $t_{\mathrm{LLM}}$ for each query, which cannot be obtained ahead of time. To address this, we use a profiling-guided approach, leveraging the observation that, despite differences in query content, the output length (and thus generation time) for a RAG pipeline often remains similar across most queries. Accordingly, for each RAG pipeline, we can measure $t_{\mathrm{LLM}}$ on a calibration set containing n queries, and get estimated $\hat{b}_{p}^{*}=B\cdot\overline{t}_{\mathrm{LLM}}$, where $\overline{t}_{\mathrm{LLM}}=\mathrm{mean}\{t_{\mathrm{LLM},1},t_{\mathrm{LLM},2},\dots,t_{\mathrm{LLM},n}\}.$ This estimated $\hat{b}_{p}^{*}$ can then be used for incoming queries, ensuring a near-optimal prefetch amount.
4.4 Implementation
TeleRAG is implemented in Python and leverages PyTorch’s [8] operator ecosystem for efficient computation. The datastore index is initially constructed using FAISS [25], and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors. These tensors are stored on disk to eliminate the need for repeated conversions.
At runtime, cluster data is loaded into a contiguous pinned memory region on the CPU, enabling non-blocking memory copies to the GPU. A fixed-size contiguous buffer on the GPU is allocated based on the user’s configuration or GPU memory capacity during runtime.
To enable concurrent CPU-GPU data transferring and LLM generation, we utilize PyTorch’s _copy(non_blocking=True) API and use separate CUDA streams for prefetching and LLM. In this way, the data copy operations will not block the GPU to perform computation, and thus the TeleRAG can do prefetching in parallel to the pre-retrieval LLM generation.
To implement the index search with GPU-CPU cooperation, for the GPU part, we use a single matrix-vector multiplication that computes distances for all prefetched vectors; for the CPU part, we utilize multithreading in Python to parallelize similarity searches across clusters. We then move the distances computed from CPU to GPU, merge with distances on GPU and perform global sorting on GPU.
5 Evaluation
Specification Value
Dataset wiki_dpr [48]
Dataset size 2.1 billion tokens
##\## of chunks 21 million
##\## of IVF cluster 4096
Embed model Contriever [35]
Embed dimension 768
Index type FLAT [^3] Original embedding without compression for the best retrieval precision.
Distance metric Inner Product
Index size 61GB
Table 2: Detailed configurations of our retrieval index.
We conduct extensive experiments to evaluate the effectiveness of TeleRAG. In this section, we describe the necessary details on how we set up the evaluations, present experiments results and provide in-depth analysis and discussions.
5.1 Datasets and RAG Pipelines
Datastore. We build a datastore based on the wiki_dpr dataset [48], which is a popular dataset contains 2.1 billions tokens from Wikipedia. Following previous works [11, 48, 64], we chunk the passages by every 100 tokens, and use Contriever [35] to generate embedding for each chunk. The embeddings have a hidden dimension of 768.
Vector index. For the baseline retrieval, we build an IVF vector index using FAISS [25] from the datastore, following setups from  [11]. See Table 2 for the detailed configurations of our vector index. As described in §4.4, we then convert the FAISS index to a customized index in PyTorch for TeleRAG.
LLMs. We evaluate TeleRAG on the Llama model family [82] in three different sizes (Llama-3.2-3B, Llama-3-8B, Llama-2-13B) to represent different use cases.
RAG pipelines. We evaluate TeleRAG with six popular RAG pipelines. Figure 7 shows the overview of each pipeline. Note that even though some pipelines lack pre-retrieval generation, the post-retrieval generation serves similar functionality for the retrieval of the next iteration. Below are brief descriptions to evaluated RAG pipelines. For iterative-based pipelines, we set the maximum iterations to 3.
Figure 7: Overview of six RAG pipelines that we evaluate.
  1. 1.
    HyDE[27] prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
  2. 2.
    SubQestion (SubQ) [^4] Implemented in LlamaIndex as SubQuestionQueryEngine[58] prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
  3. 3.
    Iterative (Iter) [^5] Implemented in LlamaIndex as MultiStepQueryEngine[58] prompts LLM to generate narrower questions first and iteratively refine them based on previous answers. At the end of each iteration, it prompts LLM to judge if the answer is good enough.
  4. 4.
    Iter-RetGen (IRG) [78] iteratively do retrieval and LLM generation for 3 iterations.
  5. 5.
    FLARE [43] iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
  6. 6.
    Self-RAG (S-RAG) [^6] We evaluate on the short-form version which only has one iteration. [11] decides whether to retrieve documents and then answers for each query. A critique stage selects the most confident answer. We use fine-tuned models based on Llama2-7B and Llama2-13B from their official implementation [12].
Evaluation Datasets. We evaluate on three commonly used question-answering datasets, NQ [54], HotpotQA [91] and TriviaQA [47]. For each dataset, we randomly sample 512 queries and report the average unless otherwise specified.
5.2 Experiment Setups
Hardware setups. We evaluate TeleRAG on two hardware environments, Desktop and Server, which equiped with represent the settings for the desktop and cloud instance use case. For Desktop, it’s paired with a RTX 4090 GPU (24GB memory) and we consider the 3B and 8B models. For Server, it’s paired with a RTX 4090 GPU (24GB memory) and we consider the 8B and 13B models. These sizes are common model size for RAG applications [11, 10]. We do not evaluate the 13B model on Desktop as its model size (26 GB) exceeds RTX4090’s memory space. Table 3 summaries the hardware configurations.
Nprobe and top-$k$. We evaluate latency with three nprobe values, 128, 256, and 512, which represent 3.1%, 6.3%, and 12.5% of the total clusters in our evaluated index. This range falls into the commonly used settings of IVF index [106, 25]. We use a top-$k$ number of 3 for retrieval, which returns 3 most relevant passages based on the retrieval query.
RAG pipeline implementation. To collect the input and output texts for each stage and have consistent results across all pipelines, we implement the RAG pipelines with the FlashRAG framework [45]. For IRG, FLARE, and S-RAG, we use the framework’s default implementations; for the other pipelines, we re-implement them using FlashRAG’s APIs.
Benchmark methodology. We follow a benchmarking methodology from  [101] and use GPT-3.5-Turbo [69] to run through each pipeline and record the input and output text of each step. During latency evaluation, we set LLM’s output length based on the recorded text. This way, we ensure a fair latency comparison across different LLM models.
Baseline systems. To evaluate the latency of each pipeline, we construct a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline. For LLM generation, we use SGLang [101], which is a state-of-the-art LLM inference engine.
Prefetching budget setups. Based on the methodology we described in §4.3, we profile each RAG pipeline with 64 random samples from NQ [54] and define the prefetch budget of each pipeline.
Max prefetching memory. We set a maximum GPU memory limit for prefetching in each configuration. For Server with H100, we allocate up to 16GB, sufficient for a high cluster hit rate at the largest tested nprobe value (512). For Desktop with RTX4090, we allocate up to 10GB and 3.75GB for the 3B and 8B models, respectively. These settings demonstrate that TeleRAG can efficiently operate using only a small fraction (up to 40% for RTX4090 and 20% for H100) of total GPU memory.
Setup Desktop Server
CPU TR 5975 EPYC 9554
CPU memory size 512GB 1.5TB
GPU RTX4090 [67] H100 [68]
GPU memory size 24GB 80GB
CPU-GPU Bus PCIe 4 PCIe 5
Bandwidth n 32 (24) GB/s n n 64 (51) GB/s n
Table 3: Hardware specifications for our setups. In bandwidth, the number in the parentheses is the actual bandwidth we measured from our system.
(a) End-to-end latency speedup with Llama-3.2-3B with a single RTX4090 GPU.
(b) End-to-end latency speedup with Llama-3-8B with a single RTX4090 GPU.
(c) End-to-end latency speedup with Llama-3-8B with a single H100 GPU.
(d) End-to-end latency speedup with Llama-2-13B with a single H100 GPU.
Figure 8: End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines, three datasets, with H100 and RTX4090 GPU.
Metric Hardware nprobe
128 256 512
Precision - 95.9% 97.6% 98.9%
CPU Latency (ms) Desktop 89 176 349
Server 64 126 251
Table 4: Retrieval precision and baseline latency of FAISS for different nprobe values. Precision is measured by comparing top-10 results against greedy search without IVF.
5.3 Latency Evaluation
Baseline retrieval. Table 4 reports the retrieval precision and the baseline CPU retrieval latency of different nprobe. To conduct this test, we use 1024 random queries from NQ. For retrieval precision, we compare against the top-10 results from the greedy search. For the latency, we report the average retrieval time. Due to the limited scalability of the CPU, a higher nprobe directly leads to a longer retrieval latency even when using the state-of-the-art implementation from FAISS.
End-to-end latency. We evaluate the end-to-end latency of all the RAG pipelines and datasets under all the hardware settings described in §5.2. Our comprehensive end-to-end speedup results are presented in Figure 8.
We can observe that TeleRAG improves the execution speed consistently with various configurations. Across different pipelines, datasets, and model/hardware setups, the average speedup is 1.21x, 1.41x, and 1.72x for nprobe values of 128, 256, and 512, respectively. We observe a more significant speedup for larger nprobe, as increased nprobe settings incur longer retrieval latencies.
In the best case, we observe around 2.6x speedup (for Iter-RetGen, HotpotQA, Llama-3.2-3B, on RTX4090). This is because the Iter-RetGen pipeline employs a retrieval step in every iteration and has a short LLM output in general, providing a larger end-to-end speedup margin for accelerating the retrieval part.
Another pipeline that has a particularly high speedup is the SubQuestion pipeline, for which TeleRAG achieves around 2.5x speedup across all the datasets with Llama-3.2-3B on RTX4090. This pipeline uses LLM to generate multiple sub-questions and perform a batched retrieval with 3 to 4 generated queries at one time. The CPU retrieval performance of batched retrieval is significantly worse than that of GPU, as it has limited parallelism. In TeleRAG, we handle the batched retrieval efficiently by leveraging GPU’s high parallel computing capacity and, therefore, achieve a high speedup for the SubQuestion pipeline.
Overall, the performance gain is stable across different datasets, with less than 3% of variance across datasets on average. For a given GPU, TeleRAG has a higher speedup with a smaller LLM because retrieval time will account for a larger fraction of the end-to-end latency. Still, we have tangible speedup in RTX4090 with Llama-3-8B model (e.g., more than 1.5x when nprobe is 512, on average), even though prefetching memory budget is only 3.75GB (space left after holding LLM, embedding model, IVF cluster centroids and misc tensors on GPU) in this case.
Speedups on retrieval. Figure 9 shows the latency speedups on retrieval on NQ. We observe a consistent speedups for all the nprobe numbers but notice the greatest speedups of retrieval often occur at nprobe 256, with average speedups of 7.21x and 7.41x on RTX4090 and H100. This is because, with a larger nprobe, the retrieval performance of TeleRAG will start to be bounded by the missed clusters on the CPU (we have a fixed prefetch budget of different nprobe numbers). Nonetheless, TeleRAG still has a large latency improvement against the pure CPU baseline.
(a) On Desktop with a RTX4090 GPU.
(b) On Server with a H100 GPU.
Figure 9: Retrieval speedup on NQ.
(a) With Llama-3-8B.
(b) With Llama-2-13B.
Figure 10: Latency breakdown for NQ on Server with a H100 GPU.
Pipeline H100 (Llm3-8B) 4090 (Llm3-3B)
Budget Hit Rate Budget Hit Rate
HyDE 9 GB $77.23\%$ 7 GB $67.16\%$
SubQ 8 GB $63.26\%$ 7 GB $58.38\%$
Iter 5 GB $80.79\%$ 3 GB $61.46\%$
IRG 4 GB $44.49\%$ 2.5 GB $29.39\%$
FLARE 6 GB $58.17\%$ 3 GB $33.01\%$
S-RAG 3 GB $36.01\%$ 1.25 GB $14.98\%$
Table 5: The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 512.
5.4 Analysis
Latency breakdown and GPU compute interference. We further show the latency breakdown of running RAG pipelines with Llama-3-8B and Llama-2-13B on Server with H100, examining the case when the GPU’s memory capacity is enough for a good amount of prefetching. Figure 10 shows the breakdown of latency into LLM time and retrieval time. We compare the baseline’s retrieval latency (FAISS CPU) with TeleRAG’s retrieval latency in three different nprobe numbers.
The first thing we observe is that, although we set the prefetching budget to the amount that the data loading time can overlap with the pre-retrieval generation time, there is a slight slowdown in LLM generation time for TeleRAG. This is potentially because GPU’s compute performance will be impacted when there is a memory-loading workload, as noticed in  [103]. However, such impact is minimal, with less than 5% slowdown in general. We observe higher impacts for Iter-RetGen and Self-RAG with Llama-3-8B, which TeleRAG’s LLM geeration time is 28% and 20% longer than the baseline. It’s because these two pipelines have a relatively shorter LLM generation time, and thus, are more vulnerable to interference on GPU compute. However, with the significant amount of speedups on retrieval, TeleRAG achieves consistent speedups across all the pipelines despite having this small impact on the LLM generation.
Prefetch budgets and cluster hit rates. Table 5 shows the prefetch budgets we set with the profiling-guided approach on RTX4090 and H100 for NQ. It also presents the averaged cluster hit rate achieved with this prefetch budget. From the table, we can see that it can generally attain a high cluster hit rate ($>50$%) for the cases where it can hold a larger amount of data for prefetching. For pipelines or scenarios that can hold for less than 3GB prefetching amount, we generally observe a relatively lower hit rate ($<35$%), limiting our benefits of reducing CPU’s search workloads. However, as observed from Figure 8, TeleRAG achieves from 1.2x to 1.5x end-to-end speedups for these pipelines. This is the combined benefit of reducing CPU workload and utilizing GPU to perform sorting on similarity distances, demonstrating the strong performance of TeleRAG across a wide range of scenarios.
Analysis on the cluster hit rate. We measure cluster hit rates across a wide range of prefetch amounts, examining scenarios with higher CPU-GPU bandwidth and sufficient GPU memory to accommodate larger prefetches. Our experiments are conducted on the HyDE pipeline with nprobe set to 512 and prefetching from 32 to 1024 clusters. We use nprobe as the prefetch metric (instead of bytes) to facilitate comparisons with optimal hit rates. We evaluate on NQ, HotpotQA, and TriviaQA, reporting average results in Figure 11.
We have three observations from Figure 11. (1) From $0\%$ to $50\%$ of nprobe, accuracy remains high, illustrating that the probability of hitting relevant clusters decreases by order of prefetching. (2) Between $50\%$ and $150\%$ of nprobe, coverage still increases rapidly but is slower than the first range. (3) After $150\%$ of nprobe, coverage plateaus at roughly $95\%$. These results suggest that prefetching the amount beyond the target nprobe is likely unnecessary, as the coverage increase is limited.
Figure 11: The cluster coverage rate for HyDE with different numbers of prefetched clusters. The target retrieval nprobe is $512$. Results are the average of NQ, HotpotQA and TriviaQA.
6 Related Work
6.1 Systems for RAG
Several prior works have attempted to build systems for RAG. RAGCache [44] proposes a caching system that stores KV caches from datastore chunks in an order-aware manner to improve the time to the first token. However, this approach only reduces prefill latency, leaving retrieval and decode times unaffected, despite the fact they typically dominate total latency [4]. Moreover, it assumes repeated use of the same document across multiple requests, limiting its scalability over large data stores. Similarly, TurboRAG [60] precomputes KV caches from the data store, but it also only optimizes prefill latency.
RaLMSpec [98] proposes speculative retrieval and batched verification, Chameleon [41] proposes a CPU-GPU-FPGA heterogeneous architecture for accelerating the retrieval process, and PipeRAG [42] proposes an algorithm-system co-design technique to overlap the retrieval and generation process by modifying the algorithms. However, these works focus on the paradigm of RAG that retrieves documents once every few tokens and are not applicable to modular RAG applications that are widely used now, as discussed in §2.1.
Speculative RAG [87] introduces drafting by smaller LLMs to reduce RAG latency, but it does not target the retrieval latency. APIServe (InferCept) [3] proposes a novel KV cache management strategy that can support the interception of LLM generation by other workloads, including retrieval. However, this work again does not focus on optimizing retrieval latency.
Unlike all these prior works, we tackle system challenges of long retrieval latency and large memory requirement for modular RAG pipelines with large-scale datastores.
6.2 Systems for Compound LLM Applications
Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities [95, 14, 85]. LLMCompiler [52] is a framework that optimizes the execution of multiple functions in large language models by enabling parallel function calling. AI Metropolis [90] accelerates LLM-based multi-agent simulations with out-of-order execution. ALTO [76] improves the performance of complex AI systems by streaming partial results. Some open-source libraries provide graph-based representation for programming this kind of application [22, 56]. RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
6.3 Vector Index
Vector index is gaining growing attention as an essential component of RAG [16, 29], and many works have been proposed to improve the efficiency with system or algorithm techniques. ScaNN proposes an anisotropic quantization method for better efficiency [32]. DiskANN and SPANN propose memory-disk hybrid indexing systems that work beyond the limitation of memory capacity [19, 38]. Other prior works propose hardware acceleration of vector index with GPUs [46], FPGAs [96, 40], TPU [20], or ray tracing hardware [104, 59]. BANG proposes a method to scale graph-based ANN beyond GPU memory [49], and Rummy allows the index to scale beyond GPU memory capacity with reordered pipelining [99]. However, these methods either require algorithm modifications or are bottlenecked by CPU-GPU bandwidth. Our proposal, focuses on the context of modular RAG applications, where queries for retrieval are usually generated by LLMs, and optimizes the latency and the GPU memory consumption without altering the algorithm of the IVF index.
7 Conclusion
In this paper, we introduced TeleRAG, an inference system that tackles the system challenges of RAG pipelins under latency-sensitive scenarios. By using lookahead retrieval, which overlaps data transfer and GPU computation for faster retrieval, a profile-guided approach to determine optimal prefetching data amount, and GPU-CPU cooperation, TeleRAG speeds up the end-to-end latency with minimal GPU memory requirement. Our evaluation shows that TeleRAG significantly improves performance compared to existing state-of-the-art solutions.

### v2

@@
 TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval
-Abstract
-Retrieval-augmented generation (RAG) extends large language models (LLMs) with external data sources to enhance factual correctness and domain coverage.
-Modern RAG pipelines rely on large datastores, leading to system challenges in latency-sensitive deployments, especially when limited GPU memory is available.
-To address these challenges, we propose TeleRAG, an efficient inference system that reduces RAG latency with minimal GPU memory requirements.
-The core innovation of TeleRAG is lookahead retrieval, a prefetching mechanism that anticipates required data and transfers it from CPU to GPU in parallel with LLM generation.
-By leveraging the modularity of RAG pipelines, the inverted file index (IVF) search algorithm and similarities between queries, TeleRAG optimally overlaps data movement and computation.
-Experimental results show that TeleRAG reduces end-to-end RAG inference latency by up to 1.72× on average compared to state-of-the-art systems, enabling faster, more memory-efficient deployments of advanced RAG applications.
-Equal contribution. Work done during internship at UW. Correspondence to: Chien-Yu Lin ¡cyulin@cs.washington.edu¿.
-1 Introduction
+Abstract.
+[…]
+RAG inference, Acceleration, Prefetching.
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: submissionid: 164
+1. Introduction
 (a) Typical pipeline stages of a RAG application.
-(b) Different scenarios for RAG and the illustration of the proposed lookahead retrieval mechanism.
-It prefetches relevant data for retrieval from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU-CPU cooperation.
-Figure 1: (a) Illustrations of RAG pipeline stages. (b) Overview of TeleRAG and comparison to baseline systems.
-Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases[29, 13, 57].
-During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations [63, 74, 50] and incorporate up-to-date or private information [36, 64].
+(b) Illustration of TeleRAG’s lookahead retrieval mechanism and comparison to different RAG systems.
+TeleRAG prefetches relevant retrieval data from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU-CPU cooperation.
+Figure 1. Overview of RAG pipeline stages and comparison of TeleRAG to baseline systems.
+Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020).
+During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and incorporate up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
 Modern RAG applications share two key characteristics.
 (1) RAG applications are built as modular pipelines as shown in Figure 1(a), where a single query undergoes multiple rounds of LLM calls and retrievals, each of which serves different functions to improve overall output quality.
-For example, query transformation rounds [27, 100, 71, 10] generally occur before retrieval to refine the user’s query with LLMs (pre-retrieval generation).
-(2) RAG’s datastores are typically large, supported by recent works demonstrating that increasing the size of the datastore positively affects the performance of RAG applications [64, 50, 15, 33, 77].
-These characteristics create significant challenges for efficient RAG inference, especially in latency-sensitive applications such as customer chatbots [18, 7, 83], financial analysis [62, 65], and emergency medical diagnosis [53, 31].
+For example, query transformation rounds (Gao et al., 2023a; Zheng et al., 2023; Peng et al., 2023; Asai et al., 2024a) generally occur before retrieval to refine the user’s query with LLMs (pre-retrieval generation).
+(2) RAG’s datastores are typically large, supported by recent works demonstrating that larger datastore consistently improves the performance of RAG applications (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n.
+d.]).
+These characteristics create several significant challenges for efficient RAG inference, especially in latency-sensitive applications such as customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024), financial analysis (Malec, 2024; MyScale, 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024).
 First, the latency of RAG systems is inherently increased due to multiple rounds of LLM generation and retrieval.
-Second, although GPU can speed up both LLM and retrieval, the combined memory demand of these processes often exceeds available GPU resources, making it expensive or infeasible to fully host both operations on GPU.
-Consequently, in local or small-scale deployments, which are common for RAG applications handling private or sensitive data, the retrieval datastore is frequently off-loaded to CPU memory to alleviate GPU constraints.
-However, while CPU off-loading resolves memory limitations, it significantly increases retrieval latency, diminishing overall system efficiency.
-Addressing the latency challenge, several recent works have been proposed to accelerate RAG inference.
-These approaches include FPGA-based retrieval acceleration [41], reuse of KV-cache [44, 60, 92], and speculative document retrieval [87, 98].
-However, none of these works directly address the substantial memory demands associated with retrieval, which remains a significant bottleneck for efficient RAG deployments.
-In this paper, we identify a key opportunity in the widely adopted inverted file index (IVF) [80] retrieval method.
-IVF reduces retrieval latency by pre-clustering the datastore into sub-clusters, limiting the search space to relevant clusters of the retrieval query during runtime.
-This approach also offers a way to reduces GPU memory usage, as only a subset of clusters are needed for each retrieval query.
-For example, after identifying relevant clusters, their data can be dynamically transferred from CPU to GPU memory and accelerate the similarity searches on GPU.
-However, this approach incurs high CPU-GPU data transfer overhead, keeping the overall retrieval latency still high [49] (see §3).
+Second, although GPUs can speed up both LLM generation and retrieval, the combination of these processes poses significant memory requirements, often exceeding available GPU resources (see §3.1).
+Hence, it is expensive or even infeasible to fully host both operations on GPUs.
+Consequently, in local or custom deployments, which are common for RAG applications with private or sensitive data, the retrieval datastore is often offloaded to CPU memory to alleviate GPU memory constraints.
+However, while CPU offloading resolves memory limitations, it significantly increases retrieval latency (see §3.1), diminishing overall system efficiency.
+This memory pressure is not limited to local deployments; even for serving scenarios in data centers, where many concurrent requests are processed together and GPUs are plentiful, managing large index sizes still presents a significant challenge for achieving high throughput with latency service level objectives (SLOs).
+Allocating additional GPU memory to the retrieval data store directly reduces the memory available for the KV cache of the LLM serving system, limiting the effective batch sizes (Kwon et al., 2023).
+In contrast, CPU retrieval would increase the per-request latency by a lot, highlighting the need for techniques that reduce GPU memory consumption while preserving low latency.
+To address the latency challenge, several recent works have been proposed to accelerate RAG inference.
+These approaches include hardware accelerators for retrieval (Jiang et al., 2023d; Quinn et al., 2025), reuse of KV-cache (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024), and speculative document retrieval (Wang et al., 2024b; Zhang et al., 2024b).
+However, these works do not address the substantial memory demands associated with retrieval.
+Therefore, designing a memory-efficient system for low-latency RAG deployments is crucial.
+In this paper, we argue that a promising direction to reduce GPU memory usage while maintaining low latency lies in improving the design of the retrieval index.
+Specifically, the inverted file index (IVF) (Sivic and Zisserman, 2003) reduces retrieval latency by pre-clustering data and limiting runtime search to relevant clusters.
+It offers a way to leverage GPU-accelerated search while not increasing memory consumption by dynamically transferring only the necessary clusters from the CPU to the GPU.
+However, this data transfer overhead is a key bottleneck, and needs to be minimized (Khan et al., 2024) (see §3.2).
 Our proposal.
-To tackle latency and memory bottlenecks in RAG inference, we introduce a novel mechanism called lookahead retrieval.
-Lookahead retrieval predicts which subsets of IVF clusters will likely be accessed during the retrieval phase and proactively transfers them from CPU to GPU during the pre-retrieval generation stage.
-The idea behind lookahead retrieval is based on the observation that queries before and after the pre-retrieval generation stage share substantial semantic overlap.
-Consequently, IVF clusters relevant to the initial input query are also likely to be relevant to the refined query.
-To validate this intuition, we rigorously analyze IVF cluster assignments across original queries and refined queries produced by the pre-retrieval generation step, evaluating on six representative RAG pipelines and three datasets.
-Our analysis confirms a strong similarity between these two query forms (see §3).
-Leveraging this insight, we propose TeleRAG, an efficient inference system designed to optimize RAG latency while minimizing GPU memory consumption.
+To tackle latency and memory bottlenecks in RAG inference, we introduce a novel mechanism called lookahead retrieval, which predicts the subsets of IVF clusters that will likely be accessed during runtime and proactively transfers them from the CPU to the GPU before the retrieval stage.
+The key observation behind lookahead retrieval is that there is substantial semantic overlap between queries before and after the pre-retrieval stage.
+Namely, IVF clusters used by the initial input query, which is available well before the retrieval stage, are also likely to be relevant to the actual retrieval query (see §3.3).
+Using this insight, we propose TeleRAG, an efficient inference system designed to optimize RAG latency while minimizing GPU memory consumption.
 TeleRAG employs lookahead retrieval to preemptively load relevant IVF clusters onto the GPU, effectively hiding CPU-GPU data transfer overhead during concurrent LLM generation.
-As illustrated in Figure 1(b), this approach significantly reduces retrieval latency without exceeding GPU memory constraints, enabling efficient execution of RAG applications in resource-limited environments.
-Although TeleRAG significantly reduces latency by processing prefetched IVF clusters on the GPU, some relevant clusters might still be missed during prefetching.
-To maintain retrieval accuracy, TeleRAG concurrently executes similarity searches for these missed clusters on the CPU.
-The results from CPU searches are seamlessly merged with GPU results, ensuring complete retrieval precision.
-Thanks to the high cluster similarity between the initial and refined queries, the number of missed clusters is typically small, allowing TeleRAG to achieve low retrieval latency through efficient CPU-GPU coordination.
-A key challenge in deploying lookahead retrieval is determining the optimal number of IVF clusters to prefetch.
-Prefetching too many clusters increases data-transfer overhead, whereas fetching too few clusters could result in higher retrieval latency due to increased CPU processing.
+As illustrated in Figure 1(b), this approach significantly reduces retrieval latency without exceeding GPU memory constraints, enabling efficient execution of RAG applications.
+To maintain retrieval accuracy while adopting lookahead retrieval, TeleRAG simultaneously searches for clusters that were not prefetched (i.e., mispredicted) from the CPU and merges the results.
+TeleRAG also supports batch and multi-GPU inference.
+A key challenge in deploying lookahead retrieval is determining the optimal number of IVF clusters to prefetch: Prefetching too many clusters increases data-transfer overhead, whereas prefetching too few could result in higher retrieval latency due to increased CPU processing.
 To address this, we propose a profile-guided approach coupled with an analytical model that dynamically determines the ideal prefetch amount based on pipeline characteristics and hardware configurations.
-Our adaptive method consistently outperforms static configurations, delivering substantial latency improvements across diverse RAG setups and hardware environments.
-Results summary.
-We evaluate TeleRAG using a 61GB Wikipedia-based datastore [5] across six popular RAG pipelines built with the Llama model family [82] (3B, 8B, and 13B).
-Remarkably, TeleRAG supports retrieval from a 61GB datastore alongside a Llama3-8B (16GB) LLM on a single RTX 4090 GPU (24GB memory), significantly outperforming the CPU retrieval baseline.
-Experiments conducted on an RTX 4090 and an H100 GPU [68] demonstrate speedups of up to 2.68x and 2.49x, with average speedups of 1.70x and 1.75x, respectively.
-These results highlight the capability of TeleRAG to efficiently handle large-scale RAG inference tasks within tight GPU memory constraints, confirming its practical value for resource-limited deployments.
+[…]
 In summary, we make the following key contributions:
@@
   • Developing TeleRAG, an efficient RAG inference system that integrates lookahead retrieval, resulting in significant acceleration of RAG with minimal GPU memory usage.
-2 Background
-Figure 2: Overview of RAG.
-2.1 RAG
-RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text [29, 13, 57].
-The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations [63, 74, 50] and generate contents based on up-to-date or private information [36, 64].
+2. Background
+Figure 2. Overview of RAG.
+2.1. RAG
+RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020).
+The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and generate contents based on up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
 RAG workflow.
-A basic RAG workflow involves several phases, including data store building, retrieval, and interactions with LLMs [24, 34, 23, 88, 84, 70, 28].
+A basic RAG workflow involves several phases, including data store building, retrieval, and interactions with LLMs (Dixit, 2024; Ilin, 2023; Databricks, 2024; Wei, 2024; Vijeaswaran et al., 2024; Pathway, 2024; Gao, 2024).
 In order to build a data store, raw data in various formats is cleaned, converted to plain text, and divided into chunks.
@@
 When a user provides a query, it is converted into a vector using the same embedding model, and the most similar chunks from the indexed database are retrieved.
-For a large-scale index, approximate algorithms such as the inverted file index (IVF) [80] (see §2.2 for detail) are commonly used to accelerate the search process.
+For a large-scale index, approximate algorithms such as the inverted file index (IVF) (Sivic and Zisserman, 2003) (see §2.2 for detail) are commonly used to accelerate the search process.
 The retrieved chunks, along with the original user query, are combined and given as a prompt to the LLM.
 The LLM then generates a response that relies on the information provided in the retrieved documents.
-This approach enables dynamic and informative responses in both single- and multi-turn dialogues [57].
+This approach enables dynamic and informative responses in both single- and multi-turn dialogues (Lewis et al., 2020).
 Modularity in modern RAG applications.
 However, naively integrating document retrieval into LLM generation can cause several issues.
-For example, the retrieval often struggles to find relevant content and may select irrelevant or mismatched chunks [27, 100, 61], and a single retrieval may not provide sufficient context, necessitating multiple retrieval rounds [29].
-To solve these issues, most state-of-the-art RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  [29, 26, 34, 28, 23, 88].
+For example, the retrieval often struggles to find relevant content and may select irrelevant or mismatched chunks (Gao et al., 2023a; Zheng et al., 2023; Ma et al., 2023), and a single retrieval may not provide sufficient context, necessitating multiple retrieval rounds (Gao et al., 2023b).
+To solve these issues, most state-of-the-art RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  (Gao et al., 2023b; Fan et al., 2024; Ilin, 2023; Gao, 2024; Databricks, 2024; Wei, 2024).
 Typically, they have the following types of steps (shown in Figure 2):
   • Pre-retrieval generation, used to assess whether retrieval is needed or to generate queries for retrieval.
-An example of a pre-retrieval technique is query transformation [61, 37, 27, 102, 93, 73, 72, 100, 58], which reformulates the original query to make it clearer and more suitable for the retrieval task.
+An example of a pre-retrieval technique is query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022), which reformulates the original query to make it clearer and more suitable for the retrieval task.
   • Retrieval, used to identify relevant documents from the vector data store. This stage takes the output of pre-retrieval generation and generates data for the next stage.
   • Post-retrieval generation, generates response based on user query and retrieved documents.
-It can also perform additional process such as summarization [39, 51] or reranking[105, 81] on the retrieved documents.
+It can also perform additional process such as summarization (Jiang et al., 2023b; Kim et al., 2023b) or reranking (Zhuang et al., 2023; Stewart and Linsdell, 2024) on the retrieved documents.
   • Judgment, dynamically determines the execution flow. For example, it decides if more iteration is needed to enhance the response. Heuristics or LLMs can be used for judgement stage.
-By proceeding through these functions, RAG applications can deliver more precise and contextually appropriate responses, significantly enhancing the capabilities of LLMs for various applications [45, 29].
-2.2 Vector Index and Inverted File (IVF)
+By proceeding through these functions, RAG applications can deliver more precise and contextually appropriate responses, significantly enhancing the capabilities of LLMs for various applications (Jin et al., 2024b; Gao et al., 2023b).
+2.2. Vector Index and Inverted File (IVF)
 The vector index is a crucial component of RAG applications that retrieves similar items from large datasets of high-dimensional vectors.
 Given the query vector $x\in\mathrm{R}^{D}$ and vector database $Y=\{y_{0},\ldots,y_{N-1}\}\subset\mathrm{R}^{D}$, which comprises $N$ vectors, the vector index aims to find the $k$ nearest neighbors of $x$ from the database:
-[…]
-where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product [46].
+$\text{k-argmin}_{i=0:N}d(x,y_{i}),$
+where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product (Johnson et al., 2019).
 To improve search efficiency, the inverted file index (IVF) algorithm is widely used for large-scale vector databases due to its simplicity and effectiveness.
 IVF partitions the data store into clusters and restricts searches to the most relevant clusters, reducing the computational cost.
-To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means [66] and partitions the database into $N_{c}$ clusters:
+To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means (Norouzi and Fleet, 2013) and partitions the database into $N_{c}$ clusters:
 $\{C_{1},C_{2},\ldots,C_{N_{c}}\}=k\text{-means}(Y,N_{c}),$
@@
 In this way, IVF reduces search space and accelerates the retrieval process.
-Here, the hyperparameter $L$ from the first step is also referred as nprobe [25].
+Here, the hyperparameter $L$ from the first step is also referred as nprobe (Douze et al., 2024).
 When nprobe is larger, the IVF retrieval will be more accurate as it search more clusters.
@@
 Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs.
-Open-source libraries offer efficient GPU implementations [46, 75].
+Open-source libraries offer efficient GPU implementations (Johnson et al., 2019; Rapidsai, 2022).
 However, IVF does not reduce the memory requirement for the index since the data for all clusters must be stored.
 As a result, IVF still requires a substantial GPU memory footprint, which becomes a bottleneck when GPU memory capacity is the constraint of RAG systems.
-Moreover, recent work reported that increasing the size of the data store positively affects RAG application performance [64, 50, 15, 33, 77], exacerbating this issue.
-2.3 Latency-Sensitive RAG Use Cases
-In this paper, we focus on reducing inference latency for a single query, a critical objective in latency-sensitive RAG applications.
-These applications span a range of real-world scenarios, such as customer chatbots [18, 7, 83], financial analysis [62, 65], autonomous driving [17, 94], and emergency medical diagnosis [53, 31], where users expect near-instantaneous responses.
-In each of these contexts, maintaining low latency per query is paramount for providing timely and accurate results.
-Batching multiple queries, a common strategy for improving throughput, is not considered in this work.
-It is because many RAG applications involve user-specific private data [86, 55, 21, 89, 30] and serve relatively small user bases, making batching infeasible.
-For large-scale RAG services [2, 1, 10], although batching is feasible, it increases the latency for each individual queries [103, 97, 79] and therefore, is not ideal for the latency sensitive applications.
-A straightforward way to reduce inference latency is to scale up GPU resources, allowing both the LLM and retrieval datastore to reside in GPU memory.
-However, this approach is often cost-prohibitive and could lead to significant hardware underutilization, particularly in the single-query scenarios that this work targets.
-Therefore, in this paper, we aim to improve single-query RAG latency without significantly increasing GPU memory usage, ensuring that both cloud-based and local deployments can satisfy stringent performance requirements without incurring excessive costs.
-3 Analyzing Latency of RAG Pipelines
+Moreover, recent work reported that increasing the size of the data store positively affects RAG application performance (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n.
+d.]), exacerbating this issue.
+2.3. Use Cases and Motivation
+In this paper, we aim to achieve low-latency RAG inference without driving up GPU memory consumption, which is relevant for both local and data center settings.
+This goal is critical for efficiently serving a wide range of real-world applications where users expect near-instantaneous responses, thus imposing strict latency SLOs.
+Examples include customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024; AWS, [n.
+d.]), financial analysis (Malec, 2024; MyScale, 2024), autonomous driving (Cai et al., 2024; Yuan et al., 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024).
+On the one hand, there is significant demand for RAG applications that involve user-specific private data in local settings (Wang and Chau, 2024; Lam et al., 2023; Chung et al., 2023; Wutschitz et al., 2023; Ghodratnama and Zakershahrak, 2023).
+In such scenarios, the available GPU resources are usually limited, and the system serves a single request at a time.
+Therefore, CPU offloading is necessary when using a large vector index, which introduces substantial latency as we will show in the next section.
+On the other hand, for large-scale RAG services in data centers (per, [n.
+d.]; gen, [n.
+d.]; Asai et al., 2024a), the goal is to serve many concurrent user requests cost-effectively, necessitating high system throughput.
+However, maximizing LLM serving throughput is often bottlenecked by the GPU memory available for the KV cache, which dictates the maximum possible batch size (Kwon et al., 2023).
+If a large retrieval datastore is loaded onto GPUs, it consumes memory that could otherwise be used for a larger KV cache, thereby limiting throughput.
+Hence, falling back to slow CPU retrieval to save GPU memory is a common option (Jiang et al., 2025), but this is not desirable for latency-sensitive applications due to latency overhead.
+In both cases, a straightforward way to reduce inference latency is scaling up GPU resources to allow both the LLM and datastore to reside in GPU memory, which is usually cost-prohibitive.
+Therefore, in this paper, we aim to design an efficient RAG inference system that can satisfy latency SLOs without requiring a significant amount of GPU memory.
+3. Analyzing Latency of RAG Pipelines
 In this section, we analyze state-of-the-art RAG applications and identify their underlying systems challenges in achieving low latency.
-To conduct the analysis, we construct a 61GB vector index with the FAISS library [25], and set the IVF clusters to 4096 following common practice [11].
-We use Llama-3-8B [6] as the LLM, and run our analysis on an H100 GPU with 80GB memory.
-§5.2 provides details about our experimental setup.
-3.1 End-to-end Latency of RAG Pipelines
-Figure 3: Latency breakdown of six RAG pipelines on NQ dataset [54].
-Figure 4: The breakdown of memory consumption at GPU and CPU for two different strategies: CPU offloading and GPU-based retrieval.
+To conduct the analysis, we construct a 61 GB vector index with the Faiss library (Douze et al., 2024), and set the IVF clusters to 4096 following common practice (Asai et al., 2023).
+We use Llama-3-8B (AI@Meta, 2024) as the LLM, and run our analysis on an RTX 4090 (24 GB memory) and H100 (80 GB memory).
+§5.2 provides more details about our experimental setups.
+3.1. End-to-end Latency of RAG Pipelines
+(a) Batch = 1.
+(b) Batch = 4.
+Figure 3. Latency breakdown of six RAG pipelines on NQ dataset (Kwiatkowski et al., 2019) with one H100 GPU. nprobe is 256.
+Figure 4.
+The breakdown of memory consumption at GPU and CPU for two different strategies: CPU offloading and GPU-based retrieval.
 The dotted line indicates the memory capacity of a RTX4090 GPU, which is a common used GPU for local deployment.
-We begin by analyzing the end-to-end latency of six representative RAG pipelines (see §5.1 for details) in two scenarios.
-In the first, the LLM runs on the GPU while retrieval is offloaded to the CPU, minimizing GPU memory usage.
-In the second, both the LLM and the vector index reside in GPU memory, enabling GPU-based retrieval for lower latency.
-If the GPU’s memory is insufficient, then CPU offloading is necessary.
-We set the nprobe to 512, a common used setting under this index’s configuration (see §5.2 for details), to measure retrieval latency.
-Figure 3 shows the breakdown of end-to-end latency into LLM computation and retrieval.
-The results are the average of 512 random data from the NQ dataset [54].
-We observe that CPU-based retrieval (first scenario) is the primary bottleneck, consuming about 56% of the total execution time on average, while GPU-based retrieval (second scenario) accounts for just 12%.
-On average, GPU retrieval is 12.5x faster than CPU retrieval, reducing overall latency by approximately 2.2x.
-Thus, accelerating retrieval on GPU is crucial for improving end-to-end latency.
+We first analyze the end-to-end latency of six representative RAG pipelines (see §5.1 for details) in two scenarios: (1) the LLM runs on the GPU, while retrieval is offloaded to the CPU, minimizing GPU memory usage, and (2) both the LLM and the vector index reside in GPU memory, enabling GPU-based retrieval for lower latency.
+If the GPU’s memory is insufficient, CPU offloading is necessary.
+We set the nprobe to 256, a commonly used setting under this index’s configuration (see §5.2 for details), to measure retrieval latency.
+Figure 3 shows the breakdown of end-to-end latency into LLM inference part and retrieval part, based on 1024 randomly sampled queries from the NQ dataset (Kwiatkowski et al., 2019).
+We observe that, in the CPU-based retrieval baseline (first scenario), the retrieval phase dominates the latency, consuming 41.1% and 60.5% of total latency for batch sizes 1 and 4, respectively.
+In contrast, GPU-accelerated retrieval in the second scenario dramatically reduces this bottleneck, accounting for only 10.5% and 28.3% of the latency in comparable configurations.
+On average, GPU retrieval is $5.96\times$ (batch size 1) or $3.87\times$ (batch size 4) faster than CPU retrieval, reducing overall latency by approximately $1.5\times$ or $1.8\times$, respectively.
+Thus, accelerating retrieval on the GPU is crucial for improving end-to-end latency.
 However, GPU acceleration comes with a significant memory cost.
@@
 Thus, GPU acceleration on retrieval is often not feasible when running on a lower-end GPU or indexing with a large datastore.
-Note that the memory footprint of the LLM and the index can be even larger and further exacerbate the memory constraint.
-In the rest of this section, we try to answer the following question: Is it possible to achieve the latency of GPU-based retrieval while using much less GPU memory?
-3.2 Opportunities of GPU-accelerated Retrieval with Limited Memory
-A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on-demand for each query, leveraging the IVF index structure that narrows the target search clusters (§2.2).
-While such a method enables faster search on the GPU, the bottleneck of retrieval is shifted to the data fetching.
-To examine the performance, we reconstruct the IVF index in PyTorch [8], leveraging its high-performance GPU operator, and implement such an on-demand fetching retrieval system.
-We measure the retrieval latency with an RTX 4090 GPU and compare it with the baseline of CPU-based retrieval.
-Figure 5 shows the breakdown of retrieval time into data fetch and search time.
-We show the performance for different amounts of data fetched (indicated by the nprobe parameter).
-We observe that the fetch time dominates the latency due to limited PCIe bandwidth between the CPU and GPU (32 GB/s).
-Despite the fact that GPU retrieval is significantly faster than CPU retrieval, due to the fetch overhead, the effective latency is slower than CPU retrieval (around 3% slow down on average across nprobe’s).
-Therefore, we must hide the data fetching latency in order to achieve a meaningful speedup to following such approach.
-Hiding the data fetch cost.
-Hiding the data fetch cost implies that CPU-to-GPU data fetch must be done before the retrieval operation is performed.
-However, predicting which data to retrieve in advance is challenging since the exact data requirements only become clear at retrieval time once we have the query produced by the pre-retrieval process (§2.1).
-Fortunately, as we show next, the original form of a query before pre-retrieval can serve as a valuable hint for the clusters required to be retrieved at the end of the pre-retrieval stage.
-This hint can guide us in predicting the clusters that should be fetched, and the data transfer can be initiated in the background along with pre-retrieval to hide its latency.
-Figure 5: Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
+Even on GPUs capable of holding the entire index in memory (e.g., H100), the index’s memory footprint limits throughput in data center serving scenario.
+Batched inference maximizes throughput, but it is bottlenecked by the KV cache capacity for LLM serving (Kwon et al., 2023).
+For the case of the example above, storing the index on the GPU leaves minimal memory for the KV cache (e.g., ~3 GB, or ~20k tokens worth), restricting batch size, while offloading the index to the CPU significantly increases the available memory (e.g., to ¿500k tokens worth).
+This issue is exacerbated since RAG typically involves long contexts (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024) and there is a trend of growing model and index sizes (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n.
+d.]), increasing memory pressure and necessitating offloading.
+However, CPU offloading is suboptimal for latency-sensitive applications as discussed above.
+Therefore, in the rest of this section, we try to answer the following question: Is it possible to achieve the latency of GPU-based retrieval while using much less GPU memory?
+Figure 5. Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
+3.2. Opportunities of GPU-accelerated Retrieval with Limited Memory
+A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on-demand for each query, leveraging the IVF index structure that narrows the search space (§2.2).
+While this method enables faster searching on the GPU, data fetching becomes the bottleneck.
+In Figure 5, we compare the latency of the on-demand fetching system against CPU retrieval on an RTX 4090 GPU.
+We show three different nprobe values that determine the amount of data fetched.
+Overall, fetch time dominates latency due to the limited PCIe bandwidth between the CPU and GPU (32 GB/s).
+Although the GPU search is substantially faster than the CPU counterpart, the fetch overhead results in a higher end-to-end latency, which is about $3\%$ slower on average across nprobe values.
+Thus, to realize any meaningful speedup with this approach, the data fetching latency must be effectively hidden.
+To hide data fetch costs, CPU-to-GPU transfer must precede retrieval, which requires predicting which data will be accessed.
+Fortunately, as we show next, the modern RAG pipelines offer a valuable hint as a query at the previous step.
 Dataset HyDE SubQ Iter IRG FLARE S-RAG
-[27] [58] [58] [78] [43] [11]
-NQ [54] 75.7% 66.2% 93.4% 85.7% 80.3% 100.0%
-HotpotQA [91] 78.6% 66.6% 86.9% 91.2% 82.2% 100.0%
-TriviaQA [47] 76.8% 65.3% 86.8% 87.5% 82.9% 100.0%
-Table 1: IVF cluster overlapping rate between the input and output of the pre-retrieval generation.
+(Gao et al., 2023a) (Liu, 2022) (Liu, 2022) (Shao et al., 2023) (Jiang et al., 2023c) (Asai et al., 2023)
+NQ (Kwiatkowski et al., 2019) 73.1% 63.2% 91.5% 83.8% 79.1% 100.0%
+HotpotQA (Yang et al., 2018) 75.3% 62.5% 89.6% 89.4% 80.2% 100.0%
+TriviaQA (Joshi et al., 2017) 73.1% 61.6% 86.2% 86.1% 81.7% 100.0%
+Table 1.
+IVF cluster overlapping rate between the input and output of the pre-retrieval generation.
 The nprobe is set to 512.
 Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
-3.3 Overlapping of IVF Clusters
+3.3. Overlapping of IVF Clusters
 While the exact data to be retrieved is only known after the pre-retrieval generation is done, we observe that there are high similarities of the IVF clusters assignments between the queries at different stages.
 Similarity of queries at different stages.
-During the pre-retrieval process of RAG pipelines (e.g., query transformation [61, 37, 27, 102, 93, 73, 72, 100, 58]), an LLM call is issued to refine an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$, which is then used for retrieval.
-This process often rewrites the query into a different format [27] or simplifies its complexities [58], which intuitively preserves the query’s core semantic content.
+During the pre-retrieval process of RAG pipelines (e.g., query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022)), an LLM call is issued to refine an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$, which is then used for retrieval.
+This process often rewrites the query into a different format (Gao et al., 2023a) or simplifies its complexities (Liu, 2022), which intuitively preserves the query’s core semantic content.
 Hence, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar.
@@
 Prediction coverage.
-To verify this hypothesis, we evaluate the average cluster coverage rate between prefetched clusters and clusters used for retrieval in three popular question-answer datasets (NQ [54], HotpotQA [91], and TriviaQA [47]) and six RAG pipelines.
-Table 1 shows the coverage when we prefetch $512$ clusters.
+To verify this hypothesis, we evaluate the average cluster coverage rate between prefetched clusters and clusters used for retrieval in three popular question-answer datasets (NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017)) and six RAG pipelines.
+Table 1 shows the coverage when we prefetch $256$ clusters.
 From the table, we observe that IVF cluster overlap rates are consistently high across a range of datasets and pipelines.
-For instance, even the lowest reported values remain above 65% (for SubQ).
+For instance, even the lowest reported values remain above 61.6% (for SubQ).
 Opportunity.
@@
 In this paper, we aim to leverage this observation to accelerate the inference latency for RAG.
-4 Design of TeleRAG
-Figure 6: (a) The overview of TeleRAG’s lookahead retrieval and comparison to the baseline CPU-offloaded retrieval.
-(b) System design of TeleRAG.
-It first loads the entire index data into CPU memory and identifies the clusters ($C_{\text{in}}$) for the query before the transformation ($q_{\text{in}}$).
-It then transfers the data of $C_{\text{in}}$ to GPU memory while running the pre-retrieval stage and generating the retrieval query $q_{\text{out}}$.
-During the search, the overlapped clusters are searched on GPU, and the missed clusters are searched on CPU.
-The retrieval results are then merged and passed to the next stage.
-Based on the high IVF cluster overlapping observation presented in § 3.3, we present TeleRAG, an efficient inference system for RAG that incorporates the idea of lookahead retrieval, which predicts and prefetches likely IVF cluster data to GPU and significantly accelerates the retrieval process with minimal GPU memory requirement.
-In this section, we describe our design overview (§4.1), an analytical model on finding the optimal prefetch data amount (§4.3), and the implementation details (§4.4).
-4.1 Overview
+Figure 6.
+(a) The overview of lookahead retrieval compared against CPU-offloaded retrieval, and (b) the system design of TeleRAG.
+After identifying clusters ($C_{\text{in}}$) for the initial query $q_{\text{in}}$, $C_{\text{in}}$ is transferred to the GPU while concurrently generating $q_{\text{out}}$.
+At retrieval time, the GPU searches prefetched clusters while the CPU handles remaining ones, before results are merged.
+4. Design of TeleRAG
+Based on the high IVF cluster overlapping observation presented in §3.3, we propose TeleRAG, an efficient RAG inference system that utilizes lookahead retrieval to prefetch a minimal set of likely IVF clusters to GPU, and accelerate the retrieval process.
+In this section, we detail the design of lookahead retrieval and implementation of TeleRAG.
+4.1. Overview
 Figure 6 presents the overview of TeleRAG with lookahead retrieval.
-In Figure 6, the input to the pre-retrieval stage is defined as $q_{\text{in}}$, and the output is denoted as $q_{\text{out}}$.
-In the subsequent retrieval stage, $q_{\text{out}}$ serves as the input query to perform the retrieval.
-We highlight the IVF clusters corresponding to $q_{\text{in}}$ with a red background ($C_{\text{in}}$), and those corresponding to $q_{\text{out}}$ with a blue background ($C_{\text{out}}$).
-Due to the semantic similarity of $q_{\text{in}}$ and $q_{\text{out}}$, there is significant overlap between $C_{\text{in}}$ and $C_{\text{out}}$3.3), marked with a purple background ($C_{\text{overlap}}$).
-Given this, the lookahead retrieval technique in TeleRAG operates in three key steps.
-First, $C_{\text{in}}$ is computed, and their data is transferred to GPU memory during pre-retrieval generation, beginning with those clusters whose centroids are closest to $q_{\text{in}}$.
-This is possible because GPUs are equipped with Direct Memory Access (DMA) engines, which can handle memory transfers concurrently with compute operations.
-Additionally, LLM generation is typically a compute-intensive task that involves minimal data transfer between the device and host memory.
-Hence, lookahead retrieval can run seamlessly alongside LLM generation.
-Second, at retrieval stage, the GPU performs a fast similarity search on the accurately predicted clusters ($C_{\text{overlap}}$) after determining $C_{\text{out}}$.
-Since similarity search essentially involves dot-product computations with high dimension embeddings, this is highly efficient on the GPU.
-Third, simultaneously to the second step, the CPU processes the similarity search for remaining clusters of $C_{\text{out}}$ that were not prefetched and is stored on CPU memory.
-Although CPUs are generally slower at performing similarity search, lookahead retrieval significantly reduces the workload assigned to the CPU, thereby accelerating the overall retrieval process.
-Additionally, TeleRAG optimizes the sorting of vector distances, which is required for the $k$-argmin operation, the final step of the IVF search where the top-$k$ indices are identified.
-Since each cluster could have more than thousands of embedding vectors, sorting is much faster on the GPU than on the CPU [9].
-In order to perform the sorting on GPU, TeleRAG passes the CPU computed distances of $C_{\text{miss}}$ to GPU first and then perform the global sorting on the combined distances of $C_{\text{overlap}}$ and $C_{\text{miss}}$.
-Unlike the original vector data, the distance values (such as L2-norm or inner product [46]) are scalars per data point and thus, are much smaller in size.
-Hence, transferring them to the GPU incurs minimal overhead, allowing the GPU to achieve end-to-end speedup on the sorting process.
-This GPU-based sorting optimization is typically absent in CPU-based retrieval implementations [25] as they assume there is no GPU in the systems.
-In summary, TeleRAG optimizes retrieval by leveraging lookahead retrieval to hide the data movement time for prefetching, GPU computation for efficient similarity search and sorting, and CPU processing for non-overlapping clusters, significantly accelerating the retrieval process.
-4.2 Prefetching Design
-Prefetch target.
-In the early design of TeleRAG, we use number of clusters as the prefetch target as it aligns with the nprobe setting of the IVF retrieval process.
-With this target, Despite we can configure the number clusters to prefetch to adapt to each pipeline’s characteristics, we found this target brings an unstable inference performance.
-It’s because the size of each cluster is often heavily skewed [59, 44], making data loading time for some clusters much longer than the others.
-To solve this, we use number of bytes, $b_{p}$, as the prefetch target.
-With $b_{p}$ as the prefeth target, it provides an upper bound on prefetching time at $\frac{b_{p}}{B}$, where $B$ stands for the CPU-GPU memory bandwidth.
-To perform prefetching, we first sort the clusters by the distance between their centroids and the query embedding vector.
-Starting with the nearest cluster, we include clusters one by one until adding another would exceed the target $b_{p}$.
-We do not split clusters to completely fill $b_{p}$, ensuring a clear separation between GPU and CPU searches.
-Prefetch for Multi-Round RAG.
-As discussed in §2.1, RAG pipelines often run through multiple rounds.
-Because each round still processes for the same input query, the IVF clusters used in retrieval are highly similar across rounds.
-In TeleRAG, we therefore perform a full prefetch only for the first round, then fetch only the additional IVF clusters needed for subsequent rounds, until reaching the memory budget.
-We currently flush the prefetched data for each new query, assuming no correlation across queries.
-We leave more advanced memory management and identification of hot clusters between queries for future work.
-4.3 Finding the Optimal Prefetch Amount
+In Figure 6, $q_{\text{in}}$ denotes the initial input to the pre-retrieval stage and $q_{\text{out}}$ denotes its output, i.e., the query given to the subsequent retrieval stage.
+We highlight the IVF clusters corresponding to $q_{\text{in}}$ and $q_{\text{out}}$ with a red background ($C_{\text{in}}$), and a blue background ($C_{\text{out}}$), respectively.
+Due to the semantic similarity of $q_{\text{in}}$ and $q_{\text{out}}$, there is significant overlap between $C_{\text{in}}$ and $C_{\text{out}}$, marked with a purple background ($C_{\text{overlap}}$).
+Given this, the proposed lookahead retrieval technique in TeleRAG operates in three key steps:
+  1. (1)
+    Pre-retrieval & data transfer: During LLM generation, identify $C_{\text{in}}$ and transfer corresponding data to GPU memory, prioritizing clusters closest to $q_{\text{in}}$.
+This leverages GPU DMA for concurrent transfer of clusters data and compute of LLM generation.
+  2. (2)
+    GPU similarity search: After identifying $C_{\text{out}}$, the GPU efficiently performs a similarity search on the predicted clusters ($C_{\text{overlap}}$) already in its memory.
+  3. (3)
+    CPU similarity search: Concurrently with the GPU search, the CPU processes the similarity search for the missed clusters ($C_{\text{miss}}$) that were not prefetched to the GPU, reducing the CPU’s overall workload.
+In summary, our lookahead retrieval speeds up retrieval by overlapping data prefetching to the GPU with LLM generation and performing concurrent similarity searches on both the GPU (for predicted clusters) and CPU (for the remainder), significantly reducing the CPU’s computational workload.
+4.2. Finding the Optimal Prefetch Amount
 A key challenge in TeleRAG is to balance the benefit of reducing retrieval latency by prefetching data against the overhead of CPU-GPU transfers.
 Prefetching more clusters reduces the subsequent retrieval time but can also extend the transfer phase; if it extends beyond the LLM generation window, we lose the advantage of overlap and potentially introduce additional delay.
-Nevertheless, even additional delay in transferring data can be worthwhile if it substantially reduces retrieval latency.
+Still, additional delays in data transfer may be worthwhile if they substantially reduce retrieval latency.
+To guide the choice of the optimal amount of data to prefetch, we develop a mathematical model based on profile information.
 Mathematical model.
-We develop a mathematical model to guide the choice of the optimal amount of data to prefetch.
-Here, we follow the notations from §4.2 where $b_{p}$ is bytes to prefetch and $B$ is the CPU-GPU bandwidth.
-Optimal amount of data to prefetch is denoted as $b_{p}^{*}$.
-To start, we let $t_{1}$ represent the combined time of prefetching and pre-retrieval LLM generation, and $t_{2}$ represents the retrieval time. We have:
-$t_{1}$ $=\max(t_{\text{LLM}},t_{p}),$ (1)
-$t_{2}$ $=\max(t_{c},t_{g}),$ (2)
-where $t_{\text{LLM}}$ is the time for LLM generation, $t_{p}$ is prefetching time, $t_{c}$ is the CPU retrieval time, and $t_{g}$ is the GPU retrieval time.
+Here, we denote $b_{p}$ as the number of bytes to prefetch and $B$ as the CPU-GPU bandwidth.
+The optimal amount of data to prefetch is denoted as $b_{p}^{*}$.
+To start, we let $t_{1}$ represent the combined time of prefetching and pre-retrieval LLM generation, and $t_{2}$ represents the retrieval time.
+We have: $t_{1}=\max(t_{\text{LLM}},t_{p})$ and $t_{2}=\max(t_{c},t_{g})$, where $t_{\text{LLM}}$ is the time for LLM generation, $t_{p}$ is prefetching time, $t_{c}$ is the CPU retrieval time, and $t_{g}$ is the GPU retrieval time.
 The objective is to minimize $t_{1}+t_{2}$.
 Since prefetching time $t_{p}$ is proportional to $b_{p}$, we express $t_{1}$ as a piecewise function:
-$t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$ (3)
+(1) $t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$
 From Eq. 1, if we prefetch fewer bytes than can be transferred during LLM generation, $t_{1}$ is effectively just $t_{\text{LLM}}$ because the transfer overlaps completely with LLM execution.
@@
 Thus, we assume $t_{c}\gg t_{g}$.
-As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed [40], such that:
-$t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$ (4)
+As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed (Jiang et al., 2023a):
+(2) $t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$
 where $r_{\mathrm{miss}}$ is the miss rate (percentage of IVF clusters are not caught on GPU), $n_{\mathrm{probe}}$ is the total number of clusters to search, and $t_{cc}$ is the CPU time to search a single cluster.
@@
 Hence,
-    $b_{p}^{*}=B\cdot t_{\text{LLM}}.$ (5)
+    (3) $b_{p}^{*}=B\cdot t_{\text{LLM}}.$
   • Case 2: $b_{p}^{*}>B\cdot t_{\text{LLM}}$.
@@
 At the minimum point, we have:
-    $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$ (6)
+    (4) $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$
     From this, we obtain:
-    $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$ (7)
+    (5) $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$
     where $\Delta r_{\mathrm{miss}}$ is the decrement of the miss rate for this round.
@@
 (2) A point determined by Eq.
-7.
+5.
 However, under typical CPU-GPU bandwidth (e.g., 55 GB/s on PCIe 5), the time spent loading additional clusters often outweighs any retrieval latency reduction from lowering $r_{\mathrm{miss}}$.
@@
 Accordingly, for each RAG pipeline, we can measure $t_{\mathrm{LLM}}$ on a calibration set containing n queries, and get estimated $\hat{b}_{p}^{*}=B\cdot\overline{t}_{\mathrm{LLM}}$, where $\overline{t}_{\mathrm{LLM}}=\mathrm{mean}\{t_{\mathrm{LLM},1},t_{\mathrm{LLM},2},\dots,t_{\mathrm{LLM},n}\}.$ This estimated $\hat{b}_{p}^{*}$ can then be used for incoming queries, ensuring a near-optimal prefetch amount.
-4.4 Implementation
-TeleRAG is implemented in Python and leverages PyTorch’s [8] operator ecosystem for efficient computation.
-The datastore index is initially constructed using FAISS [25], and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors.
-These tensors are stored on disk to eliminate the need for repeated conversions.
+4.3. Design Details
+Sorting Optimization.
+TeleRAG accelerates the final $k$-argmin sorting step of IVF search by leveraging GPU.
+Because each cluster may contain thousands of points, sorting is particularly well-suited to GPU parallelism (Arkhipov et al., 2017).
+To enable this, TeleRAG transfers the scalar distance values computed on the CPU for $C_{\text{miss}}$ to the GPU.
+Unlike full vector data, this transfer incurs negligible overhead.
+The GPU then performs a global sort over the combined distances from $C_{\text{overlap}}$ and $C_{\text{miss}}$, delivering end-to-end speedups than the CPU-only retrieval systems.
+Prefetch Target.
+TeleRAG’s prefetching mechanism targets a specific number of bytes ($b_{p}$) rather than the number of clusters to ensure more predictable performance.
+Since IVF cluster sizes are often highly uneven, targeting a fixed cluster count can lead to unstable data loading times, whereas a byte limit provides a clearer upper bound on the transfer duration based on bandwidth.
+When selecting clusters to meet the prefetch byte budget ($b_{p}$), the system adds whole clusters sequentially based on proximity to the query.
+If adding the next closest cluster would exceed the budget, it is omitted entirely rather than being partially loaded, ensuring a clean division where each cluster is processed exclusively by either GPU or CPU.
+Multi-Round Support.
+For multi-round RAG involving the same input query, the system performs a full prefetch (up to the budget) only in the first round, leveraging high cluster similarity across rounds.
+In subsequent rounds, it incrementally fetches only the additional required clusters that were not loaded previously, optimizing data transfer while respecting the memory budget.
+Batch Support.
+For batch inference, a challenge for lookahead retrieval is that each query requires a different set of IVF clusters.
+In our design, we fix the prefetching budget to the amount we got from §4.2 irrespective of batch size, and equally distribute the budget to each query in a batch.
+As a result, although the average cluster hit rate per request will decrease for larger batch sizes, we can ensure each query’s retrieval search time is equally accelerated, forming a balanced batched retrieval.
+Multi-GPU support.
+For multi-GPU systems, we perform a greedy search based on embedding similarities of queries to group similar queries into mini-batches, and distribute the mini-batches among GPUs.
+The greedy search is efficient and incurs negligible overhead: In our profiling for up to the batch size of 256, the latency to perform greedy search is less than $0.1$ s, which is minimal compared to the end-to-end latency of each batch.
+4.4. Implementation Details
+TeleRAG is implemented in Python and also leverages PyTorch’s (Ansel et al., 2024) operator ecosystem for efficient computation.
+The datastore index is initially constructed using FAISS (Douze et al., 2024), and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors.
 At runtime, cluster data is loaded into a contiguous pinned memory region on the CPU, enabling non-blocking memory copies to the GPU.
 A fixed-size contiguous buffer on the GPU is allocated based on the user’s configuration or GPU memory capacity during runtime.
-To enable concurrent CPU-GPU data transferring and LLM generation, we utilize PyTorch’s _copy(non_blocking=True) API and use separate CUDA streams for prefetching and LLM.
+To enable concurrent CPU-GPU data transferring and LLM generation, we utilize the _copy(non_blocking=True) API in PyTorch and use separate CUDA streams for prefetching and LLM.
 In this way, the data copy operations will not block the GPU to perform computation, and thus the TeleRAG can do prefetching in parallel to the pre-retrieval LLM generation.
@@
 We then move the distances computed from CPU to GPU, merge with distances on GPU and perform global sorting on GPU.
-5 Evaluation
 Specification Value
-Dataset wiki_dpr [48]
+Dataset wiki_dpr (Karpukhin et al., 2020)
 Dataset size 2.1 billion tokens
-##\## of chunks 21 million
-##\## of IVF cluster 4096
-Embed model Contriever [35]
+#\# of chunks 21 million
+#\# of IVF cluster 4096
+Embed model Contriever (Izacard et al., 2021)
 Embed dimension 768
@@
 Index size 61 GB
-Table 2: Detailed configurations of our retrieval index.
+Table 2. Detailed configurations of our retrieval index.
+5. Evaluation
 We conduct extensive experiments to evaluate the effectiveness of TeleRAG.
-In this section, we describe the necessary details on how we set up the evaluations, present experiments results and provide in-depth analysis and discussions.
-5.1 Datasets and RAG Pipelines
+In this section, we describe the necessary details on how we set up the evaluations, present experimental results, and provide in-depth analysis and discussions.
+5.1. Evaluation Datasets and RAG Models
 Datastore.
-We build a datastore based on the wiki_dpr dataset [48], which is a popular dataset contains 2.1 billions tokens from Wikipedia.
-Following previous works [11, 48, 64], we chunk the passages by every 100 tokens, and use Contriever [35] to generate embedding for each chunk.
+We build a datastore based on the wiki_dpr dataset (Karpukhin et al., 2020), a popular dataset that contains 2.1 billion tokens from Wikipedia.
+Following previous works (Asai et al., 2023; Karpukhin et al., 2020; Min et al., 2023), we chunk the passages by every 100 tokens, and use Contriever (Izacard et al., 2021) to generate an embedding for each chunk.
 The embeddings have a hidden dimension of 768.
 Vector index.
-For the baseline retrieval, we build an IVF vector index using FAISS [25] from the datastore, following setups from  [11].
-See Table 2 for the detailed configurations of our vector index.
-As described in §4.4, we then convert the FAISS index to a customized index in PyTorch for TeleRAG.
-LLMs. We evaluate TeleRAG on the Llama model family [82] in three different sizes (Llama-3.2-3B, Llama-3-8B, Llama-2-13B) to represent different use cases.
+For the baseline retrieval, we build an IVF vector index using Faiss (Douze et al., 2024) on the datastore.
+As described in §4.4, we convert the Faiss index to a customized index in PyTorch for TeleRAG.
+See Table 2 for the detailed configurations of our vector index and datastore.
+LLMs. We evaluate TeleRAG on the Llama model family (Touvron et al., 2023) in three different sizes (Llama-3.2-3B, Llama-3-8B, Llama-2-13B) to represent different use cases.
+Setup Desktop Server
+CPU Threadripper 5975 EPYC 9554
+CPU memory size 512 GB 1.5 TB
+GPU RTX4090 (NVIDIA, 2024a) H100 (NVIDIA, 2024b)
+GPU memory size 24 GB 80 GB
+CPU-GPU Bus PCIe 4 PCIe 5
+Bandwidth n 32 (24) GB/s n n 64 (51) GB/s n
+Table 3. Hardware specifications for our setups. In bandwidth, the number in the parentheses is the actual bandwidth we measured from our system.
 RAG pipelines.
@@
 Figure 7 shows the overview of each pipeline.
-Note that even though some pipelines lack pre-retrieval generation, the post-retrieval generation serves similar functionality for the retrieval of the next iteration.
-Below are brief descriptions to evaluated RAG pipelines.
-For iterative-based pipelines, we set the maximum iterations to 3.
-Figure 7: Overview of six RAG pipelines that we evaluate.
-  1. 1.
-    HyDE[27] prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
-  2. 2.
-    SubQestion (SubQ) [^4] Implemented in LlamaIndex as SubQuestionQueryEngine[58] prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
-  3. 3.
-    Iterative (Iter) [^5] Implemented in LlamaIndex as MultiStepQueryEngine[58] prompts LLM to generate narrower questions first and iteratively refine them based on previous answers.
+Note that although some pipelines lack pre-retrieval generation, the post-retrieval generation serves similar functionality for the retrieval of the next iteration.
+Below are brief descriptions of the evaluated RAG pipelines.
+  1. (1)
+    HyDE (Gao et al., 2023a) prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
+  2. (2)
+    SubQestion (SubQ) [^3] Implemented in LlamaIndex as SubQuestionQueryEngine. (Liu, 2022) prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
+  3. (3)
+    Iterative (Iter) [^4] Implemented in LlamaIndex as MultiStepQueryEngine.
+(Liu, 2022) prompts LLM to generate narrower questions first and iteratively refine them based on previous answers.
 At the end of each iteration, it prompts LLM to judge if the answer is good enough.
-  4. 4.
-    Iter-RetGen (IRG) [78] iteratively do retrieval and LLM generation for 3 iterations.
-  5. 5.
-    FLARE [43] iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
-  6. 6.
+  4. (4)
+    Iter-RetGen (IRG) (Shao et al., 2023) iteratively do retrieval and LLM generation for 3 iterations.
+  5. (5)
+    FLARE (Jiang et al., 2023c) iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
+  6. (6)
     Self-RAG (S-RAG) [^5] We evaluate on the short-form version which only has one iteration.
-[11] decides whether to retrieve documents and then answers for each query.
-A critique stage selects the most confident answer.
-We use fine-tuned models based on Llama2-7B and Llama2-13B from their official implementation [12].
+(Asai et al., 2023) uses the LLM to judge for retrieval, generate responses, and self-critique on the responses.
+We use fine-tuned models based on Llama2-7B and Llama2-13B from their official repository (Asai et al., 2024b).
+Figure 7. Overview of six RAG pipelines that we evaluate.
 Evaluation Datasets.
-We evaluate on three commonly used question-answering datasets, NQ [54], HotpotQA [91] and TriviaQA [47].
-For each dataset, we randomly sample 512 queries and report the average unless otherwise specified.
-5.2 Experiment Setups
+We use three commonly used question-answering datasets, NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017).
+For each dataset, we randomly sample 1024 queries and report the average unless otherwise specified.
+5.2. Experiment Setups
 Hardware setups.
-We evaluate TeleRAG on two hardware environments, Desktop and Server, which equiped with represent the settings for the desktop and cloud instance use case.
-For Desktop, it’s paired with a RTX 4090 GPU (24GB memory) and we consider the 3B and 8B models.
-For Server, it’s paired with a RTX 4090 GPU (24GB memory) and we consider the 8B and 13B models.
-These sizes are common model size for RAG applications [11, 10].
-We do not evaluate the 13B model on Desktop as its model size (26 GB) exceeds RTX4090’s memory space.
-Table 3 summaries the hardware configurations.
+We evaluate TeleRAG on two hardware environments, Desktop and Server, which are equipped to represent the settings for the desktop and data center use cases.
+The Desktop has the RTX4090 (24 GB memory), and we use 3B and 8B models.
+The Server has the H100 (80 GB memory), and we use 8B and 13B models.
+These sizes represent common model sizes for RAG applications (Asai et al., 2023, 2024a).
+We do not evaluate the 13B model on Desktop as its model size (26 GB) exceeds RTX4090’s memory capacity.
+Table 3 summarizes the hardware configurations.
 Nprobe and top-$k$.
-We evaluate latency with three nprobe values, 128, 256, and 512, which represent 3.1%, 6.3%, and 12.5% of the total clusters in our evaluated index.
-This range falls into the commonly used settings of IVF index [106, 25].
-We use a top-$k$ number of 3 for retrieval, which returns 3 most relevant passages based on the retrieval query.
+For the IVF index, nprobe is a hyperparameter that controls the trade-off between search efficiency and retrieval quality.
+A common heuristic is to set nprobe to $4\sqrt{N_{c}}$ (Zilliz, 2020).
+Given our index size of $N_{c}=4096$, we use nprobe = 256 ($4\sqrt{4096}$) by default, unless otherwise specified.
+For retrieval, top-$k$ denotes the number of most relevant documents to return.
+We use top-$k$ = 3, which is a standard choice in RAG.
 RAG pipeline implementation.
-To collect the input and output texts for each stage and have consistent results across all pipelines, we implement the RAG pipelines with the FlashRAG framework [45].
-For IRG, FLARE, and S-RAG, we use the framework’s default implementations; for the other pipelines, we re-implement them using FlashRAG’s APIs.
+We implement the RAG pipelines with the FlashRAG framework (Jin et al., 2024b).
+For IRG, FLARE, and S-RAG, we use the framework’s default implementations.
+For the other pipelines, we reimplement them using FlashRAG’s APIs.
 Benchmark methodology.
-We follow a benchmarking methodology from  [101] and use GPT-3.5-Turbo [69] to run through each pipeline and record the input and output text of each step.
-During latency evaluation, we set LLM’s output length based on the recorded text.
+We follow a benchmarking methodology from SGLang (Zheng et al., 2024) and use GPT-3.5-Turbo (OpenAI, [n.
+d.]) to run through each pipeline and record the input and output text of each step.
+During latency evaluation, we set the LLM’s output length based on the recorded text.
 This way, we ensure a fair latency comparison across different LLM models.
+(a) End-to-end latency speedup with Llama-3.2-3B.
+(b) End-to-end latency speedup with Llama-3-8B.
+Figure 8. End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines and three datasets, with RTX4090 GPU. nprobe is 256.
 Baseline systems.
 To evaluate the latency of each pipeline, we construct a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline.
-For LLM generation, we use SGLang [101], which is a state-of-the-art LLM inference engine.
-Prefetching budget setups. Based on the methodology we described in §4.3, we profile each RAG pipeline with 64 random samples from NQ [54] and define the prefetch budget of each pipeline.
+For LLM generation, we use SGLang (Zheng et al., 2024), which is a state-of-the-art LLM inference engine.
+For retrieval, we use a popular retrieval library, Faiss (Douze et al., 2024), as the CPU-offloaded baseline.
+[^6] This baseline resembles frameworks like LlamaIndex and LangChain, which primarily function as wrappers around LLM serving and retrieval libraries.
+We build a system from the ground up to avoid the overhead introduced by their complex layers designed for advanced functionality.
+Prefetching budget setups.
+Based on the methodology we described in §4.2, we profile each RAG pipeline with a small amount of 64 random samples from NQ (Kwiatkowski et al., 2019) and derive the prefetching budget of each pipeline.
+(a) End-to-end throughput with six pipelines and different batch sizes using Llama-3-8B.
+(b) End-to-end throughput with six pipelines and different batch sizes using Llama-2-13B.
+Figure 9.
+End-to-end throughput of TeleRAG and the baseline across six RAG pipelines on the NQ dataset using Llama-3-8B and Llama-2-13B at different batch sizes on a H100 GPU.
+The nprobe is set to 256, and the $x$-axis represents the batch size.
 Max prefetching memory.
 We set a maximum GPU memory limit for prefetching in each configuration.
-For Server with H100, we allocate up to 16GB, sufficient for a high cluster hit rate at the largest tested nprobe value (512).
-For Desktop with RTX4090, we allocate up to 10GB and 3.75GB for the 3B and 8B models, respectively.
+For Server, we allocate up to 16 GB.
+For Desktop, we allocate up to 10 GB and 3.75 GB for the 3B and 8B models, respectively.
 These settings demonstrate that TeleRAG can efficiently operate using only a small fraction (up to 40% for RTX4090 and 20% for H100) of total GPU memory.
-Setup Desktop Server
-CPU TR 5975 EPYC 9554
-CPU memory size 512GB 1.5TB
-GPU RTX4090 [67] H100 [68]
-GPU memory size 24GB 80GB
-CPU-GPU Bus PCIe 4 PCIe 5
-Bandwidth n 32 (24) GB/s n n 64 (51) GB/s n
-Table 3: Hardware specifications for our setups. In bandwidth, the number in the parentheses is the actual bandwidth we measured from our system.
-(a) End-to-end latency speedup with Llama-3.2-3B with a single RTX4090 GPU.
-(b) End-to-end latency speedup with Llama-3-8B with a single RTX4090 GPU.
-(c) End-to-end latency speedup with Llama-3-8B with a single H100 GPU.
-(d) End-to-end latency speedup with Llama-2-13B with a single H100 GPU.
-Figure 8: End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines, three datasets, with H100 and RTX4090 GPU.
-Metric Hardware nprobe
-128 256 512
-Precision - 95.9% 97.6% 98.9%
-CPU Latency (ms) Desktop 89 176 349
-Server 64 126 251
-Table 4: Retrieval precision and baseline latency of FAISS for different nprobe values. Precision is measured by comparing top-10 results against greedy search without IVF.
-5.3 Latency Evaluation
-Baseline retrieval.
-Table 4 reports the retrieval precision and the baseline CPU retrieval latency of different nprobe.
-To conduct this test, we use 1024 random queries from NQ.
-For retrieval precision, we compare against the top-10 results from the greedy search.
-For the latency, we report the average retrieval time.
-Due to the limited scalability of the CPU, a higher nprobe directly leads to a longer retrieval latency even when using the state-of-the-art implementation from FAISS.
-End-to-end latency.
-We evaluate the end-to-end latency of all the RAG pipelines and datasets under all the hardware settings described in §5.2.
-Our comprehensive end-to-end speedup results are presented in Figure 8.
-We can observe that TeleRAG improves the execution speed consistently with various configurations.
-Across different pipelines, datasets, and model/hardware setups, the average speedup is 1.21x, 1.41x, and 1.72x for nprobe values of 128, 256, and 512, respectively.
-We observe a more significant speedup for larger nprobe, as increased nprobe settings incur longer retrieval latencies.
-In the best case, we observe around 2.6x speedup (for Iter-RetGen, HotpotQA, Llama-3.2-3B, on RTX4090).
-This is because the Iter-RetGen pipeline employs a retrieval step in every iteration and has a short LLM output in general, providing a larger end-to-end speedup margin for accelerating the retrieval part.
-Another pipeline that has a particularly high speedup is the SubQuestion pipeline, for which TeleRAG achieves around 2.5x speedup across all the datasets with Llama-3.2-3B on RTX4090.
-This pipeline uses LLM to generate multiple sub-questions and perform a batched retrieval with 3 to 4 generated queries at one time.
-The CPU retrieval performance of batched retrieval is significantly worse than that of GPU, as it has limited parallelism.
-In TeleRAG, we handle the batched retrieval efficiently by leveraging GPU’s high parallel computing capacity and, therefore, achieve a high speedup for the SubQuestion pipeline.
-Overall, the performance gain is stable across different datasets, with less than 3% of variance across datasets on average.
-For a given GPU, TeleRAG has a higher speedup with a smaller LLM because retrieval time will account for a larger fraction of the end-to-end latency.
-Still, we have tangible speedup in RTX4090 with Llama-3-8B model (e.g., more than 1.5x when nprobe is 512, on average), even though prefetching memory budget is only 3.75GB (space left after holding LLM, embedding model, IVF cluster centroids and misc tensors on GPU) in this case.
-Speedups on retrieval.
-Figure 9 shows the latency speedups on retrieval on NQ.
-We observe a consistent speedups for all the nprobe numbers but notice the greatest speedups of retrieval often occur at nprobe 256, with average speedups of 7.21x and 7.41x on RTX4090 and H100.
-This is because, with a larger nprobe, the retrieval performance of TeleRAG will start to be bounded by the missed clusters on the CPU (we have a fixed prefetch budget of different nprobe numbers).
-Nonetheless, TeleRAG still has a large latency improvement against the pure CPU baseline.
-(a) On Desktop with a RTX4090 GPU.
-(b) On Server with a H100 GPU.
-Figure 9: Retrieval speedup on NQ.
-(a) With Llama-3-8B.
-(b) With Llama-2-13B.
-Figure 10: Latency breakdown for NQ on Server with a H100 GPU.
+5.3. Evaluation Results
+Single-query latency on RTX4090.
+We evaluate the end-to-end RAG latency for single queries on Desktop equipped with RTX4090, representing a typical local scenario.
+Figure 8 shows the latency reduction of TeleRAG across three datasets and two LLMs (Llama-3.2-3B and Llama-3-8B).
+As shown in Figure 8, TeleRAG consistently outperforms the baseline across all tested configurations.
+With Llama-3.2-3B, TeleRAG achieves average speedups of $1.55\times$, $1.54\times$, and $1.49\times$ on NQ, HotpotQA, and TriviaQA, respectively.
+The best speedup of $2.11\times$ is achieved in the Iter-RetGen pipeline on HotpotQA, because this pipeline involves frequent retrieval operations and generally short LLM outputs, which enhances the relative impact of retrieval acceleration.
+Another notable improvement is in the SubQuestion pipeline, where TeleRAG achieves approximately $1.85\times$ speedup across all datasets.
+This pipeline uses LLM-generated sub-questions and performs batched retrievals of 3 to 4 queries simultaneously.
+CPU-based retrieval suffers from limited parallelism in such scenarios, but TeleRAG efficiently utilizes GPU parallelism, significantly enhancing performance.
+When deploying Llama-3-8B, speedups with TeleRAG slightly decrease compared to Llama-3.2-3B, primarily due to increased LLM latency and reduced available memory for prefetching.
+Still, TeleRAG achieves approximately $1.3\times$ speedup across datasets, with a peak improvement of $1.82\times$ for Iter-RetGen pipeline on HotpotQA.
+Achieving these results with only $3.75$ GB of remaining GPU memory (after accounting for Llama-3-8B (16 GB), the embedding model (1 GB), and other miscellaneous tensors) highlights TeleRAG’s robust capability in accelerating RAG inference even under tight GPU memory constraints.
+Figure 10. Simulated throughput [^8] Results for 1 GPU are real, not simulated.of TeleRAG on NQ dataset with different number of H100 GPUs.
+Multi-query throughput on H100.
+To evaluate the TeleRAG’s performance in batched inference, we evaluate the end-to-end throughput on Server (H100) in batch sizes 1, 2, 4, and 8.
+We show the results of six evaluated RAG pipelines with Llama-3-8B and Llama-2-13B in Figure 9.
+[…]
+Figure 11. Latency breakdown for Llama-3-8B on NQ with a H100 GPU in different batch sizes. nprobe is 256.
+Figure 12. Comparison of end‑to‑end retrieval latency for two batching strategies: naive mini‑batching and similarity‑aware greedy grouping. The Greedy Group bars are too short to be visible.
+Multi-GPU Throughput.
+We also evaluate TeleRAG in a multi-GPU system to further show the scalability.
+For the evaluation, we construct a simulation framework around our single-GPU implementation.
+We first define a global batch size that the whole system will process at once, and then we construct mini-batches to distribute among each GPU.
+To form a mini-batch, we perform a greedy search based on the embedding similarity of queries, grouping similar queries into the same mini-batch to maximize the cluster hit rate.
+We present the simulated throughput of TeleRAG with a global batch size of 128 and mini-batch size of 4 with 1 to 8 GPUs.
+As shown in Figure 8, TeleRAG’s throughput scales well with the number of GPUs.
+When compared to 1 GPU results, the averaged throughput enhancements of TeleRAG are $1.7\times$, $3.1\times$, and $5.4\times$ with $2$, 44, and $8$ GPUs, respectively.
+5.4. Analysis and Sensitivity Study
+Latency breakdown.
+We further show the latency breakdown of running RAG pipelines with Llama-3-8B and Llama-2-13B on a single H100 GPU in different batch sizes in Figure 11.
+From Figure 11, we can observe that LLM latency grows sub-linearly with larger batch sizes.
+However, the latency for Faiss retrieval on CPU grows linearly with the batch size, dominating the overall latency when batch size is large.
+These results echo our findings in Figure 9, and show limited scalability of CPU retrieval in serving scenario.
+In contrast, TeleRAG significantly accelerates across all batch sizes and achieves a higher speedup from $1.4\times$ to $1.8\times$ when the batch size increases from 1 to 8.
+Ablation on mini-batch strategies.
+In our multi-GPU results, we implement a greedy grouping strategy for mini-batching based on queries’ similarity in a global batch.
+Here, we examine the benefits and overhead of this strategy in Figure 12, where we show the end-to-end retrieval latency for mini-batch size $4$ with four GPUs and a global batch of 128 queries.
+The grouping overhead is minimal and the overall retrieval latency of greedy grouping mini-batching consistently outperform naive mini-batching across all pipelines.
+(a) Llama-3.2-3B with a RTX4090 GPU.
+(b) Llama-3-8B with a H100 GPU.
+Figure 13. Single-query retrieval speedup on NQ with different nprobe values.
+Retrieval speedups across nprobe.
+Figure 13 shows the retrieval latency reduction on NQ.
+We observe consistent speedups for all nprobe values.
+The greatest speedups are achieved at nprobe 256, with average speedups of $7.21\times$ and $7.41\times$ on RTX4090 and H100, respectively.
+As a larger nprobe value is used, the retrieval performance of TeleRAG becomes constrained by missed clusters on the CPU, given our fixed prefetch budget across varying nprobe values.
+However, RAG inference with higher nprobe will result in longer latency spent on the retrieval and hence, TeleRAG still has a significant latency improvement against the CPU retrieval baseline.
 Pipeline H100 (Llm3-8B) 4090 (Llm3-3B)
 Budget Hit Rate Budget Hit Rate
-HyDE 9 GB $77.23\%$ 7 GB $67.16\%$
-SubQ 8 GB $63.26\%$ 7 GB $58.38\%$
-Iter 5 GB $80.79\%$ 3 GB $61.46\%$
-IRG 4 GB $44.49\%$ 2.5 GB $29.39\%$
-FLARE 6 GB $58.17\%$ 3 GB $33.01\%$
-S-RAG 3 GB $36.01\%$ 1.25 GB $14.98\%$
-Table 5: The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 512.
-5.4 Analysis
-Latency breakdown and GPU compute interference.
-We further show the latency breakdown of running RAG pipelines with Llama-3-8B and Llama-2-13B on Server with H100, examining the case when the GPU’s memory capacity is enough for a good amount of prefetching.
-Figure 10 shows the breakdown of latency into LLM time and retrieval time.
-We compare the baseline’s retrieval latency (FAISS CPU) with TeleRAG’s retrieval latency in three different nprobe numbers.
-The first thing we observe is that, although we set the prefetching budget to the amount that the data loading time can overlap with the pre-retrieval generation time, there is a slight slowdown in LLM generation time for TeleRAG.
-This is potentially because GPU’s compute performance will be impacted when there is a memory-loading workload, as noticed in  [103].
-However, such impact is minimal, with less than 5% slowdown in general.
-We observe higher impacts for Iter-RetGen and Self-RAG with Llama-3-8B, which TeleRAG’s LLM geeration time is 28% and 20% longer than the baseline.
-It’s because these two pipelines have a relatively shorter LLM generation time, and thus, are more vulnerable to interference on GPU compute.
-However, with the significant amount of speedups on retrieval, TeleRAG achieves consistent speedups across all the pipelines despite having this small impact on the LLM generation.
+HyDE 9 GB $91.95\%$ 7 GB $87.42\%$
+SubQ 8 GB $79.04\%$ 7 GB $76.38\%$
+Iter 5 GB $95.51\%$ 3 GB $84.59\%$
+IRG 4 GB $59.52\%$ 2.5 GB $50.34\%$
+FLARE 6 GB $79.35\%$ 3 GB $56.67\%$
+S-RAG 3 GB $71.29\%$ 1.25 GB $29.96\%$
+Table 4. The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 256.
 Prefetch budgets and cluster hit rates.
-Table 5 shows the prefetch budgets we set with the profiling-guided approach on RTX4090 and H100 for NQ.
-It also presents the averaged cluster hit rate achieved with this prefetch budget.
-From the table, we can see that it can generally attain a high cluster hit rate ($>50$%) for the cases where it can hold a larger amount of data for prefetching.
-For pipelines or scenarios that can hold for less than 3GB prefetching amount, we generally observe a relatively lower hit rate ($<35$%), limiting our benefits of reducing CPU’s search workloads.
-However, as observed from Figure 8, TeleRAG achieves from 1.2x to 1.5x end-to-end speedups for these pipelines.
-This is the combined benefit of reducing CPU workload and utilizing GPU to perform sorting on similarity distances, demonstrating the strong performance of TeleRAG across a wide range of scenarios.
-Analysis on the cluster hit rate.
-We measure cluster hit rates across a wide range of prefetch amounts, examining scenarios with higher CPU-GPU bandwidth and sufficient GPU memory to accommodate larger prefetches.
-Our experiments are conducted on the HyDE pipeline with nprobe set to 512 and prefetching from 32 to 1024 clusters.
-We use nprobe as the prefetch metric (instead of bytes) to facilitate comparisons with optimal hit rates.
-We evaluate on NQ, HotpotQA, and TriviaQA, reporting average results in Figure 11.
-We have three observations from Figure 11.
-(1) From $0\%$ to $50\%$ of nprobe, accuracy remains high, illustrating that the probability of hitting relevant clusters decreases by order of prefetching.
-(2) Between $50\%$ and $150\%$ of nprobe, coverage still increases rapidly but is slower than the first range.
-(3) After $150\%$ of nprobe, coverage plateaus at roughly $95\%$.
-These results suggest that prefetching the amount beyond the target nprobe is likely unnecessary, as the coverage increase is limited.
-Figure 11: The cluster coverage rate for HyDE with different numbers of prefetched clusters. The target retrieval nprobe is $512$. Results are the average of NQ, HotpotQA and TriviaQA.
-6 Related Work
-6.1 Systems for RAG
-Several prior works have attempted to build systems for RAG.
-RAGCache [44] proposes a caching system that stores KV caches from datastore chunks in an order-aware manner to improve the time to the first token.
-However, this approach only reduces prefill latency, leaving retrieval and decode times unaffected, despite the fact they typically dominate total latency [4].
+Table 4 shows the prefetch budgets we set with the profile-guided approach on RTX4090 and H100 for NQ.
+It also presents the average cluster hit rate achieved with this prefetch budget.
+From the table, we can see that it generally achieves a high cluster hit rate (¿5050%) when TeleRAG has a large prefetching budget.
+For cases where the budget is less than 2 GB, we observe a relatively lower hit rate (¡5050%), limiting the benefits of reducing the CPU’s search workloads.
+However, as observed from Figure 8, TeleRAG achieves from $1.2\times$ to $1.6\times$ end-to-end speedups for these pipelines, thanks to the combined benefit of reducing CPU workload and utilizing the GPU to perform sorting on similarity distances.
+6. Related Work
+6.1. Systems for RAG
+RAGCache (Jin et al., 2024a) proposes a caching system that stores KV caches from datastore chunks in an order-aware manner to improve the time to the first token.
+This approach only reduces prefill latency, leaving retrieval and decode times unaffected, despite the fact they typically dominate total latency (Agrawal et al., 2024).
 Moreover, it assumes repeated use of the same document across multiple requests, limiting its scalability over large data stores.
-Similarly, TurboRAG [60] precomputes KV caches from the data store, but it also only optimizes prefill latency.
-RaLMSpec [98] proposes speculative retrieval and batched verification, Chameleon [41] proposes a CPU-GPU-FPGA heterogeneous architecture for accelerating the retrieval process, and PipeRAG [42] proposes an algorithm-system co-design technique to overlap the retrieval and generation process by modifying the algorithms.
-However, these works focus on the paradigm of RAG that retrieves documents once every few tokens and are not applicable to modular RAG applications that are widely used now, as discussed in §2.1.
-Speculative RAG [87] introduces drafting by smaller LLMs to reduce RAG latency, but it does not target the retrieval latency.
-APIServe (InferCept) [3] proposes a novel KV cache management strategy that can support the interception of LLM generation by other workloads, including retrieval.
+Similarly, TurboRAG (Lu et al., 2024) precomputes KV caches from the data store, but it also only optimizes prefill latency.
+CacheBlend introduces a selective KV-cache fusion technique for RAG to reuse pre-computed caches (Yao et al., 2024).
+RaLMSpec (Zhang et al., 2024b) proposes speculative retrieval and batched verification, Chameleon (Jiang et al., 2023d) proposes a CPU-GPU-FPGA heterogeneous architecture for accelerating the retrieval process, and PipeRAG (Jiang et al., 2024) is an algorithm-system co-design technique to overlap the retrieval and generation by modifying the algorithms.
+However, these works focus on the paradigm of RAG that retrieves documents once every few tokens and do not apply to modular RAG applications that are widely used now, as discussed in §2.1.
+Speculative RAG (Wang et al., 2024b) introduces drafting by smaller LLMs to reduce RAG latency, but it does not target the retrieval latency.
+APIServe (InferCept) (Abhyankar et al., 2024) proposes a novel KV cache management strategy that can support the interception of LLM generation by other workloads, including retrieval.
 However, this work again does not focus on optimizing retrieval latency.
+EdgeRAG (Seemakhupt et al., 2024) reduces the memory requirement of retrieval by re-generating and caching document embeddings at runtime, targeting extremely resource-constrained environments.
+RAGO (Jiang et al., 2025) introduces an abstraction for RAG to automatically pick task placement, resource allocation, and batching policies.
 Unlike all these prior works, we tackle system challenges of long retrieval latency and large memory requirement for modular RAG pipelines with large-scale datastores.
-6.2 Systems for Compound LLM Applications
-Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities [95, 14, 85].
-LLMCompiler [52] is a framework that optimizes the execution of multiple functions in large language models by enabling parallel function calling.
-AI Metropolis [90] accelerates LLM-based multi-agent simulations with out-of-order execution.
-ALTO [76] improves the performance of complex AI systems by streaming partial results.
-Some open-source libraries provide graph-based representation for programming this kind of application [22, 56].
+6.2. Systems for Compound LLM Applications
+Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities (Zaharia et al., 2024; Berger and Zorn, 2024; Wang et al., 2024a).
+LLMCompiler (Kim et al., 2023a) is a framework that optimizes the execution of multiple functions in large language models by enabling parallel function calling.
+AI Metropolis (Xie et al., 2024) accelerates LLM-based multi-agent simulations with out-of-order execution.
 RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
-6.3 Vector Index
-Vector index is gaining growing attention as an essential component of RAG [16, 29], and many works have been proposed to improve the efficiency with system or algorithm techniques.
-ScaNN proposes an anisotropic quantization method for better efficiency [32].
-DiskANN and SPANN propose memory-disk hybrid indexing systems that work beyond the limitation of memory capacity [19, 38].
-Other prior works propose hardware acceleration of vector index with GPUs [46], FPGAs [96, 40], TPU [20], or ray tracing hardware [104, 59].
-BANG proposes a method to scale graph-based ANN beyond GPU memory [49], and Rummy allows the index to scale beyond GPU memory capacity with reordered pipelining [99].
-However, these methods either require algorithm modifications or are bottlenecked by CPU-GPU bandwidth.
+6.3. Vector Index
+Vector index is a key component of RAG (Briggs et al., [n.
+d.]; Gao et al., 2023b), and many works have been proposed to improve their efficiency.
+ScaNN proposes an anisotropic quantization method for better efficiency (Guo et al., 2020).
+DiskANN and SPANN propose memory-disk hybrid indexing systems that work beyond the limitation of memory capacity (Chen et al., 2021; Jayaram Subramanya et al., 2019).
+Other prior works propose hardware acceleration of vector index with GPUs (Johnson et al., 2019), FPGAs (Zeng et al., 2022; Jiang et al., 2023a), TPU (Chern et al., 2022), or ray tracing hardware (Zhu, 2022; Liu et al., 2023).
+BANG proposes a method to scale graph-based ANN beyond GPU memory (Khan et al., 2024), and Rummy allows the index to scale beyond GPU memory capacity with reordered pipelining (Zhang et al., 2024a).
+These methods either require algorithm modifications or are bottlenecked by CPU-GPU bandwidth.
 Our proposal, focuses on the context of modular RAG applications, where queries for retrieval are usually generated by LLMs, and optimizes the latency and the GPU memory consumption without altering the algorithm of the IVF index.
-7 Conclusion
+7. Conclusion
 In this paper, we introduced TeleRAG, an inference system that tackles the system challenges of RAG pipelins under latency-sensitive scenarios.

### v3

@@
-TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval
-Abstract.
+
+Abstract
 […]
-RAG inference, Acceleration, Prefetching.
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: submissionid: 164
-1. Introduction
-(a) Typical pipeline stages of a RAG application.
-(b) Illustration of TeleRAG’s lookahead retrieval mechanism and comparison to different RAG systems.
-TeleRAG prefetches relevant retrieval data from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU-CPU cooperation.
-Figure 1. Overview of RAG pipeline stages and comparison of TeleRAG to baseline systems.
-Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020).
-During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and incorporate up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
-Modern RAG applications share two key characteristics.
-(1) RAG applications are built as modular pipelines as shown in Figure 1(a), where a single query undergoes multiple rounds of LLM calls and retrievals, each of which serves different functions to improve overall output quality.
-For example, query transformation rounds (Gao et al., 2023a; Zheng et al., 2023; Peng et al., 2023; Asai et al., 2024a) generally occur before retrieval to refine the user’s query with LLMs (pre-retrieval generation).
-(2) RAG’s datastores are typically large, supported by recent works demonstrating that larger datastore consistently improves the performance of RAG applications (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n.
-d.]).
-These characteristics create several significant challenges for efficient RAG inference, especially in latency-sensitive applications such as customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024), financial analysis (Malec, 2024; MyScale, 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024).
-First, the latency of RAG systems is inherently increased due to multiple rounds of LLM generation and retrieval.
-Second, although GPUs can speed up both LLM generation and retrieval, the combination of these processes poses significant memory requirements, often exceeding available GPU resources (see §3.1).
-Hence, it is expensive or even infeasible to fully host both operations on GPUs.
-Consequently, in local or custom deployments, which are common for RAG applications with private or sensitive data, the retrieval datastore is often offloaded to CPU memory to alleviate GPU memory constraints.
-However, while CPU offloading resolves memory limitations, it significantly increases retrieval latency (see §3.1), diminishing overall system efficiency.
-This memory pressure is not limited to local deployments; even for serving scenarios in data centers, where many concurrent requests are processed together and GPUs are plentiful, managing large index sizes still presents a significant challenge for achieving high throughput with latency service level objectives (SLOs).
-Allocating additional GPU memory to the retrieval data store directly reduces the memory available for the KV cache of the LLM serving system, limiting the effective batch sizes (Kwon et al., 2023).
-In contrast, CPU retrieval would increase the per-request latency by a lot, highlighting the need for techniques that reduce GPU memory consumption while preserving low latency.
-To address the latency challenge, several recent works have been proposed to accelerate RAG inference.
-These approaches include hardware accelerators for retrieval (Jiang et al., 2023d; Quinn et al., 2025), reuse of KV-cache (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024), and speculative document retrieval (Wang et al., 2024b; Zhang et al., 2024b).
-However, these works do not address the substantial memory demands associated with retrieval.
-Therefore, designing a memory-efficient system for low-latency RAG deployments is crucial.
-In this paper, we argue that a promising direction to reduce GPU memory usage while maintaining low latency lies in improving the design of the retrieval index.
-Specifically, the inverted file index (IVF) (Sivic and Zisserman, 2003) reduces retrieval latency by pre-clustering data and limiting runtime search to relevant clusters.
-It offers a way to leverage GPU-accelerated search while not increasing memory consumption by dynamically transferring only the necessary clusters from the CPU to the GPU.
-However, this data transfer overhead is a key bottleneck, and needs to be minimized (Khan et al., 2024) (see §3.2).
+footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
+1 Introduction
+Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases gao2023retrieval_rag_survey; asai2024reliable; lewis2020retrieval.
+During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations mallen2022not; ram2023context; khandelwal2019generalization and incorporate up-to-date or private information izacard2023atlas; min2023silo.
+To provide accurate responses, modern RAG applications employ multiple rounds of LLM calls and retrievals for a single query gao2023retrieval_rag_survey; fan2024_rag_survey; advanced-rag-overview; modular-rag-survey; databricks-blog; azure-blog.
+As shown in Figure 1, RAG pipelines typically consist of four stages: (1) pre-retrieval generation: refines the initial user query to improve retrieval effectiveness.
+(2) retrieval: finds the relevant documents from datastore; (3) post-retrieval generation: generates a response based on the user’s query and retrieved documents; (4) judgment: Evaluates the output to determine the execution flow.
+RAG datastores are typically large, as datastore size is strongly correlated with model accuracy min2023silo; khandelwal2019generalization; borgeaud2022improving; hardt2023test; shaoscaling.
+However, a brute-force search on such a large datastore is prohibitively slow and inefficient, as most data is irrelevant to the query.
+This drives the adoption of solutions like the Inverted File Index (IVF) ivf_original, which partitions the data into clusters and restricts the search to only the most relevant ones.
+Figure 1: Typical pipeline stages of a RAG application.
+While this partitioning effectively reduces the computational burden of the search, it does not solve the memory capacity challenge.
+To accelerate the IVF retrieval process with the GPU, one straightforward approach is to keep the entire datastore in the GPU memory (case A in Figure 2).
+As the size of a datastore is large (tens to thousands GB) quinn2025_iks, such a setting is expensive and often commercially impractical for local and custom RAG deployments handling private or sensitive data.
+Even in large data centers, where user requests are often batched, this approach challenges service level objectives (SLOs) because allocating GPU memory to the datastore reduces the LLM’s key-value cache (KV cache), limiting effective batch sizes kwon2023efficient.
+An alternative approach is to load the relevant data clusters from CPU to GPU at runtime.
+However, because bandwidth between the CPU and GPU is typically limited, this approach suffers from high transfer latency.
+As a result, modern RAG systems jiang2025rago; shen2025hermes process retrieval on CPUs as they typically offer large and relatively cheap memory.
+However, CPUs are not as efficient as GPUs for highly parallelizable operations such as vector similarity search, and therefore using CPUs for search (case B in Figure 2) increases RAG system latency.
+For latency-sensitive applications such as customer chatbots vonage_reducing_rag_latency; akkiraju2024facts; ontinue_latency_ion_iq_chatbot, financial analysis harchworks_rag_financial; myscale_rag_trading, and emergency medical diagnosis klang2024assessing; apollo_medlm_rag, high latency can lead to poor user experience and even critical failures.
+To make matters worse, any latency increase of RAG systems has compounding effects due to multiple rounds of LLM generation and retrieval.
+Figure 2: Illustration of TeleRAG’s lookahead retrieval mechanism and comparison to different RAG systems.
+TeleRAG prefetches relevant retrieval data from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU–CPU cooperation.
 Our proposal.
-To tackle latency and memory bottlenecks in RAG inference, we introduce a novel mechanism called lookahead retrieval, which predicts the subsets of IVF clusters that will likely be accessed during runtime and proactively transfers them from the CPU to the GPU before the retrieval stage.
-The key observation behind lookahead retrieval is that there is substantial semantic overlap between queries before and after the pre-retrieval stage.
-Namely, IVF clusters used by the initial input query, which is available well before the retrieval stage, are also likely to be relevant to the actual retrieval query (see §3.3).
-Using this insight, we propose TeleRAG, an efficient inference system designed to optimize RAG latency while minimizing GPU memory consumption.
-TeleRAG employs lookahead retrieval to preemptively load relevant IVF clusters onto the GPU, effectively hiding CPU-GPU data transfer overhead during concurrent LLM generation.
-As illustrated in Figure 1(b), this approach significantly reduces retrieval latency without exceeding GPU memory constraints, enabling efficient execution of RAG applications.
-To maintain retrieval accuracy while adopting lookahead retrieval, TeleRAG simultaneously searches for clusters that were not prefetched (i.e., mispredicted) from the CPU and merges the results.
-TeleRAG also supports batch and multi-GPU inference.
-A key challenge in deploying lookahead retrieval is determining the optimal number of IVF clusters to prefetch: Prefetching too many clusters increases data-transfer overhead, whereas prefetching too few could result in higher retrieval latency due to increased CPU processing.
-To address this, we propose a profile-guided approach coupled with an analytical model that dynamically determines the ideal prefetch amount based on pipeline characteristics and hardware configurations.
+In this work, we observe that there is a substantial semantic overlap between a user’s initial query and the query refined by an LLM during the pre-retrieval stage (see Figure 1).
+Since both queries represent the same core information need, their relevant IVF clusters are also likely to overlap significantly (see §3.3 for detailed analysis).
+Using this insight, we propose TeleRAG4), an efficient inference system that optimizes RAG performance by leveraging GPU retrieval acceleration while minimizing GPU memory consumption.
+TeleRAG employs lookahead retrieval to proactively load relevant IVF clusters onto the GPU, hiding the CPU–GPU data transfer overhead during concurrent LLM generation.
+As illustrated in Figure 2, our approach aims to significantly reduce retrieval latency without exceeding GPU memory constraints.
+To guarantee retrieval accuracy, TeleRAG complements this prefetching with a hybrid search: any clusters missed by the lookahead are simultaneously searched on the CPU, and the results are merged.
+In addition to prefetching, TeleRAG also utilizes on-GPU caching to further reduce redundant data transfers for frequently accessed clusters.
+TeleRAG also optimizes for batch and multi-GPU inference (§4.2).
+To mitigate increased data transfer in batched scenarios, TeleRAG uses a prefetching scheduler to group similar queries, maximizing the overlap of their prefetched clusters.
+In multi-GPU settings, TeleRAG employs a cache-aware query scheduler that routes queries to the appropriate GPU to maximize the utility of its cached clusters.
 […]
 In summary, we make the following key contributions:
-  • Analyzing the correlation of the queries between the pre-retrieval generation and retrieval stages, revealing significant overlap in their corresponding IVF clusters.
-  • Proposing lookahead retrieval, which prefetches likely IVF clusters to the GPU, and hides CPU-GPU data transfer time during pre-retrieval generation.
-  • Developing TeleRAG, an efficient RAG inference system that integrates lookahead retrieval, resulting in significant acceleration of RAG with minimal GPU memory usage.
-2. Background
-Figure 2. Overview of RAG.
-2.1. RAG
-RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020).
-The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and generate contents based on up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
-RAG workflow.
-A basic RAG workflow involves several phases, including data store building, retrieval, and interactions with LLMs (Dixit, 2024; Ilin, 2023; Databricks, 2024; Wei, 2024; Vijeaswaran et al., 2024; Pathway, 2024; Gao, 2024).
-In order to build a data store, raw data in various formats is cleaned, converted to plain text, and divided into chunks.
-These chunks are then encoded into vectors using an embedding model and stored in a vector index, enabling efficient searching based on similarity.
-When a user provides a query, it is converted into a vector using the same embedding model, and the most similar chunks from the indexed database are retrieved.
-For a large-scale index, approximate algorithms such as the inverted file index (IVF) (Sivic and Zisserman, 2003) (see §2.2 for detail) are commonly used to accelerate the search process.
-The retrieved chunks, along with the original user query, are combined and given as a prompt to the LLM.
-The LLM then generates a response that relies on the information provided in the retrieved documents.
-This approach enables dynamic and informative responses in both single- and multi-turn dialogues (Lewis et al., 2020).
-Modularity in modern RAG applications.
-However, naively integrating document retrieval into LLM generation can cause several issues.
-For example, the retrieval often struggles to find relevant content and may select irrelevant or mismatched chunks (Gao et al., 2023a; Zheng et al., 2023; Ma et al., 2023), and a single retrieval may not provide sufficient context, necessitating multiple retrieval rounds (Gao et al., 2023b).
-To solve these issues, most state-of-the-art RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  (Gao et al., 2023b; Fan et al., 2024; Ilin, 2023; Gao, 2024; Databricks, 2024; Wei, 2024).
-Typically, they have the following types of steps (shown in Figure 2):
-  • Pre-retrieval generation, used to assess whether retrieval is needed or to generate queries for retrieval.
-An example of a pre-retrieval technique is query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022), which reformulates the original query to make it clearer and more suitable for the retrieval task.
-  • Retrieval, used to identify relevant documents from the vector data store. This stage takes the output of pre-retrieval generation and generates data for the next stage.
-  • Post-retrieval generation, generates response based on user query and retrieved documents.
-It can also perform additional process such as summarization (Jiang et al., 2023b; Kim et al., 2023b) or reranking (Zhuang et al., 2023; Stewart and Linsdell, 2024) on the retrieved documents.
-  • Judgment, dynamically determines the execution flow. For example, it decides if more iteration is needed to enhance the response. Heuristics or LLMs can be used for judgement stage.
-By proceeding through these functions, RAG applications can deliver more precise and contextually appropriate responses, significantly enhancing the capabilities of LLMs for various applications (Jin et al., 2024b; Gao et al., 2023b).
-2.2. Vector Index and Inverted File (IVF)
+  • Analysis of the correlation of queries between the pre-retrieval generation and retrieval stages, revealing significant overlap in their corresponding IVF clusters.
+  • Lookahead retrieval, which prefetches likely IVF clusters to the GPU, and hides CPU–GPU data transfer time during pre-retrieval generation.
+  • TeleRAG, an efficient RAG inference system that integrates lookahead retrieval and supports high-performance multi-GPU inference through prefetching and cache-aware scheduling, resulting in significant acceleration of RAG with minimal GPU memory usage.
+2 Background
+2.1 Retrieval-Augmented Generation (RAG)
+RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text gao2023retrieval_rag_survey; asai2024reliable; lewis2020retrieval.
+The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations mallen2022not; ram2023context; khandelwal2019generalization and generate contents based on up-to-date or private information izacard2023atlas; min2023silo.
+Figure 3: Overview of RAG.
+Modern RAG pipelines.
+In order to deliver precise and contextually appropriate responses, modern RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  gao2023retrieval_rag_survey; fan2024_rag_survey; advanced-rag-overview; modular-rag-survey; databricks-blog; azure-blog.
+Typically, they have the following types of steps (shown in Figure 3): (1) Pre-retrieval generation decides whether retrieval is needed or polishes queries before retrieval; for example, query transformation ma2023query; jagerman2023query; gao2022precise_hyde; zhou2022least; ye2023enhancing; press2022measuring; peng2024large; zheng2023take; llamaindex reformulates the original query to make it clearer and more suitable for retrieval.
+(2) Retrieval then identifies relevant documents for the query, taking the output of pre-retrieval generation and producing the evidence for the next stage.
+(3) Post-retrieval generation uses the user’s query and the retrieved documents to produce the final response; it can also include processes like summarization jiang2023longllmlingua; kim2023sure or reranking zhuang2023open; cohere-rerank.
+(4) Judgment component dynamically determines the execution flow (e.g., deciding whether further iteration is needed to enhance the response) using heuristics or LLMs.
+2.2 Datastore, Inverted File Index (IVF), and Retrieval
+Datastore.
+The datastore is processed by cleaning, chunking, and indexing to enable efficient retrieval.
+The raw data, in various formats, is first cleaned, converted to plain text, and divided into chunks.
+The chunks are then converted into vector embeddings using an embedding model such as Contriever izacard2021unsupervised.
+The document chunks, along with vector embeddings, are stored in a vector database, which enables efficient retrieval by searching for chunks based on vector similarities between the query and the chunk embeddings.
+IVF-based retrieval.
+Many RAG deployments use the Inverted File Index (IVF) algorithm ivf_original for efficient vector search.
+IVF partitions the vectors into clusters based on similarity, each represented by a centroid.
+The search is then restricted to only the few clusters most relevant to the query, avoiding searching the entire datastore.
+With clustering, an IVF search is a two-step process: (1) Cluster probing: The query is compared only to the cluster centroids to quickly identify the few most relevant clusters.
+(2) Searching: The system then performs a detailed search only within that small set of selected clusters and returns the final top-$k$ best-matching items.
+A parameter called nprobe douze2024faiss controls the number of clusters to check in the first step.
+A larger nprobe increases accuracy at the cost of higher latency, as more data must be searched.
+Appendix A provides a formal description of IVF.
+Latency of retrieval.
+Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs.
+Furthermore, open-source libraries offer efficient GPU implementations johnson2019billion; rapidsai.
+However, the vector database in modern RAG deployments can reach tens to thousands of GB quinn2025_iks; sella2024flash as a result of the positive correlation of datastore size with accuracy min2023silo; khandelwal2019generalization; borgeaud2022improving; hardt2023test; shaoscaling.
+The large vector databases increase the latency of RAG pipelines significantly in both local setups and data centers.
+Local setups typically consist of personal workstations or laptops with a single consumer-grade GPU (e.g., RTX4090 has 24 GB of memory), where latency is critical.
+Data center environments, on the other hand, care about both latency and throughput since they handle numerous simultaneous user queries.
+In both scenarios, GPU memory capacity becomes a performance limitation when working with vector databases.
+Consumer GPUs in personal setups lack sufficient memory to accommodate large vector databases, whereas data-center GPUs must sacrifice KV cache space (reducing throughput) to accommodate them kwon2023efficient.
+Consequently, vector databases are commonly stored in the CPU memory instead.
+Unfortunately, this approach increases retrieval latency whether IVF search operations are performed on the CPU or data must be transferred between CPU and GPU (as explored in §3).
+3 Analyzing RAG Latency
+In this section, we analyze the system challenges in state-of-the-art RAG applications.
+We constructed a 61 GB vector index with the Faiss library douze2024faiss, and set the IVF clusters to 4096 following common practice asai2023self.
+Our analysis is conducted on an RTX 4090 (24 GB) and H100 (80 GB) using Llama-3-8B llama3modelcard.
+Further setup details are in §5.2.
+3.1 End-to-end Latency of RAG Pipelines
+(a) Batch = 1.
+(b) Batch = 4.
+Figure 4: Latency breakdown of six RAG pipelines on NQ dataset kwiatkowski2019natural with one H100 GPU.
+We first analyze the end-to-end latency of six RAG pipelines (§5.1) by comparing two scenarios: (1) LLM on GPU with retrieval on CPU (low GPU memory), and (2) both LLM and the vector index on GPU (low latency).
+We set the nprobe to 256, a commonly used setting under this index’s configuration (see §5.2 for details).
+As shown in Figure 4 (using randomly sampled 1024 NQ queries kwiatkowski2019natural), the CPU-based retrieval phase is a major bottleneck, consuming 41.1% and 60.5% of total latency for batch sizes 1 and 4, respectively.
+In contrast, GPU-accelerated retrieval accounts for only 10.5% and 28.3% of the latency, respectively.
+On average, GPU retrieval is $5.96\times$ (batch size 1) and $3.87\times$ (batch size 4) faster than its CPU counterpart, reducing overall end-to-end latency by $1.5\times$ and $1.8\times$.
+Thus, accelerating GPU-based retrieval is crucial for reducing end-to-end latency.
+However, GPU acceleration has a significant memory cost.
+For instance, in our setting, with a 61 GB index and a 16 GB Llama-3-8B model, GPU retrieval requires 77 GB of GPU memory, far exceeding the 24 GB of a common RTX 4090.
+This makes GPU-accelerated retrieval often unfeasible on lower-end GPUs or with large datastores.
+Even on large GPUs (e.g., H100), storing the full index in memory limits serving throughput.
+High-throughput batched inference is bottlenecked by the LLM’s KV cache capacity kwon2023efficient.
+The index’s large memory footprint reduces the available memory for this KV cache, thereby limiting the effective batch size.
+This memory contention is exacerbated by RAG’s typically long contexts jin2024ragcache; lu2024turborag; yao2024cacheblend and the growing size of models and indexes min2023silo; khandelwal2019generalization; borgeaud2022improving; hardt2023test; quinn2025_iks, necessitating offloading.
+Therefore, in the rest of this section, we try to answer the following question: Is it possible to achieve the latency of GPU-based retrieval while using much less GPU memory?
+Figure 5: Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
+3.2 GPU-accelerated Retrieval with Runtime Transfer
+A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on demand at runtime, leveraging the IVF index to narrow the search space.
+While this enables fast GPU search, the data transfer itself becomes the new bottleneck.
+Figure 5 compares the latency of the runtime fetching system against CPU retrieval on an RTX 4090 GPU with three different nprobe values that determine the amount of data fetched.
+Overall, fetch time dominates latency due to the limited CPU–GPU PCIe bandwidth (32 GB/s).
+Although the GPU search is substantially faster, the fetch overhead results in a higher end-to-end latency ($\sim$3% slower on average).
+Thus, to achieve a meaningful speedup with this approach, the data-fetching latency must be effectively hidden.
+To hide data-fetch costs, CPU-to-GPU transfers must occur before the retrieval stage, which requires predicting which data will be accessed.
+Fortunately, modern RAG pipelines offer a valuable hint: the query from the previous step.
+Dataset HyDE SubQ Iter IRG FLARE S-RAG
+NQ 73.1% 63.2% 91.5% 83.8% 79.1% 100.0%
+HotpotQA 75.3% 62.5% 89.6% 89.4% 80.2% 100.0%
+TriviaQA 73.1% 61.6% 86.2% 86.1% 81.7% 100.0%
+Table 1: IVF cluster overlapping rate between the input and output of the pre-retrieval generation. Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
+Figure 6: The overview of TeleRAG system.
+When queries arrive, some clusters are already cached on the GPU from previous requests.
+These queries are grouped into micro-batches based on semantic similarity.
+For each micro-batch, TeleRAG prefetches the required clusters using lookahead retrieval in parallel with the pre-retrieval generation.
+During retrieval, TeleRAG performs a hybrid search: the GPU processes cache hits directly from GPU memory, while the CPU handles cache misses.
+The retrieved documents are then passed to the post-retrieval generation stage.
+3.3 Overlapping of IVF Clusters
+While the exact data to be retrieved is only known after the pre-retrieval generation is done, we observe that there are high similarities of the IVF clusters assignments between the queries at different stages.
+Similarity of queries at different stages.
+During RAG pre-retrieval stages (e.g., query transformation gao2022precise_hyde; zheng2023take; peng2023_beque; openscholar), an LLM refines an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$ for the actual retrieval.
+While this process often rewrites or simplifies the query, it intuitively preserves its core semantic content.
+Consequently, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar, suggesting their target IVF clusters will significantly overlap.
+Therefore, $q_{\text{in}}$ can serve as a valuable hint for predicting $q_{\text{out}}$.
+Prediction coverage.
+To verify this hypothesis, we evaluated the average cluster coverage (predicted vs.
+actual) in three popular QA datasets (NQ kwiatkowski2019natural, HotpotQA yang2018hotpotqa, and TriviaQA joshi2017triviaqa) and six RAG pipelines.
+As Table 1 shows, when prefetching 256 clusters, the overlap is consistently high.
+For instance, even the lowest reported coverage (for SubQ) remains above 61.6%.
+Opportunity.
+This data shows an opportunity to predict required clusters, hiding data transfer overhead during LLM generation.
+In this paper, we aim to leverage this observation to accelerate the inference latency for RAG.
+4 Design of TeleRAG
+Driven by the need for high-performance but memory-efficient RAG inference, we propose TeleRAG, an end-to-end system that accelerates retrieval through its core lookahead retrieval mechanism, which accelerates the retrieval process with a GPU but only requires a small fraction of the datastore on the GPU memory.
+Alongside this technique, TeleRAG further designs optimizations on caching, prefetching amount, and scheduling algorithms for batching and multi-GPU scenarios, forming a complete system for both local and cloud usages.
+This section details the lookahead retrieval mechanism (§4.1), and then describes how TeleRAG extends it with batching and multi-GPU support (§4.2), followed by other system-level optimizations (§4.3).
+Appendix C provides implementation details.
+4.1 Lookahead Retrieval Mechanism
+The core mechanism of TeleRAG, lookahead retrieval, is inspired by the observation in §3.3 that queries across different RAG stages are highly correlated and therefore select overlapping IVF clusters.
+Building on this insight, lookahead retrieval predicts and prefetches likely-needed clusters to the GPU during LLM generation, overlapping data transfer with computation.
+During retrieval, TeleRAG leverages lookahead retrieval to coordinate the CPU and GPU so that prefetched data are processed on the GPU, while the CPU concurrently handles any missed clusters.
+This cooperative execution forms the foundation of TeleRAG’s low-latency, memory-efficient design.
+Figure 6 shows the overview of TeleRAG with lookahead retrieval.
+Let $q_{\text{in}}$ denote the query input to the pre-retrieval stage and $q_{\text{out}}$ denote the output passed to the retrieval stage.
+The IVF clusters selected by $q_{\text{in}}$ and $q_{\text{out}}$ are highlighted with a green background ($C_{\text{in}}$) and a purple background ($C_{\text{out}}$), respectively.
+Due to the semantic similarity between $q_{\text{in}}$ and $q_{\text{out}}$, there is substantial overlap between $C_{\text{in}}$ and $C_{\text{out}}$.
+Guided by this correlation, lookahead retrieval operates in the following steps:
+  1. 1.
+    Predict & prefetch: During LLM generation, prefetch IVF clusters likely to be used in retrieval to GPU memory, identified by their distance to $q_{\text{in}}$.
+This transfer is performed asynchronously via GPU DMA, overlapping with ongoing LLM computation.
+  2. 2.
+    GPU similarity search: Once $q_{\text{out}}$ is available, the GPU efficiently searches the predicted clusters ($C_{\text{overlap}}$) already resident in GPU memory.
+  3. 3.
+    CPU similarity search: Concurrently, the CPU performs similarity search over the remaining clusters ($C_{\text{miss}}$) that were not prefetched.
+  4. 4.
+    Merge: The search results from GPU and CPU are merged on GPU. The retrieval documents are fed to LLM for post-retrieval generation on GPU.
+In summary, lookahead retrieval enables TeleRAG to accelerate retrieval by overlapping data prefetching with LLM generation and distributing the similarity search between GPU and CPU.
+This design significantly reduces CPU computation and data transfer latency, forming the backbone of TeleRAG’s efficiency.
+Prefetching amount.
+Prefetching more clusters improves hit rate and retrieval speed, but also increases CPU–GPU transfer time.
+If this transfer exceeds the pre-retrieval generation window, the overlap advantage is lost and may add latency, though slight overruns are acceptable if retrieval savings are substantial.
+Through analytical modeling and empirical profiling on modern hardware, we found prefetching up to the pre-retrieval LLM generation time ($t_{\mathrm{LLM}}$) achieves the optimal balance between latency reduction and transfer overhead.
+A detailed derivation of this result is provided in Appendix B.
+Since $t_{\mathrm{LLM}}$ varies per query, for a given pipeline and hardware bandwidth $B_{\mathrm{link}}$, we estimate its average ($\overline{t}_{\mathrm{LLM}}$) on a calibration set and set the prefetching amount to $B_{\mathrm{link}}\times\overline{t}_{\mathrm{LLM}}$.
+Figure 7: Overview of TeleRAG’s multi-GPU system design.
+The prefetching scheduler clusters semantically similar queries into micro-batches, while the cache-aware scheduler assigns these batches to GPUs based on cache locality.
+4.2 Batching and Multi-GPU System
+While §4.1 describes lookahead retrieval on a single query and single GPU basis, we now demonstrate how TeleRAG extends this mechanism to support batches and multi-GPU inference.
+The system’s multi-GPU architecture is illustrated in Figure 7.
+Batching design.
+For batched scenarios, a challenge for lookahead retrieval is that each query requires a different set of IVF clusters.
+In our design, we apply the fixed total prefetching budget as described in §4.1, and distribute it equally among all queries in the batch.
+Although the per-query hit rate drops as the batch size grows, this strategy provides balanced retrieval acceleration for each query.
+Multi-GPU support.
+In multi-GPU scenarios that serve a large number of RAG queries, TeleRAG divides a global batch of queries into micro-batches and assigns them to each available GPU.
+Each GPU then independently processes the full RAG pipeline for its assigned micro-batches.
+Scheduling optimization for prefetching.
+To compensate for the reduced hit rate in larger batches, we introduce a prefetching scheduler (see Figure 7).
+It operates by performing a greedy search on the global batch, grouping queries by the lowest L2 distance.
+While this scheduling adds latency, the greedy search is highly efficient on GPUs, incurring negligible overhead.
+Our profiling shows that for batch sizes up to 256, the search latency is less than $0.1$ s, which is minimal compared to the end-to-end batch latency.
+Cache design.
+To improve performance, TeleRAG implements a dynamic caching strategy for data clusters.
+Instead of discarding all clusters after a request is served, it retains the most frequently accessed ones in GPU memory.
+This allows subsequent requests to reuse these cached clusters, boosting the cache hit rate and prefetching efficiency.
+Cache-aware scheduling.
+To further exploit the benefits of caching, we design a cache-aware scheduler (Figure 7) that assigns micro-batches to GPUs based on cache locality.
+The scheduler employs a greedy strategy: it first selects the micro-batch with the greatest overlap across all GPU caches and assigns it to the best-matching GPU.
+It then iteratively schedules the remaining micro-batches in descending order of overlap with each GPU’s cached clusters.
+Although this scheduling introduces minor overhead, the performance gains from improved cache locality outweigh the cost (§5.4).
+4.3 System Optimizations
+GPU sorting.
+TeleRAG accelerates the final sorting stage of IVF search by leveraging the GPU.
+To achieve this, TeleRAG transfers the scalar distance values of $C_{\text{miss}}$ from the CPU to the GPU.
+Unlike transferring full vector data, this operation is lightweight and incurs negligible overhead.
+The GPU then performs a global sort over the combined distances of $C_{\text{miss}}$ and $C_{\text{overlap}}$.
+Prefetching target.
+TeleRAG prefetches a fixed byte budget ($b_{p}$), rather than a cluster count, to ensure predictable transfer times given highly uneven cluster sizes.
+The system fills this budget by adding whole clusters sequentially based on query proximity.
+If the next closest cluster exceeds the remaining budget, it is skipped entirely, ensuring a clean GPU/CPU processing split.
+Prefetching for multi-round.
+For multi-round RAG involving the same input query, the system performs a full prefetch (up to the budget) only in the first round, leveraging high cluster similarity across rounds.
+In subsequent rounds, it incrementally fetches only the additional required clusters that were not loaded previously, optimizing data transfer while respecting the memory budget.
+5 Evaluation
+We conducted extensive experiments to evaluate the effectiveness of TeleRAG.
+In this section, we describe the necessary details on evaluation setups, present experimental results, and provide in-depth analysis and discussions.
+5.1 Evaluation Datasets and RAG Models
+Datastore.
+We built a datastore based on the wiki_dpr dataset karpukhin2020dense, a popular dataset that contains 2.1 billion tokens from Wikipedia.
+Following previous works asai2023self; karpukhin2020dense; min2023silo, we chunked the passages by every 100 tokens, and used Contriever izacard2021unsupervised to generate an embedding for each chunk.
+The embeddings have a hidden dimension of 768.
+The index size is 61 GB and we clustered the embeddings into 4096 IVF clusters.
+See Appendix D for detailed configurations of the index.
+LLMs. We evaluated TeleRAG on the Llama model family touvron2023llama in two different sizes (Llama-3.2-3B, Llama-3-8B) to represent different use cases.
+Setup Desktop Server1 Server2
+CPU Threadripper 5975 EPYC 9554 EPYC 9534
+CPU mem. size 512 GB 1.5 TB 1.5 TB
+GPU RTX4090 H100 $8\timesH$200
+GPU mem. size 24 GB 80 GB 140 GB
+CPU–GPU Bus PCIe 4 PCIe 5 PCIe 5
+Bandwidth 32 GB/s 64 GB/s 64 GB/s
+Table 2: Hardware specifications for our setups.
+Figure 8: Overview of six RAG pipelines that we evaluate.
+RAG pipelines.
+We evaluated TeleRAG with six popular RAG pipelines, as depicted in Figure 8.
+Note that even in pipelines lacking a pre-retrieval stage, the post-retrieval generation serves a similar function for the next retrieval iteration.
+Below are brief descriptions of these pipelines.
+  1. 1.
+    HyDE gao2022precise_hyde prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
+  2. 2.
+    SubQuestion (SubQ)  sub_question_llamaindex prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
+  3. 3.
+    Iterative (Iter)  multistep_llamaindex prompts LLM to generate narrower questions first and iteratively refine them based on previous answers.
+At the end of each iteration, it prompts LLM to judge if the answer is good enough.
+  4. 4.
+    Iter-RetGen (IRG) shao2023enhancing iteratively does retrieval and LLM generation for 3 iterations.
+  5. 5.
+    FLARE jiang2023flare iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
+  6. 6.
+    Self-RAG (S-RAG) asai2023self uses the LLM to judge for retrieval, generate responses, and self-critique on the responses.
+We use the fine-tuned model based on Llama2-7B from their official repository selfrag_code for trace generation.
+Evaluation datasets.
+We use three commonly used question-answering datasets, NQ kwiatkowski2019natural, HotpotQA yang2018hotpotqa, and TriviaQA joshi2017triviaqa.
+For each dataset, we randomly sampled 1024 queries and reported the average.
+When cache is enabled, to ensure stable results and avoid an initial low hit rate from a cold start, we use 512 queries for warming up the cache and another different 512 queries for evaluation.
+5.2 Experiment Setups
+Hardware setups.
+We evaluated TeleRAG on three hardware environments, Desktop, Server1 and Server2, which are equipped to represent the settings for the desktop and data center use cases.
+The Desktop has the RTX4090, and we use the 3B and 8B models, which are the commonly used model sizes for RAG applications asai2023self; openscholar.
+For Server1 (H100) and Server2 (H200), we use the 8B model.
+We use Server1 for single GPU evaluation and Server2 (8 GPUs) for multi-GPU evaluation.
+Table 2 summarizes the hardware configurations.
+Nprobe and top-$k$.
+A common heuristic for the IVF index is to set nprobe to $4\sqrt{N_{c}}$ zilliz-blog.
+Given our index size of $N_{c}=4096$, we use nprobe = 256 ($4\sqrt{4096}$) by default, unless otherwise specified.
+For retrieval, we use top-$k$ = 3 for the number of documents to return, unless otherwise specified.
+RAG pipeline implementation.
+We implemented the RAG pipelines with the FlashRAG framework jin2024flashrag.
+For IRG, FLARE, and S-RAG, we used the framework’s default implementations.
+For the other pipelines, we reimplemented them using FlashRAG’s APIs.
+Benchmark methodology.
+We follow a benchmarking methodology from SGLang zheng2024_sglang and use GPT-3.5-Turbo gpt3-5-turbo to run through each pipeline and record the input and output text of each step.
+During latency evaluation, we set the LLM’s output length based on the recorded text.
+This way, we ensure a fair latency comparison across different LLM models.
+(a) End-to-end latency speedup with Llama-3.2-3B.
+(b) End-to-end latency speedup with Llama-3-8B.
+Figure 9: End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines and three datasets, with RTX4090 GPU.
+Baseline systems.
+To evaluate the latency of each pipeline, we constructed a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline.
+For LLM generation, we used SGLang zheng2024_sglang, which is a state-of-the-art LLM inference engine.
+For retrieval, we used a popular retrieval library, Faiss douze2024faiss, as the CPU-offloaded baseline.
+Prefetching budget setups.
+Based on the methodology we described in §4.1, we profiled each RAG pipeline with 64 random samples from NQ kwiatkowski2019natural and derived the prefetching budget of each pipeline.
+Since H100 and H200 have the same CPU–GPU bandwidth, we set the same prefetch budget for Server1 and Server2.
+Figure 10: End-to-end throughput of TeleRAG and the CPU-offload baseline across six RAG pipelines on the NQ dataset using Llama-3-8B, evaluated at different batch sizes on a H100 GPU.
+(a) End-to-end throughput on NQ.
+(b) End-to-end throughput on HotpotQA.
+(c) End-to-end throughput on TriviaQA.
+Figure 11: Throughput scaling of TeleRAG across multiple H200 GPUs. Global batch fixed at 128; micro-batch fixed at 4.
+Max GPU memory for retrieval.
+We set a maximum GPU memory limit for prefetching in each configuration.
+For Server1 and Server2, we allocated 12 GB and 24 GB, respectively.
+For Desktop, we allocated up to 10 GB and 3.75 GB for the 3B and 8B models, respectively.
+Cache Setup.
+The memory allocated for retrieval is shared between prefetching and caching.
+In our evaluation, we set the cache proportion to 50% (i.e., cache can use at most half of this memory).
+Since the benefit of caching is negligible on a single GPU (demonstrated in the ablation study), we enabled it only for the multi-GPU experiments on Server2.
+5.3 Evaluation Results
+Single-query latency on RTX4090.
+We evaluated the end-to-end RAG latency for single query on Desktop equipped with a RTX4090, representing the typical local scenario.
+Figure 9 shows the latency reduction of TeleRAG across three datasets and two LLMs (Llama-3.2-3B and Llama-3-8B).
+As Figure 9 shows, TeleRAG consistently outperforms the CPU-offload baseline across all evaluated configurations.
+With Llama-3.2-3B, TeleRAG achieves average speedups of $1.55\times$, $1.54\times$, and $1.49\times$ on NQ, HotpotQA, and TriviaQA, respectively.
+Among the test pipelines, the best speedup of $2.11\times$ is achieved in the Iter-RetGen pipeline on HotpotQA, because this pipeline involves frequent retrieval operations and generally short LLM outputs, which enhances the relative impact of retrieval acceleration.
+Another notable improvement is in the SubQuestion pipeline, where TeleRAG achieves approximately $1.85\times$ speedup across all datasets.
+This pipeline uses LLM-generated sub-questions and performs batched retrievals of 3 to 4 queries in its retrieval stage.
+CPU-based retrieval suffers from limited parallelism in such scenarios, but TeleRAG efficiently utilizes GPU parallelism, significantly enhancing performance.
+When deploying Llama-3-8B, speedups with TeleRAG slightly decrease compared to Llama-3.2-3B, primarily due to increased LLM latency and reduced available memory for prefetching.
+Still, TeleRAG achieves approximately $1.3\times$ speedup across datasets, with a peak improvement of $1.82\times$ for Iter-RetGen pipeline on HotpotQA.
+Achieving these results with only $3.75$ GB of remaining GPU memory (after accounting for Llama-3-8B (16 GB), the embedding model (1 GB), and other miscellaneous tensors) highlights TeleRAG’s robust capability in accelerating RAG inference even under tight GPU memory constraints.
+Figure 12: Latency breakdown for Llama-3-8B on NQ with a H100 GPU in different batch sizes. nprobe is 256.
+Figure 13: Throughput of TeleRAG on the NQ dataset with different numbers of H200 GPUs (with and without cache).
+(a) Llama-3.2-3B with a RTX4090 GPU.
+(b) Llama-3-8B with a H100 GPU.
+Figure 14: Single-query retrieval speedup on NQ with different nprobe values.
+Figure 15: Comparison of end-to-end latency for prefetching and cache-aware schedulers on 4 H200 GPUs. The overhead bars of prefetch schedule only are too short to be visible.
+Pipeline H100 (Llm3-8B) 4090 (Llm3-3B)
+Budget Hit Rate Budget Hit Rate
+HyDE 9 GB $91.95\%$ 7 GB $87.42\%$
+SubQ 8 GB $79.04\%$ 7 GB $76.38\%$
+Iter 5 GB $95.51\%$ 3 GB $84.59\%$
+IRG 4 GB $59.52\%$ 2.5 GB $50.34\%$
+FLARE 6 GB $79.35\%$ 3 GB $56.67\%$
+S-RAG 3 GB $71.29\%$ 1.25 GB $29.96\%$
+Table 3: The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 256.
+Multi-query throughput on H100.
+To illustrate TeleRAG’s performance in batched inference, we evaluated the end-to-end throughput on Server1 (H100) using batch sizes 1, 2, 4, and 8.
+The results of six RAG pipelines with Llama-3-8B are presented in Figure 10.
+[…]
+Multi-GPU throughput.
+We evaluate TeleRAG’s scalability on Server2 (H200), a multi-GPU system.
+Figure 11 reports throughput for a global batch size of 128 with a micro-batch size of 4 across 1–8 GPUs, with both the prefetching and cache schedulers enabled.
+TeleRAG scales well with the number of GPUs: on NQ, compared to single GPU case, the average speedups are $2.0\times$, $3.8\times$, and $6.5\times$ on 2, 4, and 8 GPUs, respectively.
+The mild sub-linear scaling at higher GPU counts is attributed to execution time variance across micro-batches—specifically, a long tail of higher-latency batches that create load imbalance.
+These results still demonstrate that the lookahead retrieval technique in TeleRAG can scale effectively.
+5.4 Analysis and Sensitivity Study
+Latency breakdown.
+We further show the latency breakdown of running RAG pipelines with Llama-3-8B on a single H100 GPU in different batch sizes in Figure 12.
+From Figure 12, we can observe that LLM latency grows sub-linearly with larger batch sizes.
+However, the latency for Faiss retrieval on CPU grows linearly with the batch size, dominating the overall latency when batch size is large.
+These results echo our findings in Figure 10, and show limited scalability of CPU retrieval in serving scenarios.
+In contrast, TeleRAG significantly accelerates across all batch sizes and achieves a higher speedup from $1.4\times$ to $1.8\times$ when the batch size increases from 1 to 8.
+Prefetch budgets and cluster hit rates.
+Table 3 shows the prefetch budgets we set with the profile-guided approach on RTX4090 and H100 for NQ.
+It also presents the average cluster hit rate achieved with this prefetch budget.
+From the table, we can see that it generally achieves a high cluster hit rate ($50\%$}) when TeleRAG has a large prefetching budget.
+For cases where the budget is less than 2 GB, we observe a relatively lower hit rate ($50\%$}), limiting the benefits of reducing the CPU’s search workloads.
+However, as observed from Figure 9, TeleRAG achieves from $1.2\times$ to $1.6\times$ end-to-end speedups for these pipelines, thanks to the combined benefit of reducing CPU workload and utilizing the GPU to perform sorting on similarity distances.
+Retrieval speedups across nprobe.
+Figure 14 shows the retrieval latency reduction on NQ.
+We observe consistent speedups for all nprobe values.
+The greatest speedups are achieved at nprobe 256, with average speedups of $7.21\times$ and $7.41\times$ on RTX4090 and H100, respectively.
+As a larger nprobe value is used, the retrieval performance of TeleRAG becomes constrained by missed clusters on the CPU, given our fixed prefetch budget across varying nprobe values.
+However, RAG inference with higher nprobe will result in longer latency spent on the retrieval and hence, TeleRAG still has a significant latency improvement against the CPU retrieval baseline.
+Ablation on cache.
+Figure 13 illustrates the throughput improvements enabled by caching on the NQ dataset.
+On average, caching boosts throughput of 2%, 12%, 21%, and 18% with 1, 2, 4, and 8 GPUs, respectively.
+The benefit on a single GPU is marginal because only a small cache space is used, and the limited cache capacity struggles to accommodate diverse requests.
+However, as the number of GPUs and overall cache space increases, the cache-aware scheduler effectively assigns micro-batches to the appropriate GPUs to maximize cache overlap, thereby improving prefetching hit rates and yielding significant throughput gains.
+Analysis on scheduling overhead.
+Figure 15 analyzes the benefits and overheads of the prefetching and cache-aware schedulers.
+It reports the end-to-end latency for a global batch of 128 queries with a mini-batch size of 4 on 4 H200 GPUs, comparing configurations that enable both schedulers, only the prefetching scheduler, or neither.
+As shown in Figure 15, the prefetching scheduler incurs only a minimal overhead of approximately 37 ms (thus almost invisible in the figure), while the cache-aware scheduler adds a modest overhead of about 220 ms.
+Overall, both schedulers effectively reduce end-to-end latency in most cases, with minimal additional cost.
+6 Related Work
+Systems for RAG.
+Prior works have accelerated RAG by caching the LLM’s KV cache from retrieved documents jin2024ragcache; lu2024turborag; yao2024cacheblend.
+However, these approaches primarily reduce prefill latency, failing to address the retrieval and decode latencies which often dominate.
+Other techniques, like speculative retrieval zhang2024accelerating or fine-grained pipelining jiang2024piperag, target different RAG paradigms (e.g., per-token retrieval) and do not apply to the widely-used modular pipelines discussed in §2.1.
+Unlike these approaches, TeleRAG is the first to tackle the fundamental system challenges of long retrieval latency and large memory requirements for modular RAG with large-scale datastores.
+Systems for compound LLM applications.
+Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities compound-ai-blog; sigarch; wang2024survey.
+AI Metropolis xie2024ai accelerates LLM-based multi-agent simulations with out-of-order execution.
+RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
+Vector Index.
+A separate line of work focuses on optimizing the vector index itself.
+This includes hardware-specific acceleration (GPU johnson2019billion, FPGA jiang2023co) and hybrid memory-disk systems to scale beyond RAM (DiskANN jayaram2019diskann, SPANN chen2021spann).
+More recent systems scale graph-based indexes beyond GPU memory khan2024bang; zhang2024fast.
+These methods often require significant algorithm modifications or remain bottlenecked by CPU–GPU bandwidth.
+TeleRAG is complementary, optimizing at the system-level by hiding transfer latency within the RAG application’s context, without altering the underlying IVF algorithm.
+7 Conclusion
+In this paper, we introduced TeleRAG, an inference system that improves RAG pipeline latency and throughput while imposing minimal GPU memory requirements.
+TeleRAG achieves significant acceleration by employing lookahead retrieval, which hides CPU–GPU data transfer latency by overlapping it with pre-retrieval generation, and by supporting efficient multi-GPU inference through specialized scheduling.
+Our evaluation shows that TeleRAG significantly improves performance compared to existing state-of-the-art solutions and scales effectively with increased hardware resources.
+Appendix A Background: Inverted File Index (IVF) for Vector Search
 The vector index is a crucial component of RAG applications that retrieves similar items from large datasets of high-dimensional vectors.
@@
 $\text{k-argmin}_{i=0:N}d(x,y_{i}),$
-where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product (Johnson et al., 2019).
+where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product johnson2019billion.
 To improve search efficiency, the inverted file index (IVF) algorithm is widely used for large-scale vector databases due to its simplicity and effectiveness.
 IVF partitions the data store into clusters and restricts searches to the most relevant clusters, reducing the computational cost.
-To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means (Norouzi and Fleet, 2013) and partitions the database into $N_{c}$ clusters:
+To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means norouzi2013cartesian and partitions the database into $N_{c}$ clusters:
 $\{C_{1},C_{2},\ldots,C_{N_{c}}\}=k\text{-means}(Y,N_{c}),$
@@
 In this way, IVF reduces search space and accelerates the retrieval process.
-Here, the hyperparameter $L$ from the first step is also referred as nprobe (Douze et al., 2024).
+Here, the hyperparameter $L$ from the first step is also referred as nprobe douze2024faiss.
 When nprobe is larger, the IVF retrieval will be more accurate as it search more clusters.
 However, the retrieve latency is longer as more computation and data are needed.
-Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs.
-Open-source libraries offer efficient GPU implementations (Johnson et al., 2019; Rapidsai, 2022).
-However, IVF does not reduce the memory requirement for the index since the data for all clusters must be stored.
-As a result, IVF still requires a substantial GPU memory footprint, which becomes a bottleneck when GPU memory capacity is the constraint of RAG systems.
-Moreover, recent work reported that increasing the size of the data store positively affects RAG application performance (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n.
-d.]), exacerbating this issue.
-2.3. Use Cases and Motivation
-In this paper, we aim to achieve low-latency RAG inference without driving up GPU memory consumption, which is relevant for both local and data center settings.
-This goal is critical for efficiently serving a wide range of real-world applications where users expect near-instantaneous responses, thus imposing strict latency SLOs.
-Examples include customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024; AWS, [n.
-d.]), financial analysis (Malec, 2024; MyScale, 2024), autonomous driving (Cai et al., 2024; Yuan et al., 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024).
-On the one hand, there is significant demand for RAG applications that involve user-specific private data in local settings (Wang and Chau, 2024; Lam et al., 2023; Chung et al., 2023; Wutschitz et al., 2023; Ghodratnama and Zakershahrak, 2023).
-In such scenarios, the available GPU resources are usually limited, and the system serves a single request at a time.
-Therefore, CPU offloading is necessary when using a large vector index, which introduces substantial latency as we will show in the next section.
-On the other hand, for large-scale RAG services in data centers (per, [n.
-d.]; gen, [n.
-d.]; Asai et al., 2024a), the goal is to serve many concurrent user requests cost-effectively, necessitating high system throughput.
-However, maximizing LLM serving throughput is often bottlenecked by the GPU memory available for the KV cache, which dictates the maximum possible batch size (Kwon et al., 2023).
-If a large retrieval datastore is loaded onto GPUs, it consumes memory that could otherwise be used for a larger KV cache, thereby limiting throughput.
-Hence, falling back to slow CPU retrieval to save GPU memory is a common option (Jiang et al., 2025), but this is not desirable for latency-sensitive applications due to latency overhead.
-In both cases, a straightforward way to reduce inference latency is scaling up GPU resources to allow both the LLM and datastore to reside in GPU memory, which is usually cost-prohibitive.
-Therefore, in this paper, we aim to design an efficient RAG inference system that can satisfy latency SLOs without requiring a significant amount of GPU memory.
-3. Analyzing Latency of RAG Pipelines
-In this section, we analyze state-of-the-art RAG applications and identify their underlying systems challenges in achieving low latency.
-To conduct the analysis, we construct a 61 GB vector index with the Faiss library (Douze et al., 2024), and set the IVF clusters to 4096 following common practice (Asai et al., 2023).
-We use Llama-3-8B (AI@Meta, 2024) as the LLM, and run our analysis on an RTX 4090 (24 GB memory) and H100 (80 GB memory).
-§5.2 provides more details about our experimental setups.
-3.1. End-to-end Latency of RAG Pipelines
-(a) Batch = 1.
-(b) Batch = 4.
-Figure 3. Latency breakdown of six RAG pipelines on NQ dataset (Kwiatkowski et al., 2019) with one H100 GPU. nprobe is 256.
-Figure 4.
-The breakdown of memory consumption at GPU and CPU for two different strategies: CPU offloading and GPU-based retrieval.
-The dotted line indicates the memory capacity of a RTX4090 GPU, which is a common used GPU for local deployment.
-We first analyze the end-to-end latency of six representative RAG pipelines (see §5.1 for details) in two scenarios: (1) the LLM runs on the GPU, while retrieval is offloaded to the CPU, minimizing GPU memory usage, and (2) both the LLM and the vector index reside in GPU memory, enabling GPU-based retrieval for lower latency.
-If the GPU’s memory is insufficient, CPU offloading is necessary.
-We set the nprobe to 256, a commonly used setting under this index’s configuration (see §5.2 for details), to measure retrieval latency.
-Figure 3 shows the breakdown of end-to-end latency into LLM inference part and retrieval part, based on 1024 randomly sampled queries from the NQ dataset (Kwiatkowski et al., 2019).
-We observe that, in the CPU-based retrieval baseline (first scenario), the retrieval phase dominates the latency, consuming 41.1% and 60.5% of total latency for batch sizes 1 and 4, respectively.
-In contrast, GPU-accelerated retrieval in the second scenario dramatically reduces this bottleneck, accounting for only 10.5% and 28.3% of the latency in comparable configurations.
-On average, GPU retrieval is $5.96\times$ (batch size 1) or $3.87\times$ (batch size 4) faster than CPU retrieval, reducing overall latency by approximately $1.5\times$ or $1.8\times$, respectively.
-Thus, accelerating retrieval on the GPU is crucial for improving end-to-end latency.
-However, GPU acceleration comes with a significant memory cost.
-Figure 4 shows the memory requirements for both GPU and CPU.
-As Figure 4 shows, putting both LLM weights and the retrieval index on the GPU requires around 77 GB of memory, which exceeds the capacity of consumer GPUs like the RTX 4090 with 24 GB.
-Thus, GPU acceleration on retrieval is often not feasible when running on a lower-end GPU or indexing with a large datastore.
-Even on GPUs capable of holding the entire index in memory (e.g., H100), the index’s memory footprint limits throughput in data center serving scenario.
-Batched inference maximizes throughput, but it is bottlenecked by the KV cache capacity for LLM serving (Kwon et al., 2023).
-For the case of the example above, storing the index on the GPU leaves minimal memory for the KV cache (e.g., ~3 GB, or ~20k tokens worth), restricting batch size, while offloading the index to the CPU significantly increases the available memory (e.g., to ¿500k tokens worth).
-This issue is exacerbated since RAG typically involves long contexts (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024) and there is a trend of growing model and index sizes (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n.
-d.]), increasing memory pressure and necessitating offloading.
-However, CPU offloading is suboptimal for latency-sensitive applications as discussed above.
-Therefore, in the rest of this section, we try to answer the following question: Is it possible to achieve the latency of GPU-based retrieval while using much less GPU memory?
-Figure 5. Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
-3.2. Opportunities of GPU-accelerated Retrieval with Limited Memory
-A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on-demand for each query, leveraging the IVF index structure that narrows the search space (§2.2).
-While this method enables faster searching on the GPU, data fetching becomes the bottleneck.
-In Figure 5, we compare the latency of the on-demand fetching system against CPU retrieval on an RTX 4090 GPU.
-We show three different nprobe values that determine the amount of data fetched.
-Overall, fetch time dominates latency due to the limited PCIe bandwidth between the CPU and GPU (32 GB/s).
-Although the GPU search is substantially faster than the CPU counterpart, the fetch overhead results in a higher end-to-end latency, which is about $3\%$ slower on average across nprobe values.
-Thus, to realize any meaningful speedup with this approach, the data fetching latency must be effectively hidden.
-To hide data fetch costs, CPU-to-GPU transfer must precede retrieval, which requires predicting which data will be accessed.
-Fortunately, as we show next, the modern RAG pipelines offer a valuable hint as a query at the previous step.
-Dataset HyDE SubQ Iter IRG FLARE S-RAG
-(Gao et al., 2023a) (Liu, 2022) (Liu, 2022) (Shao et al., 2023) (Jiang et al., 2023c) (Asai et al., 2023)
-NQ (Kwiatkowski et al., 2019) 73.1% 63.2% 91.5% 83.8% 79.1% 100.0%
-HotpotQA (Yang et al., 2018) 75.3% 62.5% 89.6% 89.4% 80.2% 100.0%
-TriviaQA (Joshi et al., 2017) 73.1% 61.6% 86.2% 86.1% 81.7% 100.0%
-Table 1.
-IVF cluster overlapping rate between the input and output of the pre-retrieval generation.
-The nprobe is set to 512.
-Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
-3.3. Overlapping of IVF Clusters
-While the exact data to be retrieved is only known after the pre-retrieval generation is done, we observe that there are high similarities of the IVF clusters assignments between the queries at different stages.
-Similarity of queries at different stages.
-During the pre-retrieval process of RAG pipelines (e.g., query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022)), an LLM call is issued to refine an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$, which is then used for retrieval.
-This process often rewrites the query into a different format (Gao et al., 2023a) or simplifies its complexities (Liu, 2022), which intuitively preserves the query’s core semantic content.
-Hence, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar.
-This similarity, in turn, suggests that the IVF clusters to which these queries would be assigned will overlap.
-Therefore, $q_{\text{in}}$ can serve as a valuable hint for predicting subsequent memory accesses.
-Prediction coverage.
-To verify this hypothesis, we evaluate the average cluster coverage rate between prefetched clusters and clusters used for retrieval in three popular question-answer datasets (NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017)) and six RAG pipelines.
-Table 1 shows the coverage when we prefetch $256$ clusters.
-From the table, we observe that IVF cluster overlap rates are consistently high across a range of datasets and pipelines.
-For instance, even the lowest reported values remain above 61.6% (for SubQ).
-Opportunity.
-This data shows an opportunity for predicting required clusters, which can make it possible to hide data transfer overhead during LLM generation.
-In this paper, we aim to leverage this observation to accelerate the inference latency for RAG.
-Figure 6.
-(a) The overview of lookahead retrieval compared against CPU-offloaded retrieval, and (b) the system design of TeleRAG.
-After identifying clusters ($C_{\text{in}}$) for the initial query $q_{\text{in}}$, $C_{\text{in}}$ is transferred to the GPU while concurrently generating $q_{\text{out}}$.
-At retrieval time, the GPU searches prefetched clusters while the CPU handles remaining ones, before results are merged.
-4. Design of TeleRAG
-Based on the high IVF cluster overlapping observation presented in §3.3, we propose TeleRAG, an efficient RAG inference system that utilizes lookahead retrieval to prefetch a minimal set of likely IVF clusters to GPU, and accelerate the retrieval process.
-In this section, we detail the design of lookahead retrieval and implementation of TeleRAG.
-4.1. Overview
-Figure 6 presents the overview of TeleRAG with lookahead retrieval.
-In Figure 6, $q_{\text{in}}$ denotes the initial input to the pre-retrieval stage and $q_{\text{out}}$ denotes its output, i.e., the query given to the subsequent retrieval stage.
-We highlight the IVF clusters corresponding to $q_{\text{in}}$ and $q_{\text{out}}$ with a red background ($C_{\text{in}}$), and a blue background ($C_{\text{out}}$), respectively.
-Due to the semantic similarity of $q_{\text{in}}$ and $q_{\text{out}}$, there is significant overlap between $C_{\text{in}}$ and $C_{\text{out}}$, marked with a purple background ($C_{\text{overlap}}$).
-Given this, the proposed lookahead retrieval technique in TeleRAG operates in three key steps:
-  1. (1)
-    Pre-retrieval & data transfer: During LLM generation, identify $C_{\text{in}}$ and transfer corresponding data to GPU memory, prioritizing clusters closest to $q_{\text{in}}$.
-This leverages GPU DMA for concurrent transfer of clusters data and compute of LLM generation.
-  2. (2)
-    GPU similarity search: After identifying $C_{\text{out}}$, the GPU efficiently performs a similarity search on the predicted clusters ($C_{\text{overlap}}$) already in its memory.
-  3. (3)
-    CPU similarity search: Concurrently with the GPU search, the CPU processes the similarity search for the missed clusters ($C_{\text{miss}}$) that were not prefetched to the GPU, reducing the CPU’s overall workload.
-In summary, our lookahead retrieval speeds up retrieval by overlapping data prefetching to the GPU with LLM generation and performing concurrent similarity searches on both the GPU (for predicted clusters) and CPU (for the remainder), significantly reducing the CPU’s computational workload.
-4.2. Finding the Optimal Prefetch Amount
+Appendix B Finding Prefetching Amount
 A key challenge in TeleRAG is to balance the benefit of reducing retrieval latency by prefetching data against the overhead of CPU-GPU transfers.
@@
 Since prefetching time $t_{p}$ is proportional to $b_{p}$, we express $t_{1}$ as a piecewise function:
-(1) $t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$
+$t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$ (1)
 From Eq. 1, if we prefetch fewer bytes than can be transferred during LLM generation, $t_{1}$ is effectively just $t_{\text{LLM}}$ because the transfer overlaps completely with LLM execution.
-As shown in §3.1, GPU retrieval ($t_{g}$) is generally much faster than CPU retrieval ($t_{c}$).
+As shown in the analysis section, GPU retrieval ($t_{g}$) is generally much faster than CPU retrieval ($t_{c}$).
 Thus, we assume $t_{c}\gg t_{g}$.
-As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed (Jiang et al., 2023a):
-(2) $t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$
+As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed jiang2023co:
+$t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$ (2)
 where $r_{\mathrm{miss}}$ is the miss rate (percentage of IVF clusters are not caught on GPU), $n_{\mathrm{probe}}$ is the total number of clusters to search, and $t_{cc}$ is the CPU time to search a single cluster.
@@
 Hence,
-    (3) $b_{p}^{*}=B\cdot t_{\text{LLM}}.$
+    $b_{p}^{*}=B\cdot t_{\text{LLM}}.$ (3)
   • Case 2: $b_{p}^{*}>B\cdot t_{\text{LLM}}$.
@@
 At the minimum point, we have:
-    (4) $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$
+    $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$ (4)
     From this, we obtain:
-    (5) $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$
+    $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$ (5)
     where $\Delta r_{\mathrm{miss}}$ is the decrement of the miss rate for this round.
@@
 5.
-However, under typical CPU-GPU bandwidth (e.g., 55 GB/s on PCIe 5), the time spent loading additional clusters often outweighs any retrieval latency reduction from lowering $r_{\mathrm{miss}}$.
+However, under typical CPU-GPU bandwidth (e.g., 64 GB/s on PCIe 5), the time spent loading additional clusters often outweighs any retrieval latency reduction from lowering $r_{\mathrm{miss}}$.
 Consequently, the second scenario in Eq.
@@
 Therefore, on current hardware, prefetching exactly until LLM execution ends is generally the most effective choice.
-Profiling-guided approach.
-Although $b_{p}^{*}$ can be derived from the analysis above, it depends on knowing $t_{\mathrm{LLM}}$ for each query, which cannot be obtained ahead of time.
-To address this, we use a profiling-guided approach, leveraging the observation that, despite differences in query content, the output length (and thus generation time) for a RAG pipeline often remains similar across most queries.
-Accordingly, for each RAG pipeline, we can measure $t_{\mathrm{LLM}}$ on a calibration set containing n queries, and get estimated $\hat{b}_{p}^{*}=B\cdot\overline{t}_{\mathrm{LLM}}$, where $\overline{t}_{\mathrm{LLM}}=\mathrm{mean}\{t_{\mathrm{LLM},1},t_{\mathrm{LLM},2},\dots,t_{\mathrm{LLM},n}\}.$ This estimated $\hat{b}_{p}^{*}$ can then be used for incoming queries, ensuring a near-optimal prefetch amount.
-4.3. Design Details
-Sorting Optimization.
-TeleRAG accelerates the final $k$-argmin sorting step of IVF search by leveraging GPU.
-Because each cluster may contain thousands of points, sorting is particularly well-suited to GPU parallelism (Arkhipov et al., 2017).
-To enable this, TeleRAG transfers the scalar distance values computed on the CPU for $C_{\text{miss}}$ to the GPU.
-Unlike full vector data, this transfer incurs negligible overhead.
-The GPU then performs a global sort over the combined distances from $C_{\text{overlap}}$ and $C_{\text{miss}}$, delivering end-to-end speedups than the CPU-only retrieval systems.
-Prefetch Target.
-TeleRAG’s prefetching mechanism targets a specific number of bytes ($b_{p}$) rather than the number of clusters to ensure more predictable performance.
-Since IVF cluster sizes are often highly uneven, targeting a fixed cluster count can lead to unstable data loading times, whereas a byte limit provides a clearer upper bound on the transfer duration based on bandwidth.
-When selecting clusters to meet the prefetch byte budget ($b_{p}$), the system adds whole clusters sequentially based on proximity to the query.
-If adding the next closest cluster would exceed the budget, it is omitted entirely rather than being partially loaded, ensuring a clean division where each cluster is processed exclusively by either GPU or CPU.
-Multi-Round Support.
-For multi-round RAG involving the same input query, the system performs a full prefetch (up to the budget) only in the first round, leveraging high cluster similarity across rounds.
-In subsequent rounds, it incrementally fetches only the additional required clusters that were not loaded previously, optimizing data transfer while respecting the memory budget.
-Batch Support.
-For batch inference, a challenge for lookahead retrieval is that each query requires a different set of IVF clusters.
-In our design, we fix the prefetching budget to the amount we got from §4.2 irrespective of batch size, and equally distribute the budget to each query in a batch.
-As a result, although the average cluster hit rate per request will decrease for larger batch sizes, we can ensure each query’s retrieval search time is equally accelerated, forming a balanced batched retrieval.
-Multi-GPU support.
-For multi-GPU systems, we perform a greedy search based on embedding similarities of queries to group similar queries into mini-batches, and distribute the mini-batches among GPUs.
-The greedy search is efficient and incurs negligible overhead: In our profiling for up to the batch size of 256, the latency to perform greedy search is less than $0.1$ s, which is minimal compared to the end-to-end latency of each batch.
-4.4. Implementation Details
-TeleRAG is implemented in Python and also leverages PyTorch’s (Ansel et al., 2024) operator ecosystem for efficient computation.
-The datastore index is initially constructed using FAISS (Douze et al., 2024), and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors.
+Appendix C Implementation Details
+TeleRAG is implemented in Python and also leverages PyTorch’s pytorch_2 operator ecosystem for efficient computation.
+The datastore index is initially constructed using FAISS douze2024faiss, and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors.
 At runtime, cluster data is loaded into a contiguous pinned memory region on the CPU, enabling non-blocking memory copies to the GPU.
 A fixed-size contiguous buffer on the GPU is allocated based on the user’s configuration or GPU memory capacity during runtime.
-To enable concurrent CPU-GPU data transferring and LLM generation, we utilize the _copy(non_blocking=True) API in PyTorch and use separate CUDA streams for prefetching and LLM.
-In this way, the data copy operations will not block the GPU to perform computation, and thus the TeleRAG can do prefetching in parallel to the pre-retrieval LLM generation.
-To implement the index search with GPU-CPU cooperation, for the GPU part, we use a single matrix-vector multiplication that computes distances for all prefetched vectors; for the CPU part, we utilize multithreading in Python to parallelize similarity searches across clusters.
-We then move the distances computed from CPU to GPU, merge with distances on GPU and perform global sorting on GPU.
+We use separate processes for the schedulers and services in TeleRAG to avoid bottlenecks in Python global interpreter lock.
+For the communication between the scheduler and the services, we also leverages the ZeroMQ zeromq as the communication framework.
+The message is encoded and decoded with Pickle in Python.
+For the LLM sIervice, we use SGLang zheng2024_sglang as the LLM backend.
+For the retrieval service, to implement the index search with GPU-CPU cooperation, on GPU, we use a single matrix-vector multiplication that computes distances for all prefetched vectors; on CPU, we utilize multithreading in Python to parallelize similarity searches across clusters.
+We then move the distances computed from CPU to GPU, merge with results on GPU and perform global sorting on GPU.
+To monitor the hotness of cached clusters better, we use a decay-based approach to track the hotness of cached clusters.
+Every cached cluster has a hotness value.
+Upon fetching to the GPU, the cluster is essentially cached and will be assigned with a initial value $h_{0}=h_{\text{init}}$.
+Then, when a new round of hotness update starts, the hotness will be divided with a decay factor $d$, and if the cluster is used in the previous round, an hotness update value $h_{\text{inc}}$ will be added the hotness.
+Therefore, the transition of hotness from the round $r$ to round $r+1$ can be given as:
+[…]
+In this way, the process can be parallelized easily on GPU, minimize the overhead of hotness tracking in terms of both the time and computation.
+We also set a cache fraction, which means if the cache exceeds the fraction, we need to evict some clusters.
+The cache fraction is set to $0.5$ by default.
+To make evaluation result stable and reproducible, we do cache eviction for the excessive clusters and consolidate the GPU memory after serving each batch.
+This makes sure that we will not over-evaluate the performance of the caching system.
+Both schedulers use greedy search to minimize overhead.
+The prefetching scheduler groups similar queries into mini-batches, and feeds them into the cache-aware scheduler to assign mini-batches to GPUs.
+The cache-aware scheduler will get the number of overlapped clusters on each GPU for each mini-batch, and assign the mini-batch to the GPU with the most overlapped clusters.
+Appendix D Detailed Index Configurations
 Specification Value
-Dataset wiki_dpr (Karpukhin et al., 2020)
+Dataset wiki_dpr karpukhin2020dense
 Dataset size 2.1 billion tokens
@@
 #\# of IVF cluster 4096
-Embed model Contriever (Izacard et al., 2021)
+Embed model Contriever izacard2021unsupervised
 Embed dimension 768
@@
 Index size 61 GB
-Table 2. Detailed configurations of our retrieval index.
-5. Evaluation
-We conduct extensive experiments to evaluate the effectiveness of TeleRAG.
-In this section, we describe the necessary details on how we set up the evaluations, present experimental results, and provide in-depth analysis and discussions.
-5.1. Evaluation Datasets and RAG Models
-Datastore.
-We build a datastore based on the wiki_dpr dataset (Karpukhin et al., 2020), a popular dataset that contains 2.1 billion tokens from Wikipedia.
-Following previous works (Asai et al., 2023; Karpukhin et al., 2020; Min et al., 2023), we chunk the passages by every 100 tokens, and use Contriever (Izacard et al., 2021) to generate an embedding for each chunk.
-The embeddings have a hidden dimension of 768.
-Vector index.
-For the baseline retrieval, we build an IVF vector index using Faiss (Douze et al., 2024) on the datastore.
-As described in §4.4, we convert the Faiss index to a customized index in PyTorch for TeleRAG.
-See Table 2 for the detailed configurations of our vector index and datastore.
-LLMs. We evaluate TeleRAG on the Llama model family (Touvron et al., 2023) in three different sizes (Llama-3.2-3B, Llama-3-8B, Llama-2-13B) to represent different use cases.
-Setup Desktop Server
-CPU Threadripper 5975 EPYC 9554
-CPU memory size 512 GB 1.5 TB
-GPU RTX4090 (NVIDIA, 2024a) H100 (NVIDIA, 2024b)
-GPU memory size 24 GB 80 GB
-CPU-GPU Bus PCIe 4 PCIe 5
-Bandwidth n 32 (24) GB/s n n 64 (51) GB/s n
-Table 3. Hardware specifications for our setups. In bandwidth, the number in the parentheses is the actual bandwidth we measured from our system.
-RAG pipelines.
-We evaluate TeleRAG with six popular RAG pipelines.
-Figure 7 shows the overview of each pipeline.
-Note that although some pipelines lack pre-retrieval generation, the post-retrieval generation serves similar functionality for the retrieval of the next iteration.
-Below are brief descriptions of the evaluated RAG pipelines.
-  1. (1)
-    HyDE (Gao et al., 2023a) prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
-  2. (2)
-    SubQestion (SubQ) [^3] Implemented in LlamaIndex as SubQuestionQueryEngine. (Liu, 2022) prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
-  3. (3)
-    Iterative (Iter) [^4] Implemented in LlamaIndex as MultiStepQueryEngine.
-(Liu, 2022) prompts LLM to generate narrower questions first and iteratively refine them based on previous answers.
-At the end of each iteration, it prompts LLM to judge if the answer is good enough.
-  4. (4)
-    Iter-RetGen (IRG) (Shao et al., 2023) iteratively do retrieval and LLM generation for 3 iterations.
-  5. (5)
-    FLARE (Jiang et al., 2023c) iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
-  6. (6)
-    Self-RAG (S-RAG) [^5] We evaluate on the short-form version which only has one iteration.
-(Asai et al., 2023) uses the LLM to judge for retrieval, generate responses, and self-critique on the responses.
-We use fine-tuned models based on Llama2-7B and Llama2-13B from their official repository (Asai et al., 2024b).
-Figure 7. Overview of six RAG pipelines that we evaluate.
-Evaluation Datasets.
-We use three commonly used question-answering datasets, NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017).
-For each dataset, we randomly sample 1024 queries and report the average unless otherwise specified.
-5.2. Experiment Setups
-Hardware setups.
-We evaluate TeleRAG on two hardware environments, Desktop and Server, which are equipped to represent the settings for the desktop and data center use cases.
-The Desktop has the RTX4090 (24 GB memory), and we use 3B and 8B models.
-The Server has the H100 (80 GB memory), and we use 8B and 13B models.
-These sizes represent common model sizes for RAG applications (Asai et al., 2023, 2024a).
-We do not evaluate the 13B model on Desktop as its model size (26 GB) exceeds RTX4090’s memory capacity.
-Table 3 summarizes the hardware configurations.
-Nprobe and top-$k$.
-For the IVF index, nprobe is a hyperparameter that controls the trade-off between search efficiency and retrieval quality.
-A common heuristic is to set nprobe to $4\sqrt{N_{c}}$ (Zilliz, 2020).
-Given our index size of $N_{c}=4096$, we use nprobe = 256 ($4\sqrt{4096}$) by default, unless otherwise specified.
-For retrieval, top-$k$ denotes the number of most relevant documents to return.
-We use top-$k$ = 3, which is a standard choice in RAG.
-RAG pipeline implementation.
-We implement the RAG pipelines with the FlashRAG framework (Jin et al., 2024b).
-For IRG, FLARE, and S-RAG, we use the framework’s default implementations.
-For the other pipelines, we reimplement them using FlashRAG’s APIs.
-Benchmark methodology.
-We follow a benchmarking methodology from SGLang (Zheng et al., 2024) and use GPT-3.5-Turbo (OpenAI, [n.
-d.]) to run through each pipeline and record the input and output text of each step.
-During latency evaluation, we set the LLM’s output length based on the recorded text.
-This way, we ensure a fair latency comparison across different LLM models.
-(a) End-to-end latency speedup with Llama-3.2-3B.
-(b) End-to-end latency speedup with Llama-3-8B.
-Figure 8. End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines and three datasets, with RTX4090 GPU. nprobe is 256.
-Baseline systems.
-To evaluate the latency of each pipeline, we construct a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline.
-For LLM generation, we use SGLang (Zheng et al., 2024), which is a state-of-the-art LLM inference engine.
-For retrieval, we use a popular retrieval library, Faiss (Douze et al., 2024), as the CPU-offloaded baseline.
-[^6] This baseline resembles frameworks like LlamaIndex and LangChain, which primarily function as wrappers around LLM serving and retrieval libraries.
-We build a system from the ground up to avoid the overhead introduced by their complex layers designed for advanced functionality.
-Prefetching budget setups.
-Based on the methodology we described in §4.2, we profile each RAG pipeline with a small amount of 64 random samples from NQ (Kwiatkowski et al., 2019) and derive the prefetching budget of each pipeline.
-(a) End-to-end throughput with six pipelines and different batch sizes using Llama-3-8B.
-(b) End-to-end throughput with six pipelines and different batch sizes using Llama-2-13B.
-Figure 9.
-End-to-end throughput of TeleRAG and the baseline across six RAG pipelines on the NQ dataset using Llama-3-8B and Llama-2-13B at different batch sizes on a H100 GPU.
-The nprobe is set to 256, and the $x$-axis represents the batch size.
-Max prefetching memory.
-We set a maximum GPU memory limit for prefetching in each configuration.
-For Server, we allocate up to 16 GB.
-For Desktop, we allocate up to 10 GB and 3.75 GB for the 3B and 8B models, respectively.
-These settings demonstrate that TeleRAG can efficiently operate using only a small fraction (up to 40% for RTX4090 and 20% for H100) of total GPU memory.
-5.3. Evaluation Results
-Single-query latency on RTX4090.
-We evaluate the end-to-end RAG latency for single queries on Desktop equipped with RTX4090, representing a typical local scenario.
-Figure 8 shows the latency reduction of TeleRAG across three datasets and two LLMs (Llama-3.2-3B and Llama-3-8B).
-As shown in Figure 8, TeleRAG consistently outperforms the baseline across all tested configurations.
-With Llama-3.2-3B, TeleRAG achieves average speedups of $1.55\times$, $1.54\times$, and $1.49\times$ on NQ, HotpotQA, and TriviaQA, respectively.
-The best speedup of $2.11\times$ is achieved in the Iter-RetGen pipeline on HotpotQA, because this pipeline involves frequent retrieval operations and generally short LLM outputs, which enhances the relative impact of retrieval acceleration.
-Another notable improvement is in the SubQuestion pipeline, where TeleRAG achieves approximately $1.85\times$ speedup across all datasets.
-This pipeline uses LLM-generated sub-questions and performs batched retrievals of 3 to 4 queries simultaneously.
-CPU-based retrieval suffers from limited parallelism in such scenarios, but TeleRAG efficiently utilizes GPU parallelism, significantly enhancing performance.
-When deploying Llama-3-8B, speedups with TeleRAG slightly decrease compared to Llama-3.2-3B, primarily due to increased LLM latency and reduced available memory for prefetching.
-Still, TeleRAG achieves approximately $1.3\times$ speedup across datasets, with a peak improvement of $1.82\times$ for Iter-RetGen pipeline on HotpotQA.
-Achieving these results with only $3.75$ GB of remaining GPU memory (after accounting for Llama-3-8B (16 GB), the embedding model (1 GB), and other miscellaneous tensors) highlights TeleRAG’s robust capability in accelerating RAG inference even under tight GPU memory constraints.
-Figure 10. Simulated throughput [^8] Results for 1 GPU are real, not simulated.of TeleRAG on NQ dataset with different number of H100 GPUs.
-Multi-query throughput on H100.
-To evaluate the TeleRAG’s performance in batched inference, we evaluate the end-to-end throughput on Server (H100) in batch sizes 1, 2, 4, and 8.
-We show the results of six evaluated RAG pipelines with Llama-3-8B and Llama-2-13B in Figure 9.
-[…]
-Figure 11. Latency breakdown for Llama-3-8B on NQ with a H100 GPU in different batch sizes. nprobe is 256.
-Figure 12. Comparison of end‑to‑end retrieval latency for two batching strategies: naive mini‑batching and similarity‑aware greedy grouping. The Greedy Group bars are too short to be visible.
-Multi-GPU Throughput.
-We also evaluate TeleRAG in a multi-GPU system to further show the scalability.
-For the evaluation, we construct a simulation framework around our single-GPU implementation.
-We first define a global batch size that the whole system will process at once, and then we construct mini-batches to distribute among each GPU.
-To form a mini-batch, we perform a greedy search based on the embedding similarity of queries, grouping similar queries into the same mini-batch to maximize the cluster hit rate.
-We present the simulated throughput of TeleRAG with a global batch size of 128 and mini-batch size of 4 with 1 to 8 GPUs.
-As shown in Figure 8, TeleRAG’s throughput scales well with the number of GPUs.
-When compared to 1 GPU results, the averaged throughput enhancements of TeleRAG are $1.7\times$, $3.1\times$, and $5.4\times$ with $2$, 44, and $8$ GPUs, respectively.
-5.4. Analysis and Sensitivity Study
-Latency breakdown.
-We further show the latency breakdown of running RAG pipelines with Llama-3-8B and Llama-2-13B on a single H100 GPU in different batch sizes in Figure 11.
-From Figure 11, we can observe that LLM latency grows sub-linearly with larger batch sizes.
-However, the latency for Faiss retrieval on CPU grows linearly with the batch size, dominating the overall latency when batch size is large.
-These results echo our findings in Figure 9, and show limited scalability of CPU retrieval in serving scenario.
-In contrast, TeleRAG significantly accelerates across all batch sizes and achieves a higher speedup from $1.4\times$ to $1.8\times$ when the batch size increases from 1 to 8.
-Ablation on mini-batch strategies.
-In our multi-GPU results, we implement a greedy grouping strategy for mini-batching based on queries’ similarity in a global batch.
-Here, we examine the benefits and overhead of this strategy in Figure 12, where we show the end-to-end retrieval latency for mini-batch size $4$ with four GPUs and a global batch of 128 queries.
-The grouping overhead is minimal and the overall retrieval latency of greedy grouping mini-batching consistently outperform naive mini-batching across all pipelines.
-(a) Llama-3.2-3B with a RTX4090 GPU.
-(b) Llama-3-8B with a H100 GPU.
-Figure 13. Single-query retrieval speedup on NQ with different nprobe values.
-Retrieval speedups across nprobe.
-Figure 13 shows the retrieval latency reduction on NQ.
-We observe consistent speedups for all nprobe values.
-The greatest speedups are achieved at nprobe 256, with average speedups of $7.21\times$ and $7.41\times$ on RTX4090 and H100, respectively.
-As a larger nprobe value is used, the retrieval performance of TeleRAG becomes constrained by missed clusters on the CPU, given our fixed prefetch budget across varying nprobe values.
-However, RAG inference with higher nprobe will result in longer latency spent on the retrieval and hence, TeleRAG still has a significant latency improvement against the CPU retrieval baseline.
-Pipeline H100 (Llm3-8B) 4090 (Llm3-3B)
-Budget Hit Rate Budget Hit Rate
-HyDE 9 GB $91.95\%$ 7 GB $87.42\%$
-SubQ 8 GB $79.04\%$ 7 GB $76.38\%$
-Iter 5 GB $95.51\%$ 3 GB $84.59\%$
-IRG 4 GB $59.52\%$ 2.5 GB $50.34\%$
-FLARE 6 GB $79.35\%$ 3 GB $56.67\%$
-S-RAG 3 GB $71.29\%$ 1.25 GB $29.96\%$
-Table 4. The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 256.
-Prefetch budgets and cluster hit rates.
-Table 4 shows the prefetch budgets we set with the profile-guided approach on RTX4090 and H100 for NQ.
-It also presents the average cluster hit rate achieved with this prefetch budget.
-From the table, we can see that it generally achieves a high cluster hit rate (¿5050%) when TeleRAG has a large prefetching budget.
-For cases where the budget is less than 2 GB, we observe a relatively lower hit rate (¡5050%), limiting the benefits of reducing the CPU’s search workloads.
-However, as observed from Figure 8, TeleRAG achieves from $1.2\times$ to $1.6\times$ end-to-end speedups for these pipelines, thanks to the combined benefit of reducing CPU workload and utilizing the GPU to perform sorting on similarity distances.
-6. Related Work
-6.1. Systems for RAG
-RAGCache (Jin et al., 2024a) proposes a caching system that stores KV caches from datastore chunks in an order-aware manner to improve the time to the first token.
-This approach only reduces prefill latency, leaving retrieval and decode times unaffected, despite the fact they typically dominate total latency (Agrawal et al., 2024).
-Moreover, it assumes repeated use of the same document across multiple requests, limiting its scalability over large data stores.
-Similarly, TurboRAG (Lu et al., 2024) precomputes KV caches from the data store, but it also only optimizes prefill latency.
-CacheBlend introduces a selective KV-cache fusion technique for RAG to reuse pre-computed caches (Yao et al., 2024).
-RaLMSpec (Zhang et al., 2024b) proposes speculative retrieval and batched verification, Chameleon (Jiang et al., 2023d) proposes a CPU-GPU-FPGA heterogeneous architecture for accelerating the retrieval process, and PipeRAG (Jiang et al., 2024) is an algorithm-system co-design technique to overlap the retrieval and generation by modifying the algorithms.
-However, these works focus on the paradigm of RAG that retrieves documents once every few tokens and do not apply to modular RAG applications that are widely used now, as discussed in §2.1.
-Speculative RAG (Wang et al., 2024b) introduces drafting by smaller LLMs to reduce RAG latency, but it does not target the retrieval latency.
-APIServe (InferCept) (Abhyankar et al., 2024) proposes a novel KV cache management strategy that can support the interception of LLM generation by other workloads, including retrieval.
-However, this work again does not focus on optimizing retrieval latency.
-EdgeRAG (Seemakhupt et al., 2024) reduces the memory requirement of retrieval by re-generating and caching document embeddings at runtime, targeting extremely resource-constrained environments.
-RAGO (Jiang et al., 2025) introduces an abstraction for RAG to automatically pick task placement, resource allocation, and batching policies.
-Unlike all these prior works, we tackle system challenges of long retrieval latency and large memory requirement for modular RAG pipelines with large-scale datastores.
-6.2. Systems for Compound LLM Applications
-Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities (Zaharia et al., 2024; Berger and Zorn, 2024; Wang et al., 2024a).
-LLMCompiler (Kim et al., 2023a) is a framework that optimizes the execution of multiple functions in large language models by enabling parallel function calling.
-AI Metropolis (Xie et al., 2024) accelerates LLM-based multi-agent simulations with out-of-order execution.
-RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
-6.3. Vector Index
-Vector index is a key component of RAG (Briggs et al., [n.
-d.]; Gao et al., 2023b), and many works have been proposed to improve their efficiency.
-ScaNN proposes an anisotropic quantization method for better efficiency (Guo et al., 2020).
-DiskANN and SPANN propose memory-disk hybrid indexing systems that work beyond the limitation of memory capacity (Chen et al., 2021; Jayaram Subramanya et al., 2019).
-Other prior works propose hardware acceleration of vector index with GPUs (Johnson et al., 2019), FPGAs (Zeng et al., 2022; Jiang et al., 2023a), TPU (Chern et al., 2022), or ray tracing hardware (Zhu, 2022; Liu et al., 2023).
-BANG proposes a method to scale graph-based ANN beyond GPU memory (Khan et al., 2024), and Rummy allows the index to scale beyond GPU memory capacity with reordered pipelining (Zhang et al., 2024a).
-These methods either require algorithm modifications or are bottlenecked by CPU-GPU bandwidth.
-Our proposal, focuses on the context of modular RAG applications, where queries for retrieval are usually generated by LLMs, and optimizes the latency and the GPU memory consumption without altering the algorithm of the IVF index.
-7. Conclusion
-In this paper, we introduced TeleRAG, an inference system that tackles the system challenges of RAG pipelins under latency-sensitive scenarios.
-By using lookahead retrieval, which overlaps data transfer and GPU computation for faster retrieval, a profile-guided approach to determine optimal prefetching data amount, and GPU-CPU cooperation, TeleRAG speeds up the end-to-end latency with minimal GPU memory requirement.
-Our evaluation shows that TeleRAG significantly improves performance compared to existing state-of-the-art solutions.
+Table 4: Detailed configurations of our retrieval index.
+Table 4 shows the detailed configurations of our retrieval index used in the evaluation.